### v1

# Inferring Capabilities from Task Performance with Bayesian Triangulation

Abstract
As machine learning models become more general, we need to characterise them in richer, more meaningful ways. We describe a method to infer the cognitive profile of a system from diverse experimental data. To do so, we introduce measurement layouts that model how task-instance features interact with system capabilities to affect performance. These features must be triangulated in complex ways to be able to infer capabilities from non-populational data – a challenge for traditional psychometric and inferential tools. Using the Bayesian probabilistic programming library PyMC, we infer different cognitive profiles for agents in two scenarios: 68 actual contestants in the AnimalAI Olympics and 30 synthetic agents for O-PIAAGETS, an object permanence battery. We showcase the potential for capability-oriented evaluation.
1 Introduction
What does success or failure tell us about a system’s capabilities? In isolation, performance results are hard to interpret: most interesting tasks require several different abilities to be successfully combined. instances. For example, an agent exploring a 3D environment to find a reward needs to understand that the reward still exists when occluded (object permanence). It must then remember where it is, and successfully navigate to it. If the agent fails, we cannot know the cause: lack of object permanence, limited memory or navigation skills. This leads to poor predictability about the performance of the agent for new tasks, in or out of the training distribution.
Figure 1: This task instance requires the agent to: (a) understand that the reward still exists when occluded (object permanence), (b) remember which of the three walls it is occluded behind while it navigates towards it (memory), and (c) reach it over ramps and around the wall (navigation).
Figure 2: Triangulation in the measurement layouts using results from task instances like this one. Bottom-up inference from three tasks (in green and red for success and failure respectively) leading to the cognitive profile. Top-down inference (in blue) predicting failure for the fourth task.
However, suppose we also knew that the agent performs well on tasks that involve complex navigation, and the same for tasks involving memory. We could now infer that the system’s failure is likely a failure of object permanence by triangulation, a robust inferential method in formal epistemology (Heesen, Bright, and Zucker 2019) and causal (Bayesian) reasoning.
This bottom-up inference is illustrated in green and red in Fig. 2. We can use this information to draw conclusions about where the system is safe to deploy, and to direct efforts to improve it. We can also predict behaviour in new scenarios: for example, the agent will likely fail at a new task with high object permanence demands, as illustrated in blue in Fig. 2.
This example demonstrates that we can leverage the fact that different tasks—and their constituent instances—have different demands to help us extract conclusions about system capabilities from patterns of task performance. However, these conclusions require an analytical approach that facilitates complex inferences. If we were to examine each task or instance in isolation or to simply aggregate performance across all tasks, this triangulation of abilities would not be possible. To address this, we propose a novel Bayesian approach to evaluation that enables simultaneous inferences about multiple abilities from patterns of task performance. This approach has a series of principles facilitating a capability-oriented approach to the evaluation of machine learning models:
1. We do not simply break out performance by ‘categories’. Related work employs a ‘taxonomy of tasks’ (see e.g., Table 1 in (Liang et al. 2022)), each placed into one or more categories. This limits the inferential capacity and predictability of these category aggregates for new tasks. Instead, we evaluate based on instance demands.
2. Unlike psychometric approaches such as factor analysis (FA), structural equation modelling (SEM) or item response theory (IRT), we do not rely on populational data (Gustafsson and Undheim 1996). Instead, we infer the cognitive profile of a single subject from their performance data alone.
3. We do not ‘reify’ extracted latent variables as capabilities, which has been done hierarchically elsewhere, (e.g., the Cattell-Horn-Carroll model (Keith and Reynolds 2010)).
Instead, we identify general linking functions that help derive the requisite capabilities from the demands of the tasks.
4. Probabilistic models currently used in cognitive modelling (Lee 2011) are often simple at the level of the individual and cannot convey the required probabilistic expressions connecting individual performance, capabilities and task features. Instead, complex interactions allow for Bayesian triangulation which we use for informative evaluation. In this paper we find an operational solution to these inferential challenges in the form of Measurement Layouts: specialised and semantically-rich hierarchical Bayesian networks. Using the inference power of the No U-Turn Sampler (Hoffman and Gelman 2014) in PyMC (Salvatier, Wiecki, and Fonnesbeck 2016), a probabilistic programming engine, we can triangulate from performance and the task’s cognitive demands to the capability and bias levels for each subject (their cognitive profile). Then, performance can be inferred downstream for new tasks.
We derive the cognitive profiles in two scenarios: the contestants in the AnimalAI Olympics (Crosby et al. 2020) and several synthetic agents for O-PIAAGETS (Voudouris et al. 2022), an object permanence battery. We show how systems with similar overall performance can have very different capability levels and biases.
2 Background
AI systems are often evaluated on one or more standard benchmarks. Because of the remarkably broad capabilities of modern systems, benchmarks are becoming larger and more diverse. For instance, BIG-bench (Srivastava et al. 2022) contains over 200 tasks for evaluating language models. Aggregating many tasks into a single score does not allow for nuanced conclusions about where or why a system fails (Burnell et al. 2023), and has little value for predicting system performance a specific unseen task.
Analysing 200 tasks separately would lead to an unwieldy and hard to interpret set of results. One option is the categorisation of tasks across several dimensions, providing a moderate level of insight into where a model succeeds or fails (Osband et al. 2019; Ilievski et al. 2021; Crosby et al. 2020; Liang et al. 2022). However, conclusions are highly dependent on how instances are assigned to categories, and predictive value is limited inside and across categories.
An widely-used approach that does not require any taxonomy of capabilities is to extract and reify them from the data. Given experimental data for many tasks and subjects, we can use Factor Analysis (FA) to infer those factors that best explain the observed variance in the population. Some structure can be inferred from FA, as in the classical hierarchical models (Cattell-Horn-Carroll model (Keith and Reynolds 2010)), but today it is common to use more powerful techniques such as Structural Equation Modelling (SEM) (Ullman and Bentler 2012). However, SEM assumes linear relationships between variables and that dependent variables have normally distributed sample means. These assumptions are unlikely in the real world. Further, SEM is a populational technique requiring a large sample of individuals in order to confirm structural relationships.
A more granular approach is to identify the difficulty of each task instance and the ability of each system based on the relationship between task difficulty and the system’s performance. This idea is the cornerstone of Item Response Theory (IRT) (Embretson and Reise 2000; De Ayala 2009).
When several abilities exist, we must rely on multidimensional IRT(Reckase 2009) However, IRT is also populational, and derives abilities and difficulties relative to population averages, therefore the parameters of the same item can change if we add new items, or the abilities of a system may change if we add more systems. This poses a problem for AI evaluation, where the population of individual systems is constantly changing and advancing.
Still, with FA, SEM, IRT or other techniques, patterns of performance are mapped to capabilities, making up a multi-dimensional characterisation of the subject that is sometimes called a cognitive profile (Letteri 1980). Cognitive profiles make it possible to identify differences that might go unnoticed in aggregated scores, make predictions about future performance, and gain insights into the relationships between different abilities (e.g., deficits in mental rotation are one cause of poor spatial navigation).
Most psychological assessment is populational. Exceptions can mostly be found in developmental psychology and animal cognition (Moran et al. 2017; Bergman and Andersson 2015; Shaw and Schmelz 2017). The scarcity of subjects and the difficulty of experiments results in an emphasis on carefully designed tasks that ensure a systematic coverage of the relevant features and possible distractors. On the other hand, if the cognitive mechanisms for solving the task are accurate and enough variation is introduced, as in the example of Fig. 2, the results can be sufficient to estimate the capabilities of a single individual, independently of the other evaluated individuals. This is the type of approach that we need to evaluate ML models and AI systems.
Fortunately, there is recent movement towards producing training and test variations, augmented examples or corruptions to test systems, through adversarial learning (Zellers et al. 2018; Nie et al. 2019; Rozen et al. 2019; Kiela et al. 2021), procedural content generation (Risi and Togelius 2020) and other domain-specific techniques (Sugawara, Stenetorp, and Aizawa 2020). These variations have also been inspired by software testing, such as (Marijan and Gotlieb 2020) and (Ribeiro et al. 2020). Some of this effort results in annotated datasets, especially for reinforcement learning — from the inspiration behind the AnimalAI Olympics (Crosby et al. 2020) to the fully-systematic O-PIAAGETS (Voudouris et al. 2022). However, the next step of using these annotated data for capability-oriented evaluation has not been taken yet.
Not using populational data has to be compensated by a larger number of varied instances. When many dimensions of variation are introduced, the inference problem becomes complex. Traditional Bayesian Networks (BN) (Ben-Gal 2008) can be used for arbitrarily sophisticated probability distributions. For complex networks, though, performing the inference calculation is computationally intractable (Cooper 1990). Approximation techniques for Bayesian inference (BI) have been introduced, including the Markov Chain Monte Carlo family (Andrieu et al. 2003), allowing for the inference of latent values in more complex BNs, particularly those with multiple levels of dependency and complex relations between variables. Taking advantage of these recent advances in Bayesian approximation, evaluation procedures, and computing power allows us to extend the benefits of capability-oriented evaluation to machine learning for the first time, enabling more robust, thorough, and principled evaluations of ML systems.
3 Measurement Layouts
Performance can be understood as a function of task demands and capability levels. The precise relation between the two depends on the task features and the tested capabilities. For instance, in a memory problem, the number of objects to remember in a particular task demands a memory capability level at least as high as the number of objects. However, the difference between task demands and capability levels is likely not sufficient to account for all variance in observed performance. Cognitive biases, such as a preference or aversion to apparently irrelevant features like colour, can also affect performance. Finally, AI systems can exhibit a lack of robustness caused by some unaccounted factors or noise. Cognitive profiles need to include all of these. Let us define more precisely a task characterisation and the corresponding cognitive profile. We will then detail how they are connected by the measurement layouts.
Definition 3.1.
A task characterisation is a set of observable, usually constructed, meta-features $X$, expressing cognitive demands and other high-level properties of the task.
For instance, in an image classification problem, the level of blur of the image or the level of clutter of the background could constitute perceptual demands. Many other properties may also, in practice, affect performance.
Definition 3.2.
The cognitive profile of a system is a tuple $\langle C,B,R\rangle$. Where $C$, $B$, and $R$ are vectors of capability levels, bias, and robustness values respectively.
Capability levels represent what the system can do, bias values represent some other preferences or limitations that may impact performance in a less monotonic way, and robustness levels account for reliability issues, unexplained or random effects (noise) on either the agent or environment.
Definition 3.3.
A measurement layout is a directed acyclic graph that connects, through linking functions, the meta-features of a task characterisation with the cognitive profile of a system, in order to predict observed performance.
We define measurement layouts as Hierarchical Bayesian Networks (HBN) (Murphy 2023) with the following characteristics: Within the measurement layout we include a node for every meta-feature in every task instance and a node for each element of the system’s cognitive profile. All of them are roots of the graph. A connection $A\rightarrow B$ denotes that $B$ is conditionally dependent on $A$. Linking functions are used to formally define relationships between dependent nodes.
All nodes within the HBN define a probability distribution. We can classify nodes into three categories depending on the provenance of their summary statistics:
Metafeatures: These come from the task characterisation and are fixed, observable values. They are represented as a fixed Dirac-delta distribution.
Cognitive Profile Nodes: These are elements of the cognitive profile relevant to assessed task’s demands. This includes required capabilities and likely biasses. These often combine with meta-features and feed in to derived nodes. The range and distribution of these elements derives from the task characteristics. To make our inferences as data-driven as possible these begin with uninformative priors.
Derived Nodes: These are instance-level inferences. They are combinations of meta-features and/or cognitive profile elements that have their parameters derived from dependent nodes. Information to compute these values is encoded in the linking function for the node, the output of which is used as the summary statistics of the node’s distribution.
One derived node is the observable performance for the task—the only leaf node—but most are Intermediate Non-Observable Nodes (INON), which represent intermediate performances or effects. All three node types were seen in Fig. 2 (right). The connective structure can be enriched by domain knowledge determining which cognitive profile elements and meta-features are expected to affect the values of instance-level inference nodes. These relationships can be higher-order and complex, with certain INON nodes requiring the nuanced confluence of many cognitive profile elements and meta-features. Complex connective structure (and the requirement for increased levels of domain knowledge) can arise from the need to accurately distinguish between success, failure, and the multitude of confounding behaviour that a carefully designed benchmark must account for.
By a linking function, we refer to any mathematical expression that maps values from the outputs of one node’s probability distribution to the input of another.
For instance, it could be a sigmoid function of the difference between capabilities and demands. It could also be a product of (non-compensatory) capabilities, representing a quantitative and differentiable expression for a logical ‘and’. Alternatively, it could be some additive linear equation, representing a quantitative expression of a logical ‘or’, when some preceding nodes are compensatory—or a generalised mean in between. The functions can scale the incoming nodes, so that we adjust the ranges for capabilities, biases and noise. The possibilities are endless; in practice we use only a few kinds of linking functions. We give the specific equations for the linking functions that we used used in this work in Tables 3 and 5 in the appendix. These are used in the two scenarios that follow.
A key conceptual difference between our measurement layouts and typical use cases for HBNs is that our approach is trying to capture a hierarchical dependency relation on capabilities and demands, rather than encoding information on hyper-priors. The linking functions between nodes are often non-linear, conceived from domain knowledge of the task and applying ideas from cognitive modelling.
For instance, one of the key decisions is to determine when some capabilities are independent, and if so, whether they are compensatory or not, being realised with additive or multiplicative expressions respectively. If they are not independent, then an arrow has to connect them with a linking function expressing this dependency mathematically.
Given the cognitive profile of a system, and the characterisation of a new task, we can use the measurement layout to anticipate performance by forward (top-down) inference. But before that we need to infer the cognitive profile from the observed performance of other tasks by backward (bottom-up) Bayesian inference. For this, we utilise PyMC’s inference engine and the No U-Turn Sampler (Hoffman and Gelman 2014) approach to BI (see Appendix A for a detailed description). For this to accurately predict a system’s cognitive profile, the test battery
requires instances to jointly control for alternative explanations permitting performance results to ‘triangulate’ latent capabilities.
4 Cognitive Profiles for the AAI Olympics
To illustrate how a measurement layout identifies cognitive profiles, we first select a domain with simple task characterisations that can be straightforwardly related to cognitive profiles to predict performance.
The Animal-AI Olympics (AAIO) (Crosby et al. 2020) meets all these requirements. The AAIO was an open competition evaluating AI agents in a 3D environment, where agents had to get a reward by locating it and navigating in an arena.
Participants were asked to build systems (usually reinforcement learning agents) and train them on their own custom-built tasks in the AAI Environment. Screenshots for two tasks are shown in Fig. 7 located in Appendix B along with more information about AAIO and the environment. We select 69 tasks from the competition test set that evaluate simple goal-directed behaviour. We include 68 participant agents for which we know nothing about their design.
4.1 Measurement Layout
In this domain, following the analysis performed in (Burnell et al. 2022), the set of meta-features $X$ of the instances that characterise task instances are rewardSize, rewardDistance, Xpos and rewardBehind. These represent the reward’s size, distance, and if it is left, right, in front, or behind the agent.
The cognitive profile is composed of two capabilities, navigationAbility and visualAbility, respectively representing the skill to move around purposefully and the visual acuity of the agent, one bias rightLeftBias, representing any preference for left or right movements, and one robustness level noiseLevel, accounting for any unexplained variability.
We then compare the profile of the agent to the demands of the task to derive performance. This is our measurement layout, as can be seen in Fig. 4 (left). The metafeatures $X$ appear as observed (shaded) nodes in the model, appearing within the plate, with 69 instances. The elements of the cognitive profile appear as external latent (unshaded) variables.
The derived nodes lie at the intersection of the meta-features and the external cognitive profile nodes.
Figure 3: Measurement layout for the Animal AI Olympics.
Figure 4: Cognitive profiles of the 68 agents in AAIO, with names shown for some of them. The $x$-axis and $y$-axis show the inferred means for navigationAbility and visualAbility respectively, with their standard deviations as error bars in grey. The radius of each point represents the average performance, while the colour represents the noiseLevel, with red representing higher noise than blue.
Not visible in the figure are the linking functions associated with each arrow. The variable rightLeftEffect is calculated as the product between rightLeftBias and XPos, since left/right bias and XPos are represented with negative and positive values respectively, so the product gives a positive result when both go in the same direction. This rightLeftEffect, jointly with rewardBehind and rewardDistance affect navigationPerformance, a non-observable intermediate performance variable. The exact formulation of how these three variables are integrated can be found in Appendix C.2,
but broadly the linking function subtracts the effects of reward distance and position relative to the agent’s viewpoint (e.g., behind) from the navigation ability and adds the right-left bias to this. The linking function then applies a logistic function to this margin to give a value between 0 and 1 for navigationPerformance. The process to derive visualPerformance is similar, simply subtracting rewardSize ( inverted so the higher the smaller) from visualAbility.
Note that the ranges of these two abilities are given by the ranges of the task metafeatures, with navigationAbility in $[0,5.3]$ and with visualAbility in $[0,1.9]$ respectively, representing how distant and how big a reward—with their corresponding distance and size units—the agent can obtain. Finally, taskPerformance is Bernoulli-distributed on a product of navigationPerformance and visualPerformance adding the noiseLevel. We use the product because navigation and visual acuity are non-compensatory (i.e., navigation skills are of no use if the agent cannot see, and vice versa).
4.2 Inference and Analysis
Using the measurement layout and the 69 instances, we can use PyMC to infer a very simple cognitive profile for each of the 68 entrants of the competition. We run the NUTS inference algorithm with two sampling chains of size 1000 each, finding no inference divergences.
Fig. 4 (right) summarises the inferred capabilities and noise levels. We see the two main capabilities are correlated for this population, going from low performance in the bottom left area (e.g., ‘ice-play’ with 4% accuracy) to high performance in the top right area (e.g., ‘Trrrr’ with 99% accuracy). In these extremes, capabilities are nonexistent (close to 0) or saturated (to the maximum 5.3 and 1.9 respectively), as expected. But in the middle range, there are agents that have noticeable differences between their capabilities (e.g., ‘cocel’). Some have strong biases, such as ‘Thursday’ and ‘KoozytHiperdyne’ with left-right bias values of $1.14$ and $-0.78$ respectively (not shown on the plot). This can have an effect on navigation ability but not on visual ability.
The standard deviations of the abilities, shown as error bars, and the level of noise are useful to explain situations both when behaviour is conformant or not with the task characteristics. For instance, ‘Juohmaru’, with 54% accuracy, is a very conformant agent: performance decreases as rewards get smaller and further away, and when they are initially located behind from the agent. We see that navigation ability is $2.50\pm 1.08$ and visual ability is $0.92\pm 0.44$, with a medium level of noise $0.26\pm 0.17$. But ‘sparklemotion’, with 36% accuracy, is very non-conformant, with abilities $0.40\pm 0.38$ and $0.59\pm 0.45$, and higher noise $0.31\pm 0.13$. Appendix C contains full characteristic grids of these two agents in Fig. 8 (demonstrating the conformant / non-conformant behaviour) and all inferred results in Fig. 9.
Note that we have inferred two different capabilities for each agent as well reliability measure of these estimates: how well they can predict performance (depending on the quality of the estimates, biases and remaining noise) for each particular instance given its meta-features. We do this without having different sub-batteries for each capability.
Our model can also predict performance on unseen tasks instances. To evaluate this, we randomly held out 20% of the task instances from the backwards inference stage and evaluated the predicted probabilities of success on them. Due to the small size of the dataset, we repeated this process 15 times and took the mean Brier score for each agent. Fig. 5(a) summarises these scores, which are lower than predictions based on aggregate performance for agents with success rates in the range $[0.2,0.7]$, exactly where there is variability to explain. Full results and discussion in Appendix E.
5 Cognitive Profiles for Object Permanence
Our analysis of the AAI Olympics used a relatively simple scenario as a proof of concept. But one limitation of this analysis is that we do not know the ”ground truth” of the agents’ abilities, so it is difficult to determine precisely how effective our approach was. It also remains unclear how well our approach works in complex scenarios with many abilities and metafeatures. Therefore, we next sought to demonstrate the scalability of our approach by analysing data from a more complex scenario—a rich test battery for evaluating Object Permanence. To ensure that we knew the ground truth abilities of the agents, we created a synthetic dataset by simulating the performance of 30 hypothetical agents based on specified cognitive profiles 111because of the complexity of the abilities that affect performance on this testbed, it would be extremely difficult to create a varied population of real agents with specific levels of abilities. This scenario tests whether our methods can recover complex cognitive profiles from performance data of agents for which we know the ground truth.
To do so, the authors of this paper were divided into two teams, Team A and Team B. Team A generated the synthetic dataset, while Team B defined the measurement layout and attempted to recover the cognitive profiles of the synthetic agents. Both teams were familiar with O-PIAAGETS and the general abilities, bias and elements that are combined in the battery. This two team approach allowed us to generate synthetic agents while keeping Team B blind to their profiles, removing the potential for bias and making the modelling effort more akin to a real-world scenario.
5.1 The Testbed and Synthetic Agents
We applied our measurement methodology to the Object-Permanence In Animal-AI: Generalisable Test Suites (O-PIAAGETS, (Voudouris et al. 2022)). Object Permanence (OP) is the understanding that objects continue to exist when they go out of view. Behaviourally, an agent has object permanence if they behave as though objects continue to exist even when they cannot directly perceive them. It is a key component of physical common-sense (Shanahan et al. 2020; Lake et al. 2017; Baillargeon et al. 2011), allowing biological agents to more successfully predict and interact with their environment. Object permanence is well studied in cognitive science, where several experimental designs are used to detect its presence in human and non-human animals (Scholl 2007; Flombaum and Scholl 2006; Hare and Tomasello 2005; Chiandetti and Vallortigara 2011). Many of these paradigms are represented in O-PIAAGETS, which contains over 13,000 instances. We selected three paradigms that test for allocentric object permanence, where objects are occluded independently of the observer’s actions. We also include basic control tasks within the test-suite to evaluate the agent’s non-OP capabilities and make the inference process more robust. Details about the test-bed design criteria, the included tasks, and why object permanence is difficult to identify are given in the appendix in D.1 and D.2.
Figure 5: Visual representation of the measurement layout for the object permanence task.
Team A generated 30 synthetic agents of four different types to ensure a diversity of behaviour. Reference Agents act as reference points, e.g., a perfect agent, an agent with $x$success, etc.
Achilles Heel Agents have low ability on some dimension, but otherwise perform well and can be said to have object permanence.
Context-Specific Agents have mixed OP performance, doing well in some specific tests/contexts but not others.
Fraudster Agents behave in a sophisticated manner that might look like object permanence, but actually is not (Uzgiris and Hunt 1975). Tables  6 and  7 in appendix D.4 show the specifics of each agent. This information was not disclosed to Team B.
5.2 Measurement Layout
Object permanence is a difficult capability to evaluate—combined with the need to visually identify goals, navigate to the goal, avoid hazards, and to remember the location of the object; there are many possible factors influencing overall performance that need to be disambiguated.
We present the measurement layout for the OP tasks in Fig. 5. The same broad structure as for the AAIO measurement layout is expanded upon. In the OP task-suite, there are additional challenges to navigation in the form of ramps, platforms, and lava. Subsequently, nodes for the agent’s capability (and instance-level performance) at handling each of these must be added to the layout. Meta-features for the presence of each of these challenges are needed. To represent object permanence specifically, we add extra meta-features: Occluder Presence, Time Under Occlusion, and Number of Positions (denoting the number of search-location choices available to the agent in forced choice instances). We further add capability nodes for Object Permanence and Memory Ability (each with an associated Performance node). Memory Performance, Occluder Presence and Number of Positions all feed into object permanence performance. The linking function for OP performance is the non-complementary product of Memory Performance and the logistic of the OP margin. The OP margin is the difference between Object Permanence and the product of Number of Positions and Time Under Occlusion. If no occluder was present the linking function for Object Permanence Performance returns $1$.
5.3 Inference and Analysis
Team B took the synthetic agent results and inferred the cognitive profiles using the PyMC implementation of the measurement layout. Overall, our model was successful at inferring system capabilities from patterns of performance. We evaluated the model in three ways. The first was a qualitative comparison of the cognitive profile descriptions generated by Team B with the criteria used to generate the agents. The model facilitated a correct description of the agent in around 75% of cases. For the “Achilles Heel” agents in particular, which had object permanence abilities but a specific weakness, the model was highly successful at identifying the cause of failure. It was, however unable to spot “fraudsters” making use of strategies (such as “go to the last seen location of the reward”) to mimic OP capability. Appendix D.4 contains a detailed breakdown of inferred parameters of the cognitive profiles of each agent, more information on the experimental set up, and the full qualitative analysis comparing inferred capabilities to the ground truth capabilities of the agents designed by Team A.
Capability OP FlatNav Visual Lava Platform Ramp Memory
Overall RMSE 0.28 0.18 0.16 0.2 0.28 0.38 0.15
Included RMSE 0.13 0.11 0.24 0.17 0.134 0.27 0.16
Table 1: RMSE between the ground truth cognitive profile for Team A’s synthetic agents and the inferred values by Team B’s measurement layout. Overall: All 30 agents. Included: 13 agents generated using only the features included in the model.
(a) AAIO Tasks
(b) O-PIAAGETS tasks.
Figure 6: Brier score of model predictions against agent success rates.
The second validation involved calculating the RMSE of the model for each ability (normalised to be in $[0,1]$), relative to the ground truth generated by Team A. Given in Table 1, model errors were modest across most abilities. Understandably, the model struggled most with agents that were created using features not captured by the model. For example, agent 23 was generated to have robust object permanence only in cup tasks (passing only 10% of other types). As our model did not include task type as a meta-feature, it erroneously took the non-cup task failures as evidence the agent lacks OP as well as struggling with platforms. However, given that the cup tasks are the only OP tasks without platforms, this is not an unreasonable assessment. By contrast, as the bottom row of the table shows, the model was accurate for agents generated using only features included in the model. This highlights the importance of having a measurement layout capturing the important meta-features that most affect performance, and also of designing test-batteries that allow confounding factors to be disentangled.
For our third validation, we show that our model can be used for accurate prediction of agent success on unseen task instances. Taking instance meta-features and applying forward inference with our model yields a probability of success. We compare this to the probability predicted by aggregate success using a Brier score. The full results are given in Appendix E, but a summary is given in Figure 5(b). The model is consistently a better predictor of success than relying on the aggregate measure. We see better prediction performance for agents with less certain success rates. These are the agents for which we are more interested in improving prediction as there is more capacity to reduce uncertainty.
6 Discussion
Increasingly more machine learning systems are (pre-)trained once and deployed for a range of tasks, Task-oriented evaluation based on aggregated performance cannot robustly evaluate these systems across the many situations in which they might be deployed, nor predict behaviour for new tasks and changing distributions. However, we can often identify abstract capabilities that are not directly measurable, such as ‘object permanence’ in agents or ‘handling negation’ in language models. In the behavioural sciences this is a common approach, but many techniques are populational. Triangulation for single individuals by contrasting against the task demands is common in animal cognition research, but it is unusual to see complex inferences from thousands of items, as is common in machine learning. In this paper we have integrated several elements to make this inferential exercise possible and shown it can identify the capabilities and biases of each model—its cognitive profile—independently.
This unleashes a new way of evaluating general-purpose systems, by identifying areas of weakness or reasons for task failure. This is especially useful for hierarchical capabilities with multiple levels of dependency. We can speed up the debugging process by isolating skill-gaps or biases responsible for failure, and improve confidence in predictions of performance on new tasks. Instead of hoping that the average performance will hold on new tasks, our approach enables the use of task demands and the inferred cognitive profile to determine whether system deployment is safe and worthwhile.
There are several limitations of this work. If many parameters must be inferred simultaneously then many evaluation instances are needed. We advocate incremental approaches, where each capability is inferred with subsets of tasks before moving to tasks with more complex dependencies. Nonetheless, we have shown that even a posteriori, as in the AAIO, inferences can be achieved simultaneously. Here we covered only the domain of agentic systems, but the same ideas can be extrapolated to language, vision and other domains. Ultimately, the biggest challenge for the construction of measurement layouts is the need for detailed domain knowledge and robust evaluation batteries. Both can be costly, but are urgently needed to improve AI evaluation. Currently, evaluation is focused on advancing SOTA performance on an increasingly small selection of benchmarks that often lack construct validity and are limited in scope (Koch et al. 2021; Raji et al. 2021).
Our approach also highlights the need for a clear distinction between training datasets, which can be massive, unannotated, and built in an adversarial way (Kiela et al. 2021) (Jia and Liang 2017; Zellers et al. 2019), and evaluation batteries, which should be devised carefully with systematic variation across the elements that might affect performance. Our vision is that, from well-constructed batteries such as O-PIAAGETS, building measurement layouts and triangulating with them should become routine.
Appendices
Here we include some supplementary material to the paper. First we include more details about PyMC, the Bayesian inference engine we are using and the decisions made for the estimation algorithm, and then we give more details and results for the two scenarios; first, the Animal AI Olympics and second, O-PIAAGETS.
Appendix A PyMC
PyMC (Salvatier, Wiecki, and Fonnesbeck 2016) is a probabilistic programming language for building Bayesian models and fitting them with Markov Chain Monte Carlo (MCMC) methods (Andrieu et al. 2003). Complex Bayesian models often require solving intractable integrals for inference such as the normalisation constant in Bayes’ theorem,
[…]
among others. Monte Carlo (MC) Integration provides a method to approximate these intractable integrals using random sampling.
In order to evaluate
[…]
we can rewrite it as an expectation
[…]
for any density function $p$ such that $p(x)>0$.
MC Integration takes a random sample $X=\{x_{i}\}_{i=1}^{N}\in\Omega$ with distribution $p(x)$ and uses the Law of Large Numbers to approximate the integral as:
$I\approx\frac{1}{N}\sum_{i=1}^{N}\frac{f(x_{i})}{p(x_{i})}$
MC Integration is reliant on being able to sample efficiently from $\Omega$ according to $p(x)$. MCMC algorithms aim to do this, scaling to complex, hierarchical models by building a Markov Chain with equilibrium distributions approximating the desired probability density function $p$. Sampling from the Markov Chain can then be used as a stand-in for sampling $p(x)$. The prototypical example of an MCMC algorithm is the Metropolis Hastings algorithm (Hastings 1970). Hamiltonian Monte Carlo methods (Betancourt and Girolami 2013) improve on typical MCMC algorithms by replacing the random-walk element exploring the Markov Chain state-space with Hamiltonian Mechanics.
PyMC requires all expressions to be differentiable, which means that functions such as maximum and minimum are implemented as generalised means, and logical operators are implemented with products and sums.
A.1 Measurement Layouts in PyMC
In our work, we utilise the No U-Turn Sampler (Hoffman and Gelman 2014) built into PyMC, an extension to Hamiltonian MC methods which eliminates the need to set particular hyper-parameters.
With the NUTS sampler we use 2 sampling chains, tuned for $1000$ samples that are discarded in order to optimise the step-size within NUTS, and then draw $1000$ samples to estimate the posterior.
We define the model once for each scenario and then we estimate the parameters of the cognitive profile and some other INON elements for each system being evaluated independently.
As mentioned in the introduction, data and code will become available upon publication.
Appendix B Animal AI Environment and Competition
Animal-AI is an environment for training and testing AI systems. It is built in the Unity engine (Juliani et al. 2018). Animal-AI provides a first-person perspective in a 3D environment. The environment can contain a number of pre-specified objects, representing rewards, obstacles, tools, or dangers. Tasks can be defined using a domain-specific language to procedurally generate a large number of tasks instances. The agent’s interaction with the world is limited to moving around in the 3D space, though it can exploit physics interactions to exert more influence over the environment. Fig. 7 shows two screenshots from separate task instances.
In the Animal-AI Olympics (Crosby et al. 2020), an example arena containing the possible objects was provided to contestants. They were instructed that their agents would be evaluated on a variety of problems constructed from the objects in the example arena. The contestants then needed to build an agent—using any method of their choosing—that they believed would succeed at a wide variety of tasks. The tasks that ultimately made up the evaluation suite were based on adaptations of tasks from comparative and developmental literature. More information can be found at (http://animalai.org/AAI/).
For our work with the AAIO data-set we utilised a subset of tasks instances that were solvable with simple goal-directed behaviour. The specific task instances we utilised were:
[1-1-1, 1-1-2, 1-1-3, 1-2-1, 1-2-2, 1-2-3, 1-3-1, 1-3-2, 1-3-3, 1-4-1, 1-4-2, 1-4-3, 1-5-1, 1-5-2, 1-5-3, 1-6-1, 1-6-2, 1-6-3, 1-7-1, 1-7-2, 1-7-3, 1-8-1, 1-8-2, 1-8-3, 1-9-1, 1-9-2, 1-9-3, 1-10-1, 1-10-2, 1-10-3, 1-11-1, 1-11-2, 1-11-3, 1-12-1, 1-12-2, 1-12-3, 1-13-1, 1-13-2, 1-13-3, 1-14-1, 1-14-2, 1-14-3, 1-15-1, 1-15-2, 1-15-3, 1-16-1, 1-16-2, 1-16-3, 1-17-1, 1-17-2, 1-17-3, 7-1-1, 7-1-2, 7-1-3, 7-2-1, 7-2-2, 7-2-3, 7-3-1, 7-3-2, 7-3-3, 7-4-1, 7-4-2, 7-4-3, 7-5-1, 7-5-2, 7-5-3, 7-6-1, 7-6-2, 7-6-3]
The types of arenas that these instances correspond to can be found at (http://animalai.org/AAI/testbed).
Figure 7: Screenshots from two different instances in the Animal AI Olympics that we include in our analysis.
Appendix C AAIO Agent Results
This section gives more details for the first domain we examined in Section 4, the Animal-AI Olympics competition dataset. We also give more detail about the specifics of the measurement layouts.
C.1 Patterns of Performance
Figure 8: Agent characteristic grid displaying results for some selected agents. Top row, from left to right: ‘ironbar’, ‘sparklemotion’, ‘Juohmaru’. Bottom row, from left to right: ‘y.yang’, ‘daydayup’, ‘41Animals’. Note that the scales of the $x$-axis and $y$-axis are normalised and not the same as in the main body of text.
To show the patterns of performance of several agents, we use a simple visualisation known as agent characteristic grids, as described in (Burnell et al. 2022). The plots place each feature as a dimension and the success rate represented in colours from green (success) to red (failure). Values are grouped into bins of appropriate width in the desired dimensions for analysis. Each cell also shows the number of observed instances in that bin. Fig. 8 shows some of the agents discussed in the paper, as well as some other agents that will serve us to illustrate a few points.
For instance, ‘ironbar’ has 99% accuracy (the grid is mostly green everywhere) for this set of tasks (like ‘Trrrr’, the winner of the competition) and is identified as saturating both capabilities in Fig. 4. The next one, ‘sparklemotion’, with 36% accuracy, is very non-conformant, with navigationAbility $0.40\pm 0.38$ and visualAbility $0.59\pm 0.45$, and high noise $0.31\pm 0.13$. Proportionally, navigationAbility is worse as the grid looks even less conformant on the distance of the reward than the reward size. The rightmost on the top row is ‘Juohmaru’, with 54% accuracy, which is a very conformant agent, with red concentrating on the top and right of the grid (the objectively most difficult instances). Its navigation ability is $2.50\pm 1.08$ and visual ability is $0.92\pm 0.44$, and medium noise: $0.26\pm 0.17$.
We have specifically chosen some particular agents in the second row for which the extracted cognitive profile does not capture the whole performance patterns completely.
Leftmost on the second row, ‘y-yang’ has 70% accuracy and a very strange grid, with red cells in a diagonal, indicating a very non-conformant agent. This translates into navigationAbility $3.82\pm 0.45$ and visualAbility $1.68\pm 0.21$, but medium noise $0.22\pm 0.17$. This is so because overall the agent has high accuracy, so a bit of noise can explain the few anomalies on the top right of the grid.
In the middle of the second row is ‘daydayup’ with 52% accuracy. This agent shows a higher navigation ability than visual ability, as can be seen in the grid, where the size of the reward is the predictive feature and reward distance is less relevant. This translates into navigationAbility $3.48\pm 1.28$ (high than average but with uncertainty) and visualAbility $0.92\pm 0.44$ (low), with some a high level of random noise affecting performance $0.48\pm 0.20$.
Finally, ‘41Animals’ with 87% accuracy, shows relatively high values and low deviations for both navigationAbility $4.37\pm 0.59$ and visualAbility $1.81\pm 0.11$, with very little noise $0.05\pm 0.04$ despite having some non-conformant behaviour in the diagonal. In this case the left-right bias is higher ($0.45\pm 0.91$) and may account for this.
In general, for those agents where accuracy is too low or too high, the differences are more subtle. Of course, the interesting part of Fig. 9 is the agents in the middle. This suggests that estimating capabilities can benefit from a more focused analysis of instances around the demand thresholds. For ’Juohmaru’, this would amount to selecting instances by difficulty, instead of simply selecting all instances for analysis.
C.2 Animal AI Olympics Measurement Layout
Table 2 provides more information about the task meta-features that we utilised to construct the measurement layouts. Note that these ranges determine the ranges of the capabilities, so that we have units. For instance, if distance to the reward were measured in metres then the navigation capability would have metres as a unit. Let the navigation capability of an agent be, say, 20 metres. Since it is defined with a logistic function of distance, that means that the agent can achieve excellent performance when well below 20, performing at chance around the 20 metre mark, and failing every instance when the reward is well beyond 20 metres away.
Dimension
Description
Range/Values
Reward Size
The size of the reward (the sign is modified so lower is larger)
$[0,1.9]$
Reward Distance
Euclidean distance from the agent’s start position to the reward
$[0,5.3]$
Reward Behind
Whether the reward originates behind (1), left or right (0.5) or in front (0) of the agent
$\{0,0.5,1\}$
XPos
Whether the reward is to the left ($-1$), the right (1) or is centred (0) relative to the agent’s start position
$\{-1,0,1\}$
Table 2: AAIO Task Characterisation details. The dimensions on which instances vary in our measurement layouts.
Name Type Range/Values Prior / Linking Function
rewardSize Meta-Feature $[0,1.9]$
rewardDistance Meta-Feature $[0,5.3]$
rewardBehind Meta-Feature $\{0,0.5,1\}$
XPos Meta-Feature $\{-1,0,1\}$
navigationAbility Capability $[0,5.3]$ $ScaledBeta(1,1,0,5.3)$
visualAbility Capability $[0,1.9]$ $ScaledBeta(1,1,0,1.9)$
rightleftBias Bias [,][$\infty$,$\infty$][ - ∞ , ∞ ] $\mathcal{N}(0,1)$
noiseLevel Robustness $[0,1]$ $Uniform(0,1)$
rightleftEffect INON $[-1,1]$ ${\sf rightleftBias}\times{\sf XPos}$
navigationPerformance INON $[0,1]$ $\sigma({\sf NavigationAbility}-{\sf rewardDistance}\times(\frac{1}{2}{\sf rewardBehind}+1)+{\sf rightleftEffect})$
visualPerformance INON $[0,1]$ $\sigma({\sf visualAbility}-{\sf rewardSize})$
taskPerformance Observed $[0,1]$ $p\sim Bernoulli(\omega[{\sf noiseLevel}]({\sf navigationPerformance}\times{\sf visualPerformance},\nu))$
Table 3: Nodes in the Animal AI Measurement Layout, including range/values, priors and linking function.
The measurement layout for this suite of tasks consists of two capability nodes (navigation and visual acuity), one bias node (whether the goal is initially on the left or the right of the agent), and one robustness node (stochastic noise on success/failure). Each of the capability and bias nodes has an associated instance-performance node, its intermediate non-observable node (INON). The noise is applied at the end.
In Table 3 we have complete details of all the nodes in the measurement layout. The function $\sigma$ represents the standard logistic function:
$\sigma(x)=\frac{1}{1+e^{-x}}$
while $\omega$ represents a weighting:
$\omega[\alpha](b,c)=(1-\alpha)\cdot b+\alpha\cdot c$
The value $\nu$ represents a noise prior that represents a constant model set to 1 -- average performance. Finally, $ScaledBeta$ represents a scaled beta distribution to an interval different from $[0,1]$. Note that a beta distribution has two parameters that need to be estimated.
Figure 9: Inferred cognitive profile elements, along with the success rate, for the 68 agents in the AAIO. The gradient colour red-yellow-green denotes how high the means are (green is highest) while the gradient circles white-grey-black denote how high the standard deviations are (black is highest).
Finally, Fig. 9 shows the complete inference results for the 68 agents in the AAIO. The agents are ordered by the aggregate success rate at the collection of task instances. We see that, for the agents in the middle of the table in particular, our approach is able to decompose agent capabilities at this task into the constituent navigation and visual abilities. Bias has a more diverse behaviour throughout the table, with strong biases for extremely low scoring agents too, such as with agents ‘ahorizon’ and ‘Octopus’.
Equally, we can see that the detected bias can account for poor performance. Compare ‘Thursday’ and ‘MadStorks’ which share a similar overall success rate (0.30 vs. 0.28) despite the fact that ‘Thursday’ has much more navigation ability than ‘MadStorks’ and they have similar visual abilities. This is explained by ‘Thursday”s much larger bias towards moving to the right (0.73 vs 0.4) even if the reward is to the left.
Appendix D O-PIAAGETS Battery
Now we look at the object permanence test battery in more detail. We look at the overall design criteria for the test-bed, the types of tasks involved, and how the tasks are characterised. We also give full details for the measurement layout devised for this task and an extended view of the results inferred by our approach.
D.1 Design Criteria
Object Permanence (OP), and similar cognitive capabilities such as episodic memory and theory of mind, are difficult to robustly identify in AI systems because there are often several competing alternative explanations for why an agent is succeeding or failing at a particular task. We offer Bayesian triangulation as a solution to this problem.
Imagine a simple task where a reward is observed to roll leftwards behind an occluder, and go fully out of sight. An agent with OP ought to pass this task, since it: (a) wants to obtain the reward, (b) understands that the reward continues to exist despite not being able to see it, and (c) is able to infer the location of the occluded reward based on its previously observed trajectory. However, an agent with OP might fail this task, if for instance it is bad at navigating, and takes a very circuitous route to its goal and runs out of time before it can reach the reward. Conversely, an agent lacking OP might pass this task, simply because it has been rewarded previously for going forwards some number of steps and then turning left. This is the problem of the underdetermination of theory by behaviour (Stanford 2006), in which an observed behaviour can be explained in multiple ways. The use of well-informed measurement layouts based on internally valid test batteries allows us to more robustly eliminate incorrect alternative explanations and triangulate on the underlying causal explanations of AI system behaviour.
There are many competing desiderata to consider when designing a test-suite for the evaluation of OP in AI systems.
  1. 1.
    Cognitive capabilities such as OP can be difficult to place on an ordinal scale. This is because it is often not clear what affects behavioural performance in humans and other animals, and to what extent certain instance dimensions are relevant to good performance.
  2. 2.
    The relationships between cognitive capabilities can be complex and heterarchical, with single instances testing multiple cognitive capabilities with complex interactions. A carefully thought-out measurement layout is required, informed by cognitive science.
  3. 3.
    Instances may not be defined on all dimensions, introducing the possibility of missing data.
  4. 4.
    Instances are not necessarily evenly distributed across the levels of a dimension, leading to the potential for sampling biases.
These considerations were used as design criteria for the test-bed presented in the next subsection.
D.2 Object Permanence Tasks
The tasks that we use in our OP test-bed are based on a subset of the O-PIAAGETS object permanence test battery introduced in (Voudouris et al. 2022). The battery in its entirety is vast (including over 13,000 task instances requiring object permanence and over 250,000 basic task instances that don’t require OP, but help control for other potential explanations. Here, we adapt the following three separate task types from O-PIAAGETS.
Primate Cognition Test Battery Cup Tasks
Herrmann et al. developed the Primate Cognition Test Battery (PCTB) for testing aspects of physical and psychological common-sense in primates (Herrmann et al. 2007). One task in this battery was designed to assess OP. The participant is presented with three identical inverted cups, under which the experimenter hides a reward. Once the reward has been occluded, the participant is invited to lift a cup. If they fail to locate the reward, the trial ends. The reasoning here is that if the participant correctly locates the reward without any evidence of trial-and-error, there is evidence that they have OP, understanding that the reward continues to exist even when it is out of view.
Figure 10: Primate Cognition Test Battery 3-Cup (’Cup’) Task. Top: Top-down view; bottom: Agent-view. Red arrow indicate reward location, grey arrows indicate agent location.
O-PIAAGETS has a battery of 598 instances based on this experimental set-up, including 359 that require object permanence to be solved and 239 that do not. In the OP tests, the agent observes a green reward drop from a height behind an occluder. They must traverse a ramp, sometimes avoiding lava, to obtain it. If they traverse the wrong ramp, they cannot return, and they will fail the instance, making this a forced choice test. See Fig. 10. In the instances that do not require OP, the reward is always visible, either through transparent objects or due to the lack of occluders. These instances serve as controls, so that we can evaluate whether poor performance on tasks requiring OP can be explained not by a lack of object permanence per se, but rather by poor performance traversing ramps, avoiding lava, or navigating towards rewards.
Primate Cognition Test Battery Grid Tasks
Extending the general structure for testing OP, Crosby et al. developed a task where rewards can be occluded in one of up to 12 different positions, through holes in the floor (Crosby et al. 2020). O-PIAAGETS contains 430 instances based on this design, with 4, 8, or 12 holes (see Fig. 11). 191 instances require OP to be solved, with the remaining 239 not requiring OP, where the holes are not deep enough to occlude the goal. These latter instances serve as controls.
Figure 11: Primate Cognition Test Battery Grid (’Grid’) tasks. Top: Top-down view; bottom: Agent-view. Red arrow indicate reward location, grey arrows indicate agent location. Reproduced from Voudouris et al., 2022.
Chiandetti & Vallortigara Chick Tasks
Chiandetti and Vallortigara developed a series of tests for investigating physical common-sense in day-old chicks (Chiandetti and Vallortigara 2011), which Crosby et al. developed for use in the Animal-AI Environment (Crosby et al. 2020). In the example in Fig. 12, the agent observes a reward rolling away (panel A). Then the lights go out, blocking all visual input for several time steps (panel B). During this lights out period, the reward is deflected either left and rolls behind the occluder. When the lights turn on, the reward is fully occluded (panel C). If the agent understands that the reward continues to exist even when it is out of view, they ought to reason that it is behind the occluder on the left, as there is nowhere it could be hiding on the right.
Figure 12: Chiandetti and Vallortigara (’CV’) Task. Top: Reward Rolling away; middle: ”Lights out” period; bottom: Reward occluded. Reproduced from Voudouris et al., 2022
We used 36 instances from O-PIAAGETS that tested OP in this way, and included 1195 further tasks that did not require OP to be solved, to serve as controls.
Basic Tasks
All of these tasks require an agent to be able to navigate towards green goals, around obstacles, while traversing platforms and ramps, and avoiding lava. When testing for OP in biological agents such as primates or chicks, it is often safe to assume that they understand the value of rewards such as food or that they are able to navigate or search for rewarding objects. This is not a safe assumption when evaluating artificial agents, particularly those trained purely with techniques from machine learning. Therefore, the testbed must include instances that test for these basic skills. We included a further 521 instances testing basic abilities to navigate towards reward, avoid lava, and use ramps. These are extensions and variations of the tests used in section  4.
In total, we included 2188 unique instances from O-PIAAGETS, varying along the 12 dimensions outlined in Table  4. Note the asterisks in the table, denoting that a particular feature was not used in the measurement layouts used for evaluation in this paper. This occurred because Team B either didn’t realise the feature was present in the environment or because it was missed as a possible source of bias or capability. Nevertheless, this actually highlights a strength of our evaluation framework; imperfect measurement layouts can still be useful for both inferring agent capabilities and predicting future agent performance.
Dimension
Description
Range/Values
Reward Visibility*
Is the reward visible at the choice point?
${0,1}$
Time Under Occlusion
Time reward is occluded
$[0,4]$
Positions
Number of alternative reward places
$[1,11]$
Reward Distance
Manhattan distance to the goal
$[0,56]$
Right/left Position
Is reward left, right or straight from the agent?
$\{-1,0,1\}$
Reward Size
How large is the reward?
$[0.5,4]$
Occluder Presence
Are there any occluders hiding the rewards?
$\{0,1\}$
Lava Presence
Is there any lava in the arena?
${0,1}$
Platform Presence
Are there any platforms in the arena?
${0,1}$
Ramp Presence
Are there any ramps in the arena?
${0,1}$
Transparent Walls
Are there any transparent walls in the arena?
${0,1}$
Occluder Redness*
What is the R value of occluders in the arena?
$[0,255]$
Occluder Greenness*
What is the G value of occluders in the arena?
$[0,255]$
Occluder Blueness*
What is the B value of occluders in the arena?
$[0,255]$
Table 4: Object Permanence Task Characterisation details. The dimensions that instances vary on. The dimensions with an asterisk are not included in the measurement layout presented in this paper.
This test-bed satisfies our criteria for capturing and evaluating OP that we outlined above:
  1. 1.
    Object permanence is difficult to place on an ordinal scale. In cognitive science, it is usually discussed in terms of its presence or absence.
  2. 2.
    Object permanence relies on working memory capacity and spatial cognition, and successful performance on tests of OP often require good navigational skills. The dimensions defining each task vary on their relevance to OP and these precursor abilities (working memory, spatial cognition, navigational ability). There is therefore a complex interaction between latent capabilities and task dimensions.
  3. 3.
    Instances were not defined on all dimensions, introducing missingness. For example, in many instances, there were no occluders, so Occluder Colour is undefined.
  4. 4.
    Due to the structure of Animal-AI and O-PIAAGETS, some instance-types are represented more than others. For example, we have 36 Chiandetti & Vallortigara Chick Tasks testing OP, compared to 359 Primate Cognition Test Battery Cup Tasks, almost 10 times more.
D.3 Object Permanence Measurement Layout
This section extends the information about the measurement layout in section 5.2 and Fig. 5. The measurement layout for OP consists of seven capability nodes, one bias node, and one robustness node (stochastic noise on success/failure). As in the AAIO case, each of the capability and bias nodes has an associated instance-performance node, an intermediate non-observable node (INON). The noise is applied at the end to the final performance probability.
Name Type Range/Values Prior/Linking Function
rampPresence Meta-feature $\{0,1\}$
lavaPresence Meta-feature $\{0,1\}$
platformPresence Meta-feature $\{0,1\}$
goalDistance Meta-feature $[0,56]$
goalSize Meta-feature $[0.4,5]$
rightleftPosition Meta-feature $\{-1,0,1\}$
occluderPresence Meta-feature $\{0,1\}$
timeUnderOcc Meta-feature $[0,4]$
mumberOfPositions Meta-feature $[1,11]$
OPAbility Capability $[0,48.4]$ $Uniform(0,48.4)$
memoryAbility Capability $[0,4.4]$ $Uniform(0,4.4)$
visualAbility Capability $[0,6]$ $Uniform(0,6)$
rampAbility Capability $[0,1]$ $Beta(1,1)$
lavaAbility Capability $[0,1]$ $Beta(1,1)$
platformAbility Capability $[0,1]$ $Beta(1,1)$
flatNavAbility Capability $[0,56]$ $\mathcal{N}(0,1)$
rightleftBias Bias [,][$\infty$,$\infty$][ - ∞ , ∞ ] $\mathcal{N}(0,1)$
noiseLevel Robustness $[0,1]$ $Uniform(0,1)$
rampPerformance INON $[0,1]$ $\mu({\sf rampAbility},{\sf rampPresence})$
platformPerformance INON $[0,1]$ $\mu({\sf platformAbility},{\sf platformPresence})$
lavaPerformance INON $[0,1]$ $\mu({\sf lavaAbility},{\sf lavaPresence})$
flatnavPerformance INON $[0,1]$ $\sigma({\sf flatNavAbility}-{\sf goalDistance}+{\sf rightLeftEffect})$
navigationPerformance INON $[0,1]$ ${\sf rampPerformance}\times{\sf platformPerformance}\times{\sf lavaPerformance}\times{\sf flatnavPerformance}$
memoryPerformance INON $[0,1]$ $\sigma({\sf memoryAbility}-{\sf timeUnderOcc})$
OPPerformance INON $[0,1]$ ${\sf memoryPerformance}\times\sigma(48.4-((48.4-({\sf OPAbility}-{\sf 4}\times{\sf numPositions}))\times{\sf occluderPresence}))$
visualAcuityPerformance INON $[0,1]$ $\sigma({\sf visualAbility}-({\sf 5}-{\sf goalSize}))$
rightleftEffect INON [,][$\infty$,$\infty$][ - ∞ , ∞ ] ${\sf rightLeftBias}\times{\sf rightLeftPosition}$
taskPerformance Observed $[0,1]$ $p\sim Bernoulli(\omega[{\sf noiseLevel}]({\sf OPPerformance}\times{\sf navigationPerformance}\times{\sf visualAcuityPerformance},\nu))$
Table 5: Nodes in the Object Permanence Measurement Layout, including priors and linking function.
Table 5 gives the complete details, broken down by node type. As in the AAIO case, the function $\sigma$ is the standard logistic function while $\omega$ is a parameterised weighting function, where the value $\nu$ represents a noise prior that represents a constant model set to 1 -- average performance. The function $\mu$ is a margin, defined as
$\mu(a,b)=1-((1-a)\cdot b)$
where $a$ is an ability and $b$ a binary feature. This means that when the binary feature is 0 then the margin represents $p(success)=1$. If the binary feature = 1 then $p(success)=a$. The units of some of the abilities derive from the meta-features, like the memory ability, which is seconds, as it comes from the opposing metafeature timeUnderOcc. Similarly, visualAbility has the same units as goalSize, and flatNavAbility has the same units as goalDistance. OPAbility comes from a product between timeUnderOcc and numPositions so its unit is seconds times positions. Clearly, the most complicated linking function is the one for OPPerformance, which basically is opposed to (subtracted by) the maximum time under occlusion (4) times the number of positions. This is subtracted from the maximum OPAbility (48.4) and multiplied by the occluderPresence which is subtracted again from the maximum OPAbility (48.4). A logistic function is applied as usual but then this is multiplied by memoryPerformance, representing the dependency between object permanence and memory.
D.4 Synthetic Agents and Comparison
The synthetic agents Team A designed can be found in Table  6 and Table  7.
Name Agent Class Agent Description
Agent 1
Reference
Passes every instance.
Agent 2
Reference
Passes each instance with a probability of 0.5.
Agent 3
Reference
Passes each instance with a probability of 0.1.
Agent 4
Reference
Approximates a random walker. Passes a small proportion of instances. It is more likely to pass ’safe’ instances that don’t have lots of dangerous holes to fall down or lava to run into.
Agent 5
Reference
Passes each instance with a probability of 0.9. This is a more human-like failure rate than agent 1.
Agent 6
Achilles Heel
Has OP, passing an instance with a probability of 0.95 unless there is lava present, in which case it passes with a probability of 0.1.
Agent 7
Achilles Heel
Has OP, passing an instance with a probability of 0.95 unless there is lava or an occluder with a ‘redness‘ over 200 present, in which case it passes with a probability of 0.1.
Agent 8
Achilles Heel
Has OP, passing an instance with a probability of 0.95 unless there is an occluder with a ‘greenness‘ over 200 present, in which case it passes with a probability of 0.1.
Agent 9
Achilles Heel
Has OP, passing an instance with a probability of 0.95 unless there is a ramp present, in which case it passes with a probability of 0.1.
Agent 10
Achilles Heel
Has OP but lacks fine control of actions. If there is lava present, it only passes those instances with a probability of 0.5. It often falls into the wrong holes on its way to the further holes, so as distance to goal increases, the probability of passing Grid instances increases from 0.6 in increments of 0.1. This is only for the OP tasks. For the non-OP instances and the remaining instances, it solves an instance with a probability of 0.8.
Agent 11
Achilles Heel
Has OP but poor visual acuity. If goals are 0.5 or smaller, they only pass the instance with a probability of 0.1. For goals of size 1, the probability is 0.3, it is 0.5 for size 1.5, 0.7 for size 2, 0.8 for size 2.5, and 0.9 for a goal size 3 or larger.
Agent 12
Achilles Heel
Has OP but is biased to the right. Probability of passing a task is 0.95 when the goal is to the right, 0.3 when it is to the left, and 0.6 when it is centrally located.
Agent 13
Achilles Heel
Has OP but is biased to the left. Probability of passing a task is 0.95 when the goal is to the left, 0.3 when it is to the right, and 0.6 when it is centrally located.
Agent 14
Achilles Heel
Has OP, but is easily confused when there are several positions that are similar looking (either identical or mirror images of each other).
Agent 15
Fraudster
Does not have OP, but goes to where they last saw the reward. Deterministically passes all instances where this policy works (all tasks except some of CV Chick tasks), and fails the rest.
Agent 16
Fraudster
Does not have OP, but goes to where they last saw the reward. Passes all instances where this policy works with probability of 0.95 (all tasks except some of CV Chick tasks), and passes the rest with probability 0.25.
Table 6: The agents Team A designed and synthesised performances for.
Name Agent Class Agent Description
Agent 17
Achilles Heel
Has OP, but has low memory, so struggles when the goal is occluded for longer.
Agent 18
Achilles Heel
Has OP, but is easily confused when there are more positions that the goal could be occluded in. Passes instances with more than 10 other positions available with probability of 0.3. Between 5 and 10, probability of 0.6. Between 3 and 5, probability of 0.75. Less than 3, passes with probability of 0.95.
Agent 19
Context-Specific
Non-robust OP. It can solve the navigation tasks and all the CV tasks, but fails the other instances.
Agent 20
Context-Specific
Non-robust OP. It solves the navigation tasks and all the CV tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
Agent 21
Context-Specific
Non-robust OP. It can solve the navigation tasks and all the Grid tasks, but fails the other instances.
Agent 22
Context-Specific
Non-robust OP. It can solve the navigation tasks and all the Grid tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
Agent 23
Context-Specific
Non-robust OP. It can solve the navigation tasks and all the 3 cup tasks, but fails the other instances.
Agent 24
Context-Specific
Non-robust OP. It can solve the navigation tasks and all the 3 cup tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
Agent 25
Context-Specific
Non-robust OP. It can solve the navigation tasks and all the CV and grid tasks, but fails the other instances.
Agent 26
Cognitive Pathology
Non-robust OP. It can solve the navigation tasks and all the CV and grid tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
Agent 27
Cognitive Pathology
Non-robust OP. It can solve the navigation tasks and all the 3 Cup and grid tasks, but fails the other instances.
Agent 28
Cognitive Pathology
Non-robust OP. It can solve the navigation tasks and all the 3 cup and grid tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
Agent 29
Cognitive Pathology
Non-robust OP. It can solve the navigation tasks and all the CV and 3 cup tasks, but fails the other instances.
Agent 30
Cognitive Pathology
Non-robust OP. It can solve the navigation tasks and all the CV and 3 cup tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
Table 7: The agents Team A designed and synthesised performances for (continued).
Figs. 13 and 14 show the models inferred values for each element of the cognitive profile. Fig 13 gives the mean of the inferred distribution and Fig .14 the standard deviation. These values were used by Team B to infer the qualitative type of agent being evaluated. Multiple aspects of the values inferred by the model need to be considered to make a proper assessment of each agent—particularly since Team B did not know a priori the exact types of agents that would exist. For example, Agent 2 has relatively low scores for most abilities, a high standard deviation for most abilities, and a value of $0.5$ for the mean and s.d of overall success. This led Team B to believe that this agent was succeeding the tasks at random with 50% probability. Equally, when looking at Agents 13 and 14, the right-left bias was high (with low standard deviation) leading Team B to believe that these agents were massively favouring movement in particular directions.
In Fig. 15 we perform a qualitative comparison of the ground truth for the designed agents and the inferred profiles. The left side of the figure shows the ground truth characteristics of the agents generated by Team A. The right side of the figure shows Team B’s conclusions about the characteristics of the agents based on the outputs of the model. The colours denote how well Team B’s conclusions align with the ground truth, with green being more accurate and red being less accurate.
Using the measurement layout, Team B were able to recover all reference agents accurately. Team B were also able to recover the majority of the ‘Achilles Heel’ agents with high accuracy. Notable failures are the identification of a slight weakness with Lava for Agent 8 and a slight weakness with ramps for Agent 14. Agent 8 had a 90% probability of failing a task containing an occluder with a greenness value over 200, which has no bearing on the presence or absence of lava. It is possible that this small bias towards passing tasks without lava is a spurious result derived from the random values sampled to generate the performances. Agent 14 struggles with similar looking locations for reward locations, of which there are more for the grid tasks and the cup tasks, which both contain ramps by necessity, potentially explaining this ramp bias.
Team B were able to recover the cognitive profiles of Agents 17 and 18 to some degree. Team B inferred that Agent 17 struggled with memory, but because OP ability requires memory in the measurement layout, low memory ability was conflated with slightly lower OP. Agent 18 had worse performance when the reward could be hidden in more places. Similarly, the measurement layout stipulates that OP ability is required to track objects in multiple locations, so poorer performance on tasks with more hiding places was interpreted as a lower OP ability, a plausible inference.
Team B were able to partially recover the cognitive profiles of the context-specific agents. The relatively lower accuracy here was because task type was not included in the measurement layout. Therefore, many of the biases recovered by Team B reflect the particular construction of the paradigms used by Team A (see legend of Fig.  15).
Team B did not accurately recover the cognitive profiles of the ‘fraudster’ agents. Both Agents 15 and 16 follow the policy of going to where the goal was last seen. This results in generally good performance in all tasks except the Chiandetti & Vallortigara tasks, where this policy would result in falling into lava. Therefore, the performance distributions look similar to a high performing object permanence agent. To successfully discriminate these sophisticated agents, we would need to weight the importance of certain sub-groups of instances, such that, for example, Chiandetti & Vallortigara tasks are of most inferential importance in the case where there is good performance on the remaining paradigms. This requires targeted inference and diligent use and application of measurement layouts.
Overall, however, Team B was able to extract a lot of information about a the majority of the synthetic agents—despite an imperfectly designed measurement layout. Also noteworthy is that when the measurement provides correct inferences, they tend to be consistent across that particular agent. That is, if the measurement layout identifies a particular weakness in an agent, it won’t also pick a second, nonexistent weakness also.
Figure 13: Mean inference results for the cognitive profile elements, along with the mean success rate, for the 30 synthetic agents in the OP task.
Figure 14: Standard Deviations of inferred cognitive profile elements, along with the standard deviation of the success rate, for the 30 synthetic agents in the OP task.
Figure 15: Qualitative analysis of the 30 synthetic agents, comparing the characteristics as they were designed with the cognitive profile that we have identified with our model decision for the measurement layout. In a gradient from green-yellow-red, with green representing a successful identification and red an unsuccessful one.
Appendix E Prediction Results
This section gives the full results for prediction on unseen instances. The synthetic data set was partitioned into an 80/20 train-test split. The measurement layout models (one for each agent) were trained on the same training partition. For each instance from the test-set we apply forward inference to yield a success probability.
We evaluate these probabilities using a Brier score (Brier 1950), also making use of the decomposition into Calibration and Refinement (Murphy 1973). These scores are given in Table 8, with the better performing approach for prediction in bold for each agent. As expected, the aggregate predictions have no refinement at all, and the values are close to 0.25 for balanced cases (the lower numbers only appear when the proportion of right/wrong performance is more imbalanced). The results from the measurement layout are much better here, as this method can give refined results per instance. When we look at calibration, however, we see aggregate predictions are almost perfect, since train and test come from the same distribution. But the calibration from the measurement layout is also quite good, and overall there are many more wins for the measurement layouts approach. Moreover, it is important to emphasise that unlike this table, train and test do not have to come from the same distribution. For instance, if we removed the easy instances for many of the tasks, we would get similar results for the measurement layout prediction approach (calibration and refinement should not change dramatically), but the results for the aggregate prediction approach would be much worse, since the calibration would be completely lost.
Looking at the particular agent properties that cause the measurement layouts to be more or less capable at prediction can provide some insight into the strengths and weaknesses of our approach. Agent 1 was designated as the perfect agent that will always pass every instance. It is no surprise then that predictions using the aggregate success rate were very successful, achieving the best possible Brier score. Our model’s predictions were still highly accurate (Brier score of 0.022), but the model lacked the extreme certainty of the aggregate measure—likely because using Bayes’ rule, beginning with an uninformed prior it’s impossible to update the posterior to exclude one of the outcomes.
Agent 5 is another case where the aggregate is better at prediction than our measurement layouts. Agent 5 passed all instances with probability 0.9. We believe our model struggled to perform as well as the aggregate here due to there not being sufficiently many instances to move the initial prior (uniform distribution) to the more extreme distribution predicting success at a rate of 0.9. Our model was able to cope with other fixed pass rates (such as with Agent 2, 3, and 4), so this may just be a matter of degree. It’s worth noting that for any fixed pass rate Agent, it’s actually impossible to predict more accurately than the aggregate in expectation. Hence For these types of agents the Aggregate Brier score is minimised, and for many fixed pass rate agents, our measurement layouts matched the aggregate’s Brier score.
Agents 7 and 8 were predicted equally or better by the Aggregate measure than our model. These agents passed all instances with probability 0.95, unless the object was occluded by an object with a lot of red or green present in its RGB colour value (red for 7, green for 8). Given that the measurement layout did not include colour values, it’s easy to see why they would struggle to predict this complex behaviour accurately. On the other hand, they did not do much worse than the aggregate prediction, with Brier scores of 0.138 vs 0.138 and 0.109 vs 0.103 respectively.
The only other agents that our measurement layouts were worse at predicting than the aggregate were agents 15 and 16. These two agents were designed as ‘Fraudster’ agents that mimic object permanence by going to the last position where they saw the reward. This is sufficient to solve many object permanence tasks except for some of the CV chick tasks. When this ‘Fraudster’ policy would be successful, these agents succeed with a probability of 0.95, if the policy would not be applicable, the agents succeeds with probability 00 (for agent 15) or $0.25$ (for agent 16). We believe it is the high overall success rate that causes our model to be less predictive here, similar to Agent 5 (and to a lesser extent Agent 7 and 8), again because it takes more task instances to move an uninformative policy to extremes such as predicting success with a probability greater than $0.9$.
Agent ML BS Agg. BS ML Cal. Agg. Cal. ML Ref. Agg. Ref.
1 0.022 0.000 0.022 0.000 0.000 0.000
2 0.250 0.250 0.000 0.000 0.250 0.250
3 0.078 0.078 0.001 0.001 0.077 0.077
4 0.022 0.022 0.000 0.000 0.022 0.022
5 0.099 0.089 0.011 0.000 0.087 0.089
6 0.067 0.250 0.011 0.000 0.056 0.250
7 0.138 0.138 0.012 0.000 0.125 0.137
8 0.109 0.103 0.008 0.000 0.101 0.103
9 0.060 0.240 0.005 0.000 0.055 0.239
10 0.211 0.237 0.002 0.000 0.209 0.236
11 0.176 0.242 0.012 0.000 0.164 0.242
12 0.218 0.235 0.003 0.000 0.215 0.235
13 0.222 0.234 0.004 0.001 0.218 0.234
14 0.152 0.166 0.008 0.000 0.145 0.166
15 0.026 0.007 0.020 0.000 0.007 0.007
16 0.077 0.064 0.012 0.000 0.063 0.064
17 0.090 0.166 0.022 0.000 0.068 0.166
18 0.069 0.072 0.008 0.000 0.061 0.072
19 0.056 0.249 0.004 0.000 0.051 0.249
20 0.177 0.246 0.016 0.000 0.162 0.246
21 0.156 0.249 0.023 0.001 0.132 0.248
22 0.204 0.248 0.013 0.004 0.191 0.245
23 0.012 0.250 0.011 0.003 0.000 0.247
24 0.165 0.245 0.005 0.000 0.159 0.245
25 0.149 0.204 0.023 0.000 0.125 0.204
26 0.229 0.248 0.012 0.001 0.217 0.247
27 0.084 0.192 0.034 0.002 0.048 0.190
28 0.186 0.242 0.028 0.003 0.158 0.239
29 0.098 0.152 0.056 0.000 0.040 0.152
30 0.205 0.232 0.013 0.000 0.191 0.232
Table 8: Brier Scores (BS) and the breakdown into calibration (Cal.) and refinement (Ref.) for the Measurement Layouts (ML) and aggregate predictions (Agg.)
Appendix F Methodology for Designing Measurement Layouts
For complex, multifaceted tasks measurement layouts have the potential to grow and become large and complex themselves. It’s therefore important that to apply them to other domains (and reap the benefits of the additional evaluation information they provide) to have a general methodology for constructing the appropriate measurement layout structure. To this end, we developed a methodology for constructing measurement layouts that we wish to explicate here more clearly.
As stated in the main paper, the core of our approach is to identify the demands of a task, and to compare these against the capabilities of the system being evaluated. Tasks are composed of multiple instances that vary in the degree of demands. In our methodology we tie demands with meta-features in order to give the demands values that are tied to objective properties of the task. For each demand in a task, a system will have varying levels of performance at dealing with the demand based on its capability. we model this as a probability of the system successfully handling the aspect of the task with its meta-feature value. This is where the linking function comes into play. We often leveraged the sigmoid function $\sigma(capabiltiy-demand)$. This approach is beneficial because it ties capabilities to probabilistic performance based on the demand. This also causes the capabilities to have the same unit as the demand, and allows us to interpret a capability with value $x$ as consistently succeeding on the demand with meta-feature $x$ in $50\%$ of instances.
Certain demands also require that other demands have been met. This requires a dependency understanding of the demands of the task. Performances for dependent task demands can be incorporated into the sigmoid margin calculation or can have an effect on the resulting performance value. The exact approach for combining demand performances will depend on the task itself, and whether the demands are compensatory—whether performance can be achieved with one capability compensating for lacking another. Our environments were non-compensatory, so we often opted for taking the product of probabilities. However, more complex relations such as taking the minimum or maximum of a set of probabilities is available by utilising a generalised mean. The other major aspect that we need to identify to construct the measurement layout is the sources of possible bias. In principle, this is similar to creating the structure for capabilities: We still identify meta-features that may affect performance, however, now we are looking for those that are not necessary for the task. We still include a bias ”performance” for each instance (how present was the bias in this instance, and in which direction was it directed?) , but this is now included in the sigmoid function for the relevant capability it may affect: $\sigma(capability-demand+bias)$. Once all of the demands and biases are accounted for, a final combination can give an output for overall task performance.
Breaking a complex task down into demands and possible biases is not always straightforward. Care, and domain knowledge of what constitutes success in the task, needs to be utilised to identify the core broad capabilities required, and how these break down to identifiable meta-features. We recommend an iterative approach: If a particular measurement layout cannot explain system behaviour reliably (with a low standard deviation on the resulting posterior distributions), then there is either too few evaluation instances, or the measurement layout is missing an important capability or bias.
When evaluating Artificial Intelligence systems, and for that matter non-human Biological Intelligences, we need to be careful to avoid anthropomorphising the system. We do not want to mistakenly ascribe capabilities or explain behaviours based on a crude facsimile of human behaviour. AI systems, in particular, have the potential to solve tasks using incredibly different internal methods and skills. We have been careful to avoid this mistake in our work by focusing on the required demands for a task rather than any model for how the system internally processes meta-features.
Appendix G Computation Code, and Datasets
Our experiments did not require a large amount of compute, We ran all of our experiments on Google Colaboratory (Google 2017) using the free tier of service without utilising the GPU. All corresponding code, datasets and raw results will be released on publication to ensure ease of replication.

### v2

@@
 Abstract
-As machine learning models become more general, we need to characterise them in richer, more meaningful ways.
+As machine learning models become more general, we need to characterise their capabilities in richer, more interpretable ways that move beyond aggregated statistics on static benchmarks.
 We describe a method to infer the cognitive profile of a system from diverse experimental data.
-To do so, we introduce measurement layouts that model how task-instance features interact with system capabilities to affect performance.
-These features must be triangulated in complex ways to be able to infer capabilities from non-populational data – a challenge for traditional psychometric and inferential tools.
-Using the Bayesian probabilistic programming library PyMC, we infer different cognitive profiles for agents in two scenarios: 68 actual contestants in the AnimalAI Olympics and 30 synthetic agents for O-PIAAGETS, an object permanence battery.
-We showcase the potential for capability-oriented evaluation.
+To do so, we introduce measurement layouts which model how task-instance features interact with system capabilities to explain performance.
+System capabilities can be estimated using Bayesian triangulation, inferring their value based on the performance of a model on tasks with different features.
+Our approach accurately recovers the cognitive profiles of hand-crafted behavioural agents, as well as estimating the cognitive profiles for deep reinforcement learning agents and human children in a virtual game environment.
+These cognitive profiles are significantly richer than aggregated benchmark statistics, summarising multiple distinct capabilities that explain behaviour, and are also significantly more predictive, accurately estimating performance on new, held-out tasks.
 1 Introduction
 What does success or failure tell us about a system’s capabilities?
-In isolation, performance results are hard to interpret: most interesting tasks require several different abilities to be successfully combined.
-instances.
-For example, an agent exploring a 3D environment to find a reward needs to understand that the reward still exists when occluded (object permanence).
+In isolation, performance results are hard to interpret: most interesting tasks require several different capabilities to be successfully combined.
+For example, consider an agent exploring a 3D environment similar to that in Figure 1(a), where in order for the agent to find the reward, it needs to understand that the reward still exists when occluded (known as object permanence).
 It must then remember where it is, and successfully navigate to it.
-If the agent fails, we cannot know the cause: lack of object permanence, limited memory or navigation skills.
-This leads to poor predictability about the performance of the agent for new tasks, in or out of the training distribution.
-Figure 1: This task instance requires the agent to: (a) understand that the reward still exists when occluded (object permanence), (b) remember which of the three walls it is occluded behind while it navigates towards it (memory), and (c) reach it over ramps and around the wall (navigation).
-Figure 2: Triangulation in the measurement layouts using results from task instances like this one.
-Bottom-up inference from three tasks (in green and red for success and failure respectively) leading to the cognitive profile.
-Top-down inference (in blue) predicting failure for the fourth task.
-However, suppose we also knew that the agent performs well on tasks that involve complex navigation, and the same for tasks involving memory.
-We could now infer that the system’s failure is likely a failure of object permanence by triangulation, a robust inferential method in formal epistemology (Heesen, Bright, and Zucker 2019) and causal (Bayesian) reasoning.
+If the agent fails on an instance of this task, the cause is unclear—is it a failure of object permanence, limited memory, or navigation skills?
+Without an answer to this question, it is difficult to predict performance for new tasks with different features.
+One way to approach this issue is to collect the performance of the agent on a benchmark of arbitrarily varied tasks (with and without occlusion, shorter or longer distances to the goal, etc.), and compute the proportion of successes across that set.
+This aggregated statistic approximates the expected probability of success on a new task instance, but it does so blind to its specific features.
+Moreover, this approach fails to explain why the agent passes or fails on a certain task.
+(a) A three-cup instance. The agent must (a) retain the object’s existence when occluded, (b) remember which wall hides it, and (c) navigate around obstacles.
+(b) Bayesian triangulation: bottom-up inference (green/red) yields a cognitive profile; top-down inference (blue) predicts performance on a new task.
+Figure 1: (1(a)) shows a representative task instance; (1(b)) shows how Measurement Layouts infer and predict capabilities from instance-level results.
+Suppose, however, we also knew that the agent performs well on tasks that involve complex navigation, and the same for tasks involving memory.
+We could now infer, by triangulation, that the system’s failure is likely a failure of object permanence, a robust inferential method in formal epistemology (Heesen et al., 2019) and causal (Bayesian) reasoning.
 This bottom-up inference is illustrated in green and red in Fig.
-2.
+1(b).
 We can use this information to draw conclusions about where the system is safe to deploy, and to direct efforts to improve it.
 We can also predict behaviour in new scenarios: for example, the agent will likely fail at a new task with high object permanence demands, as illustrated in blue in Fig.
-2.
+1(b).
 This example demonstrates that we can leverage the fact that different tasks—and their constituent instances—have different demands to help us extract conclusions about system capabilities from patterns of task performance.
@@
 If we were to examine each task or instance in isolation or to simply aggregate performance across all tasks, this triangulation of abilities would not be possible.
-To address this, we propose a novel Bayesian approach to evaluation that enables simultaneous inferences about multiple abilities from patterns of task performance.
-This approach has a series of principles facilitating a capability-oriented approach to the evaluation of machine learning models:
-1.
-We do not simply break out performance by ‘categories’.
-Related work employs a ‘taxonomy of tasks’ (see e.g., Table 1 in (Liang et al.
-2022)), each placed into one or more categories.
-This limits the inferential capacity and predictability of these category aggregates for new tasks.
-Instead, we evaluate based on instance demands.
-2.
-Unlike psychometric approaches such as factor analysis (FA), structural equation modelling (SEM) or item response theory (IRT), we do not rely on populational data (Gustafsson and Undheim 1996).
-Instead, we infer the cognitive profile of a single subject from their performance data alone.
-3. We do not ‘reify’ extracted latent variables as capabilities, which has been done hierarchically elsewhere, (e.g., the Cattell-Horn-Carroll model (Keith and Reynolds 2010)).
-Instead, we identify general linking functions that help derive the requisite capabilities from the demands of the tasks.
-4.
-Probabilistic models currently used in cognitive modelling (Lee 2011) are often simple at the level of the individual and cannot convey the required probabilistic expressions connecting individual performance, capabilities and task features.
-Instead, complex interactions allow for Bayesian triangulation which we use for informative evaluation.
+To address this, we propose a novel Bayesian approach to evaluation that enables simultaneous inferences about multiple capabilities from patterns of task performance.
+This approach is underpinned by several psychometric principles that we elucidate in Appendix A.
 In this paper we find an operational solution to these inferential challenges in the form of Measurement Layouts: specialised and semantically-rich hierarchical Bayesian networks.
-Using the inference power of the No U-Turn Sampler (Hoffman and Gelman 2014) in PyMC (Salvatier, Wiecki, and Fonnesbeck 2016), a probabilistic programming engine, we can triangulate from performance and the task’s cognitive demands to the capability and bias levels for each subject (their cognitive profile).
+Using the inference power of the No U-Turn Sampler (Hoffman and Gelman, 2014) in PyMC (Salvatier et al., 2016a), a probabilistic programming engine, we can triangulate from performance and the task’s demands to the capability and bias levels for each subject (their cognitive profile).
 Then, performance can be inferred downstream for new tasks.
-We derive the cognitive profiles in two scenarios: the contestants in the AnimalAI Olympics (Crosby et al.
-2020) and several synthetic agents for O-PIAAGETS (Voudouris et al.
-2022), an object permanence battery.
-We show how systems with similar overall performance can have very different capability levels and biases.
-2 Background
-AI systems are often evaluated on one or more standard benchmarks.
-Because of the remarkably broad capabilities of modern systems, benchmarks are becoming larger and more diverse.
-For instance, BIG-bench (Srivastava et al.
-2022) contains over 200 tasks for evaluating language models.
-Aggregating many tasks into a single score does not allow for nuanced conclusions about where or why a system fails (Burnell et al.
-2023), and has little value for predicting system performance a specific unseen task.
-Analysing 200 tasks separately would lead to an unwieldy and hard to interpret set of results.
-One option is the categorisation of tasks across several dimensions, providing a moderate level of insight into where a model succeeds or fails (Osband et al.
-2019; Ilievski et al.
-2021; Crosby et al.
-2020; Liang et al.
-2022).
-However, conclusions are highly dependent on how instances are assigned to categories, and predictive value is limited inside and across categories.
-An widely-used approach that does not require any taxonomy of capabilities is to extract and reify them from the data.
-Given experimental data for many tasks and subjects, we can use Factor Analysis (FA) to infer those factors that best explain the observed variance in the population.
-Some structure can be inferred from FA, as in the classical hierarchical models (Cattell-Horn-Carroll model (Keith and Reynolds 2010)), but today it is common to use more powerful techniques such as Structural Equation Modelling (SEM) (Ullman and Bentler 2012).
-However, SEM assumes linear relationships between variables and that dependent variables have normally distributed sample means.
-These assumptions are unlikely in the real world.
-Further, SEM is a populational technique requiring a large sample of individuals in order to confirm structural relationships.
-A more granular approach is to identify the difficulty of each task instance and the ability of each system based on the relationship between task difficulty and the system’s performance.
-This idea is the cornerstone of Item Response Theory (IRT) (Embretson and Reise 2000; De Ayala 2009).
-When several abilities exist, we must rely on multidimensional IRT(Reckase 2009) However, IRT is also populational, and derives abilities and difficulties relative to population averages, therefore the parameters of the same item can change if we add new items, or the abilities of a system may change if we add more systems.
-This poses a problem for AI evaluation, where the population of individual systems is constantly changing and advancing.
-Still, with FA, SEM, IRT or other techniques, patterns of performance are mapped to capabilities, making up a multi-dimensional characterisation of the subject that is sometimes called a cognitive profile (Letteri 1980).
-Cognitive profiles make it possible to identify differences that might go unnoticed in aggregated scores, make predictions about future performance, and gain insights into the relationships between different abilities (e.g., deficits in mental rotation are one cause of poor spatial navigation).
-Most psychological assessment is populational.
-Exceptions can mostly be found in developmental psychology and animal cognition (Moran et al.
-2017; Bergman and Andersson 2015; Shaw and Schmelz 2017).
-The scarcity of subjects and the difficulty of experiments results in an emphasis on carefully designed tasks that ensure a systematic coverage of the relevant features and possible distractors.
-On the other hand, if the cognitive mechanisms for solving the task are accurate and enough variation is introduced, as in the example of Fig.
-2, the results can be sufficient to estimate the capabilities of a single individual, independently of the other evaluated individuals.
-This is the type of approach that we need to evaluate ML models and AI systems.
-Fortunately, there is recent movement towards producing training and test variations, augmented examples or corruptions to test systems, through adversarial learning (Zellers et al.
-2018; Nie et al.
-2019; Rozen et al.
-2019; Kiela et al.
-2021), procedural content generation (Risi and Togelius 2020) and other domain-specific techniques (Sugawara, Stenetorp, and Aizawa 2020).
-These variations have also been inspired by software testing, such as (Marijan and Gotlieb 2020) and (Ribeiro et al.
-2020).
-Some of this effort results in annotated datasets, especially for reinforcement learning — from the inspiration behind the AnimalAI Olympics (Crosby et al.
-2020) to the fully-systematic O-PIAAGETS (Voudouris et al.
-2022).
-However, the next step of using these annotated data for capability-oriented evaluation has not been taken yet.
-Not using populational data has to be compensated by a larger number of varied instances.
-When many dimensions of variation are introduced, the inference problem becomes complex.
-Traditional Bayesian Networks (BN) (Ben-Gal 2008) can be used for arbitrarily sophisticated probability distributions.
-For complex networks, though, performing the inference calculation is computationally intractable (Cooper 1990).
-Approximation techniques for Bayesian inference (BI) have been introduced, including the Markov Chain Monte Carlo family (Andrieu et al.
-2003), allowing for the inference of latent values in more complex BNs, particularly those with multiple levels of dependency and complex relations between variables.
-Taking advantage of these recent advances in Bayesian approximation, evaluation procedures, and computing power allows us to extend the benefits of capability-oriented evaluation to machine learning for the first time, enabling more robust, thorough, and principled evaluations of ML systems.
+Not only are cognitive profiles much more explanatory of what causes failure or success, but we find that they are also much more predictive than the traditional approach of using aggregated benchmark statistics to predict performance.
+Our approach is agnostic to the task or type of system being evaluated, so long as there are instance-level task demands that can be related to distinct capabilities.
+To showcase our method, we construct measurement layouts and cognitive profiles for agentic systems in two scenarios: a simple navigation task and a more complex object permanence task.
+All code and data associated with our method can be found on GitHub here
 2 Measurement Layouts
@@
 Cognitive profiles need to include all of these.
-Let us define more precisely a task characterisation and the corresponding cognitive profile.
-We will then detail how they are connected by the measurement layouts.
-Definition 3.1.
-A task characterisation is a set of observable, usually constructed, meta-features $X$, expressing cognitive demands and other high-level properties of the task.
-For instance, in an image classification problem, the level of blur of the image or the level of clutter of the background could constitute perceptual demands.
+We define a task characterisation as the set of observable, usually constructed, meta-features $X$, expressing cognitive demands and other high-level properties of the task.
+For example, in an image classification problem, the level of blur of the image or the level of clutter of the background could constitute perceptual demands.
 Many other properties may also, in practice, affect performance.
-Definition 3.2.
-The cognitive profile of a system is a tuple $\langle C,B,R\rangle$. Where $C$, $B$, and $R$ are vectors of capability levels, bias, and robustness values respectively.
-Capability levels represent what the system can do, bias values represent some other preferences or limitations that may impact performance in a less monotonic way, and robustness levels account for reliability issues, unexplained or random effects (noise) on either the agent or environment.
-Definition 3.3.
-A measurement layout is a directed acyclic graph that connects, through linking functions, the meta-features of a task characterisation with the cognitive profile of a system, in order to predict observed performance.
-We define measurement layouts as Hierarchical Bayesian Networks (HBN) (Murphy 2023) with the following characteristics: Within the measurement layout we include a node for every meta-feature in every task instance and a node for each element of the system’s cognitive profile.
+Similarly, we characterise the cognitive profile of an AI system as a tuple $\langle C,B,R\rangle$.
+Where $C$, $B$, and $R$ are vectors of capability levels, bias, and robustness values.
+Capability levels represent what the system can do, bias values represent some other preferences or limitations that may impact performance in a less monotonic way, and robustness levels account for reliability issues, unexplained or random effects (noise).
+We define a measurement layout as a directed acyclic graph that connects, through linking functions, the meta-features of a task characterisation with the cognitive profile of a system, in order to predict observed performance.
+We implement measurement layouts as Hierarchical Bayesian Networks (HBN) (Murphy, 2023) with the following characteristics: Within the measurement layout we include a node for every meta-feature in every task instance and a node for each element of the system’s cognitive profile.
 All of them are roots of the graph.
@@
 Linking functions are used to formally define relationships between dependent nodes.
-All nodes within the HBN define a probability distribution. We can classify nodes into three categories depending on the provenance of their summary statistics:
-Metafeatures: These come from the task characterisation and are fixed, observable values. They are represented as a fixed Dirac-delta distribution.
-Cognitive Profile Nodes: These are elements of the cognitive profile relevant to assessed task’s demands.
-This includes required capabilities and likely biasses.
-These often combine with meta-features and feed in to derived nodes.
-The range and distribution of these elements derives from the task characteristics.
-To make our inferences as data-driven as possible these begin with uninformative priors.
-Derived Nodes: These are instance-level inferences.
-They are combinations of meta-features and/or cognitive profile elements that have their parameters derived from dependent nodes.
-Information to compute these values is encoded in the linking function for the node, the output of which is used as the summary statistics of the node’s distribution.
-One derived node is the observable performance for the task—the only leaf node—but most are Intermediate Non-Observable Nodes (INON), which represent intermediate performances or effects.
+Nodes can encode meta-features, latent capabilities, and derived nodes such as observable performance or Intermediate Non-Observable Nodes (INON) for representing intermediate marginal performances.
 All three node types were seen in Fig.
-2 (right).
+1(b).
 The connective structure can be enriched by domain knowledge determining which cognitive profile elements and meta-features are expected to affect the values of instance-level inference nodes.
-These relationships can be higher-order and complex, with certain INON nodes requiring the nuanced confluence of many cognitive profile elements and meta-features.
-Complex connective structure (and the requirement for increased levels of domain knowledge) can arise from the need to accurately distinguish between success, failure, and the multitude of confounding behaviour that a carefully designed benchmark must account for.
-By a linking function, we refer to any mathematical expression that maps values from the outputs of one node’s probability distribution to the input of another.
-For instance, it could be a sigmoid function of the difference between capabilities and demands.
-It could also be a product of (non-compensatory) capabilities, representing a quantitative and differentiable expression for a logical ‘and’.
-Alternatively, it could be some additive linear equation, representing a quantitative expression of a logical ‘or’, when some preceding nodes are compensatory—or a generalised mean in between.
+These relationships can be higher-order and complex, with certain INONs requiring complex combinations of many cognitive profile elements and meta-features.
+The need for complexity can arise from the many confounding behaviours that a well-designed benchmark needs to account for.
+By a linking function, we refer to any differentiable function that maps the values from the outputs of one (or multiple) node’s probability distribution to the input of another.
+For example, it could be a sigmoid function of the difference between capabilities and demands.
 The functions can scale the incoming nodes, so that we adjust the ranges for capabilities, biases and noise.
-The possibilities are endless; in practice we use only a few kinds of linking functions.
-We give the specific equations for the linking functions that we used used in this work in Tables 3 and 5 in the appendix.
-These are used in the two scenarios that follow.
-A key conceptual difference between our measurement layouts and typical use cases for HBNs is that our approach is trying to capture a hierarchical dependency relation on capabilities and demands, rather than encoding information on hyper-priors.
+We give the specific equations for the linking functions (and node formulations) that we used in this work in Tables 11 and 3 in the appendix.
+A key conceptual difference between our measurement layouts and typical uses for HBNs is that our approach is trying to capture a hierarchical dependency relation on capabilities and demands, rather than encoding information on hyper-priors.
 The linking functions between nodes are often non-linear, conceived from domain knowledge of the task and applying ideas from cognitive modelling.
-For instance, one of the key decisions is to determine when some capabilities are independent, and if so, whether they are compensatory or not, being realised with additive or multiplicative expressions respectively.
-If they are not independent, then an arrow has to connect them with a linking function expressing this dependency mathematically.
-Given the cognitive profile of a system, and the characterisation of a new task, we can use the measurement layout to anticipate performance by forward (top-down) inference.
-But before that we need to infer the cognitive profile from the observed performance of other tasks by backward (bottom-up) Bayesian inference.
-For this, we utilise PyMC’s inference engine and the No U-Turn Sampler (Hoffman and Gelman 2014) approach to BI (see Appendix A for a detailed description).
-For this to accurately predict a system’s cognitive profile, the test battery
-requires instances to jointly control for alternative explanations permitting performance results to ‘triangulate’ latent capabilities.
-4 Cognitive Profiles for the AAI Olympics
-To illustrate how a measurement layout identifies cognitive profiles, we first select a domain with simple task characterisations that can be straightforwardly related to cognitive profiles to predict performance.
-The Animal-AI Olympics (AAIO) (Crosby et al.
-2020) meets all these requirements.
-The AAIO was an open competition evaluating AI agents in a 3D environment, where agents had to get a reward by locating it and navigating in an arena.
-Participants were asked to build systems (usually reinforcement learning agents) and train them on their own custom-built tasks in the AAI Environment.
-Screenshots for two tasks are shown in Fig.
-7 located in Appendix B along with more information about AAIO and the environment.
-We select 69 tasks from the competition test set that evaluate simple goal-directed behaviour.
-We include 68 participant agents for which we know nothing about their design.
-4.1 Measurement Layout
-In this domain, following the analysis performed in (Burnell et al.
-2022), the set of meta-features $X$ of the instances that characterise task instances are rewardSize, rewardDistance, Xpos and rewardBehind.
-These represent the reward’s size, distance, and if it is left, right, in front, or behind the agent.
-The cognitive profile is composed of two capabilities, navigationAbility and visualAbility, respectively representing the skill to move around purposefully and the visual acuity of the agent, one bias rightLeftBias, representing any preference for left or right movements, and one robustness level noiseLevel, accounting for any unexplained variability.
-We then compare the profile of the agent to the demands of the task to derive performance.
-This is our measurement layout, as can be seen in Fig.
-4 (left).
-The metafeatures $X$ appear as observed (shaded) nodes in the model, appearing within the plate, with 69 instances.
-The elements of the cognitive profile appear as external latent (unshaded) variables.
-The derived nodes lie at the intersection of the meta-features and the external cognitive profile nodes.
-Figure 3: Measurement layout for the Animal AI Olympics.
-Figure 4: Cognitive profiles of the 68 agents in AAIO, with names shown for some of them.
-The $x$-axis and $y$-axis show the inferred means for navigationAbility and visualAbility respectively, with their standard deviations as error bars in grey.
-The radius of each point represents the average performance, while the colour represents the noiseLevel, with red representing higher noise than blue.
-Not visible in the figure are the linking functions associated with each arrow.
-The variable rightLeftEffect is calculated as the product between rightLeftBias and XPos, since left/right bias and XPos are represented with negative and positive values respectively, so the product gives a positive result when both go in the same direction.
-This rightLeftEffect, jointly with rewardBehind and rewardDistance affect navigationPerformance, a non-observable intermediate performance variable.
-The exact formulation of how these three variables are integrated can be found in Appendix C.2,
-but broadly the linking function subtracts the effects of reward distance and position relative to the agent’s viewpoint (e.g., behind) from the navigation ability and adds the right-left bias to this.
-The linking function then applies a logistic function to this margin to give a value between 0 and 1 for navigationPerformance.
-The process to derive visualPerformance is similar, simply subtracting rewardSize ( inverted so the higher the smaller) from visualAbility.
-Note that the ranges of these two abilities are given by the ranges of the task metafeatures, with navigationAbility in $[0,5.3]$ and with visualAbility in $[0,1.9]$ respectively, representing how distant and how big a reward—with their corresponding distance and size units—the agent can obtain.
-Finally, taskPerformance is Bernoulli-distributed on a product of navigationPerformance and visualPerformance adding the noiseLevel.
-We use the product because navigation and visual acuity are non-compensatory (i.e., navigation skills are of no use if the agent cannot see, and vice versa).
-4.2 Inference and Analysis
-Using the measurement layout and the 69 instances, we can use PyMC to infer a very simple cognitive profile for each of the 68 entrants of the competition.
-We run the NUTS inference algorithm with two sampling chains of size 1000 each, finding no inference divergences.
-Fig.
-4 (right) summarises the inferred capabilities and noise levels.
-We see the two main capabilities are correlated for this population, going from low performance in the bottom left area (e.g., ‘ice-play’ with 4% accuracy) to high performance in the top right area (e.g., ‘Trrrr’ with 99% accuracy).
-In these extremes, capabilities are nonexistent (close to 0) or saturated (to the maximum 5.3 and 1.9 respectively), as expected.
-But in the middle range, there are agents that have noticeable differences between their capabilities (e.g., ‘cocel’).
-Some have strong biases, such as ‘Thursday’ and ‘KoozytHiperdyne’ with left-right bias values of $1.14$ and $-0.78$ respectively (not shown on the plot).
-This can have an effect on navigation ability but not on visual ability.
-The standard deviations of the abilities, shown as error bars, and the level of noise are useful to explain situations both when behaviour is conformant or not with the task characteristics.
-For instance, ‘Juohmaru’, with 54% accuracy, is a very conformant agent: performance decreases as rewards get smaller and further away, and when they are initially located behind from the agent.
-We see that navigation ability is $2.50\pm 1.08$ and visual ability is $0.92\pm 0.44$, with a medium level of noise $0.26\pm 0.17$.
-But ‘sparklemotion’, with 36% accuracy, is very non-conformant, with abilities $0.40\pm 0.38$ and $0.59\pm 0.45$, and higher noise $0.31\pm 0.13$.
-Appendix C contains full characteristic grids of these two agents in Fig.
-8 (demonstrating the conformant / non-conformant behaviour) and all inferred results in Fig.
-9.
-Note that we have inferred two different capabilities for each agent as well reliability measure of these estimates: how well they can predict performance (depending on the quality of the estimates, biases and remaining noise) for each particular instance given its meta-features.
-We do this without having different sub-batteries for each capability.
-Our model can also predict performance on unseen tasks instances.
-To evaluate this, we randomly held out 20% of the task instances from the backwards inference stage and evaluated the predicted probabilities of success on them.
-Due to the small size of the dataset, we repeated this process 15 times and took the mean Brier score for each agent.
-Fig.
-5(a) summarises these scores, which are lower than predictions based on aggregate performance for agents with success rates in the range $[0.2,0.7]$, exactly where there is variability to explain.
-Full results and discussion in Appendix E.
-5 Cognitive Profiles for Object Permanence
-Our analysis of the AAI Olympics used a relatively simple scenario as a proof of concept.
-But one limitation of this analysis is that we do not know the ”ground truth” of the agents’ abilities, so it is difficult to determine precisely how effective our approach was.
-It also remains unclear how well our approach works in complex scenarios with many abilities and metafeatures.
-Therefore, we next sought to demonstrate the scalability of our approach by analysing data from a more complex scenario—a rich test battery for evaluating Object Permanence.
-To ensure that we knew the ground truth abilities of the agents, we created a synthetic dataset by simulating the performance of 30 hypothetical agents based on specified cognitive profiles 111because of the complexity of the abilities that affect performance on this testbed, it would be extremely difficult to create a varied population of real agents with specific levels of abilities.
-This scenario tests whether our methods can recover complex cognitive profiles from performance data of agents for which we know the ground truth.
-To do so, the authors of this paper were divided into two teams, Team A and Team B.
-Team A generated the synthetic dataset, while Team B defined the measurement layout and attempted to recover the cognitive profiles of the synthetic agents.
-Both teams were familiar with O-PIAAGETS and the general abilities, bias and elements that are combined in the battery.
-This two team approach allowed us to generate synthetic agents while keeping Team B blind to their profiles, removing the potential for bias and making the modelling effort more akin to a real-world scenario.
-5.1 The Testbed and Synthetic Agents
-We applied our measurement methodology to the Object-Permanence In Animal-AI: Generalisable Test Suites (O-PIAAGETS, (Voudouris et al.
-2022)).
-Object Permanence (OP) is the understanding that objects continue to exist when they go out of view.
-Behaviourally, an agent has object permanence if they behave as though objects continue to exist even when they cannot directly perceive them.
-It is a key component of physical common-sense (Shanahan et al.
-2020; Lake et al.
-2017; Baillargeon et al.
-2011), allowing biological agents to more successfully predict and interact with their environment.
-Object permanence is well studied in cognitive science, where several experimental designs are used to detect its presence in human and non-human animals (Scholl 2007; Flombaum and Scholl 2006; Hare and Tomasello 2005; Chiandetti and Vallortigara 2011).
-Many of these paradigms are represented in O-PIAAGETS, which contains over 13,000 instances.
-We selected three paradigms that test for allocentric object permanence, where objects are occluded independently of the observer’s actions.
-We also include basic control tasks within the test-suite to evaluate the agent’s non-OP capabilities and make the inference process more robust.
-Details about the test-bed design criteria, the included tasks, and why object permanence is difficult to identify are given in the appendix in D.1 and D.2.
+One of the key decisions is to determine how capabilities interact and the extent to which they are compensatory, being realised with additive or multiplicative expressions.
+We use PyMC (Salvatier et al., 2016b) to implement and fit measurement layouts in practice using the No U-Turn Sampler (Hoffman and Gelman, 2014), although any probabilistic programming language can be used for this.
+3 Experiment 1: A Simple Navigation Task
+Figure 2: A: The average final score of each agent, varying pixel input size and navigational noise.
+B: A measurement layout relating two task demands and a capability to performance.
+C: The Brier Scores on the held-out test set of performances, as predicted by the average performance on the train set (Aggregate; Orange) and the measurement layout (Model; Blue), against the expected Brier Score on the test set (orange line).
+The measurement layout outperforms the aggregate for every agent.
+Note: No agent passed more than around 50% of the instances.
+To demonstrate the power of the measurement layout approach, we design a simple experiment with agents with predefined capabilities.
+These agents interact with a three-dimensional embodied environment, Animal-AI (Beyret et al., 2019; Voudouris et al., 2025), and can move forwards, backwards, and turn; their goal is to obtain a green reward in each episode.
+They receive an $n\times n$ RGB image, and turn around until there is a green pixel directly ahead, and then they move towards it.
+We can vary the resolution of their image observations, directly manipulating their capability to see, i.e., their visual acuity.
+We can also vary the speed with which they move towards green rewards, by introducing random stationary actions once they have located a green reward, directly manipulating their navigation capability.
+We generate 1000 agents with different visual acuities and navigation capabilities, and test them on 1000 tasks in Animal-AI consisting of green rewards of different sizes placed at different distances from the agent.
+The average final reward across these 1000 tasks is presented in Figure 2A, demonstrating the effect of manipulating the size of the observation image and the amount of navigational noise.
+Figure 2B presents the measurement layout for this task, relating the goal distance and the goal size to two latent capabilities: visual acuity and navigation capability.
+We relate visual acuity to the quantity $\frac{\text{distance}}{\text{size}}$.
+Smaller, more distant rewards are harder, and therefore require higher visual acuity, than larger, closer rewards.
+Navigation capability is only related to goal distance.
+We use half-normal priors on these latent capabilities, reflecting the fact that there is a meaningful 0 value for both navigation capability and visual acuity, representing no capability.
+We fitted the measurement layout in Figure 2A using 4 chains, with 1000 warm up draws and 2000 sample draws to ensure convergence.
+Figure 2C shows how our measurement layout outperforms the aggregate as a predictor of success for all agents.
+Figure 3 shows that the inferred posterior estimates of the capability neatly correspond to the relative visual acuity and navigation capability of a sample of agents.
+Figure 3: Abilities for a selected subset of vision agents, with means (points) and 95% Highest Density Intervals.
+Left: The navigation abilities for agents with $\times32$ and $\times40$ pixel inputs, for noise levels of 0.0, 0.3, 0.8, and 0.9.
+Differences in navigation are recovered given a pixel input size.
+Right: The visual acuities for agents with $\times4$, $\times20$, and $\times40$ pixel inputs, for low (0.0) and high (0.8) navigational noise.
+Differences in visual acuity are recovered given a navigational noise level.
+Navigation ability is dominated by visual acuity because the agent can only navigate (with noise) to a goal that it can see.
+4 Experiment 2: An Object Permanence Task
+Our second experiment explores a more complex set of capabilities, including Object Permanence.
+To ensure that we knew the ground truth abilities of the agents, we created a synthetic dataset by simulating the performance of 30 synthetic agents based on specified cognitive profiles.
+To achieve this, the team was split into Team A who designed the synthetic agents, and Team B who built the measurement layout without knowledge of the agents built by Team A.
+More details of the specific testbed and the synthetic agents can be found in Appendix C.2.
+We present the measurement layout for the object permanence tasks in Fig.
+4.
+In the object permanence benchmark (Voudouris et al., 2022), there are additional challenges to navigation in the form of ramps, platforms, and lava.
+Subsequently, meta-features for the presence of these challenges, as well as nodes for the agent’s capability (and instance-level performance) at handling them must be added to the measurement layout.
+In order to represent object permanence specifically, we add extra meta-features corresponding to key OP demands.
+A full mathematical description is given in Appendix D.2
 Figure 4: Visual representation of the measurement layout for the object permanence task.
-Team A generated 30 synthetic agents of four different types to ensure a diversity of behaviour. Reference Agents act as reference points, e.g., a perfect agent, an agent with $x$success, etc.
-Achilles Heel Agents have low ability on some dimension, but otherwise perform well and can be said to have object permanence.
-Context-Specific Agents have mixed OP performance, doing well in some specific tests/contexts but not others.
-Fraudster Agents behave in a sophisticated manner that might look like object permanence, but actually is not (Uzgiris and Hunt 1975).
-Tables  6 and  7 in appendix D.4 show the specifics of each agent.
-This information was not disclosed to Team B.
-5.2 Measurement Layout
-Object permanence is a difficult capability to evaluate—combined with the need to visually identify goals, navigate to the goal, avoid hazards, and to remember the location of the object; there are many possible factors influencing overall performance that need to be disambiguated.
-We present the measurement layout for the OP tasks in Fig.
-5.
-The same broad structure as for the AAIO measurement layout is expanded upon.
-In the OP task-suite, there are additional challenges to navigation in the form of ramps, platforms, and lava.
-Subsequently, nodes for the agent’s capability (and instance-level performance) at handling each of these must be added to the layout.
-Meta-features for the presence of each of these challenges are needed.
-To represent object permanence specifically, we add extra meta-features: Occluder Presence, Time Under Occlusion, and Number of Positions (denoting the number of search-location choices available to the agent in forced choice instances).
-We further add capability nodes for Object Permanence and Memory Ability (each with an associated Performance node).
-Memory Performance, Occluder Presence and Number of Positions all feed into object permanence performance.
-The linking function for OP performance is the non-complementary product of Memory Performance and the logistic of the OP margin.
-The OP margin is the difference between Object Permanence and the product of Number of Positions and Time Under Occlusion.
-If no occluder was present the linking function for Object Permanence Performance returns $1$.
-5.3 Inference and Analysis
-Team B took the synthetic agent results and inferred the cognitive profiles using the PyMC implementation of the measurement layout.
+As before, the cognitive profiles of agents were inferred with PyMC using 4 chains, 1000 warm-up samples, and 2000 sample draws, noting one divergences (agent 3) across all agents.
 Overall, our model was successful at inferring system capabilities from patterns of performance.
-We evaluated the model in three ways.
-The first was a qualitative comparison of the cognitive profile descriptions generated by Team B with the criteria used to generate the agents.
-The model facilitated a correct description of the agent in around 75% of cases.
-For the “Achilles Heel” agents in particular, which had object permanence abilities but a specific weakness, the model was highly successful at identifying the cause of failure.
+We evaluated the model in two ways, as we see next.
+Qualitative Evaluation
+The first validation was a qualitative comparison of the cognitive profile descriptions generated by Team B with the criteria used to generate the agents.
+The measurement layout facilitated a correct description of the agent in around 75% of cases.
+Where agents had been designed to have a specific weakness (“Achilles Heel” agents), the model was highly successful at identifying the cause of failure.
+None of the Achilles Heel agents (6,9,11,13) showed degraded performance on Object Permanence when other factors caused failure on specific tasks (Figure 5(b)).
+Agents without Object Permanence (Agent 3) have a reduced score.
+Agent 5 passes all tests with high probability and is given a high OP score.
+However, each Achilles heel agent had its weakness identified in the relevant capability (Figure 6).
 It was, however unable to spot “fraudsters” making use of strategies (such as “go to the last seen location of the reward”) to mimic OP capability.
-Appendix D.4 contains a detailed breakdown of inferred parameters of the cognitive profiles of each agent, more information on the experimental set up, and the full qualitative analysis comparing inferred capabilities to the ground truth capabilities of the agents designed by Team A.
-Capability OP FlatNav Visual Lava Platform Ramp Memory
-Overall RMSE 0.28 0.18 0.16 0.2 0.28 0.38 0.15
-Included RMSE 0.13 0.11 0.24 0.17 0.134 0.27 0.16
-Table 1: RMSE between the ground truth cognitive profile for Team A’s synthetic agents and the inferred values by Team B’s measurement layout.
-Overall: All 30 agents.
-Included: 13 agents generated using only the features included in the model.
-(a) AAIO Tasks
-(b) O-PIAAGETS tasks.
-Figure 6: Brier score of model predictions against agent success rates.
-The second validation involved calculating the RMSE of the model for each ability (normalised to be in $[0,1]$), relative to the ground truth generated by Team A.
-Given in Table 1, model errors were modest across most abilities.
-Understandably, the model struggled most with agents that were created using features not captured by the model.
-For example, agent 23 was generated to have robust object permanence only in cup tasks (passing only 10% of other types).
-As our model did not include task type as a meta-feature, it erroneously took the non-cup task failures as evidence the agent lacks OP as well as struggling with platforms.
-However, given that the cup tasks are the only OP tasks without platforms, this is not an unreasonable assessment.
-By contrast, as the bottom row of the table shows, the model was accurate for agents generated using only features included in the model.
-This highlights the importance of having a measurement layout capturing the important meta-features that most affect performance, and also of designing test-batteries that allow confounding factors to be disentangled.
-For our third validation, we show that our model can be used for accurate prediction of agent success on unseen task instances.
+Appendix E.1 contains a detailed breakdown of inferred parameters of the cognitive profiles of each agent, more information on the experimental set up, and the full qualitative analysis comparing inferred capabilities to the ground truth capabilities of the agents designed by Team A.
+E.2 contains a deeper analysis of the results, looking at different categories of agent in more detail, and E.3 contains the full forest plots for all agents and core capabilities.
+Figure 5: a) Brier score of model predictions against agent success rates within O-PIAAGETS.b) Posterior means and 95% HDI for a selection of interesting agents on Object Permanence capability.
+Figure 6: The posterior means and 95% Highest Density Intervals for a selection of interesting agent types on four capabilities: Lava, Ramps, Right-Left Bias, and Visual Acuity.
+Predictive Evaluation
+For our second validation, we show that our model can be used for accurate prediction of agent success on unseen task instances.
 Taking instance meta-features and applying forward inference with our model yields a probability of success.
-We compare this to the probability predicted by aggregate success using a Brier score.
-The full results are given in Appendix E, but a summary is given in Figure 5(b).
-The model is consistently a better predictor of success than relying on the aggregate measure.
+We compare this to the probability predicted by aggregate success.
+A full analysis is available in Appendix E.3.
+However, broadly, we see that the model is consistently a better predictor of success (lower Brier score; Figure 5(a)) than relying on the aggregate measure.
 We see better prediction performance for agents with less certain success rates.
-These are the agents for which we are more interested in improving prediction as there is more capacity to reduce uncertainty.
+These are the agents for which we are more interested in improving prediction.
+5 Experiment 3: Extending Measurement Layouts To Real Data
+In our third experiment, we adapted the previous measurement layout to construct cognitive profiles for real agents solving object permanence tasks, using data from an independent study (Voudouris et al., 2024).
+We fitted cognitive profiles for deep reinforcement learning agents (Proximal Policy Optimization, PPO, Schulman et al.
+(2017); Dreamer-v3, Hafner et al.
+(2025)) on five different training curricula (Voudouris et al., 2024), as well as human children aged 4–7 years old, a simple heuristic agent similar to those described in Section 3, and a random action agent that moves stochastically in the environment.
+These behavioural data were more complex than those generated for Experiment 2: the agents were evaluated on whether they entered the same location as the occluded reward (choice response) as well as whether they ultimately obtained the reward (success response).
+We therefore generalised the measurement layout to incorporate this multivariate outcome.
+We also made some adjustments to the capabilities and meta-features to be inferred, to account for differences in the datasets.
+A detailed account of the updated measurement layout is given in Appendix D.3, but is also presented in Figure 7.
+We also fitted measurement layouts for children both with and without (ablated) a visual acuity capability to account for the lack of variance in the goal size meta-feature.
+Our measurement layouts were fitted to 10 deep reinforcement learning agents, 1 random agent, 1 heuristic agent, and the combined dataset of all instances from children aged 4–7 (n=1608; with and without visual acuity ablation).
+Four chains were used for each, with a 1000 sample warm-up and 2000 draws.
+Figure 7: The updated measurement layout to accommodate the meta-features and observed variables collected in Voudouris et al. (2024).
+Figure 8: Brier Scores for Success (Left) and Choice (Right). Measurement layouts are more predictive for most agents.
+Figure 8 shows the Brier Scores for the measurement layouts on the held-out test set compared to using the average performance on the train set as the predictor.
+All measurement layouts are better than the Brier Score except for the Dreamer 4, Dreamer 5, PPO 4, and PPO 5 for the success response variable.
+This is likely because performance is so low for these agents that the aggregate measure is very predictive of overall performance.
+Noticeably, we are able to fit a predictive model of the behaviour of children, who were capable of completing the task and also had high levels of variability due to the between-subjects design.
+Figure 9 presents the estimated posteriors for three capabilities: object permanence, navigation, and visual acuity.
+All agents except children have low object permanence capability.
+In contrast, the children are poor navigators, compared to the reinforcement learning agents and particularly the heuristic agent.
+The heuristic agent is useful sanity check—it was designed to navigate only to observable goals.
+We were successfully able to recover the fact that it fails when goals are occluded but is nevertheless a good navigator.
+On visual acuity, it seems that all the reinforcement learning agents were heavily downsampling visual inputs, thus affecting their visual acuity compared to the random and heuristic agents.
+A measurement layout with the visual acuity capability ablated is also shown for children, since they were presented with tasks with no variance in goal size.
+Figure 9: The posterior means and 95% Highest Density Intervals for 13 agent types on three capabilities: object permanence, navigation, and visual acuity.
+6 Constructing Effective Measurement Layouts
+For reasons of both scalability and ease of application, it is prudent to have a generalised methodology for constructing measurement layouts.
+The core of our approach is to identify the demands of a task, and to compare these against the capabilities of the system being evaluated.
+As demands we use the meta-features of each task instance to ensure demands are objective properties of the task.
+We limit demands to meta-features that 1) should meaningfully affect performance (actual difficulty indicators) and 2) suspected to affect performance but should not (potential bias indicators).
+Each demand is related to a capability to provide a margin and typically will use a linking function of $\sigma(capability-demand)$.
+This approach is beneficial because it ties capabilities to bounded, probabilistic performance based on the demand.
+This also causes the capabilities to have the same unit as the demand, and allows us to interpret a capability with value $x$ as consistently succeeding on the demand with meta-feature $x$ in $50\%$ of instances, higher if capability exceeds demand, and lower otherwise.
+Of course, with more knowledge about the task, more intricate linking functions can be used to capture this knowledge succinctly.
+In order to link demands and capabilities with task success we need to understand the dependency structure of the task; that is, which demands need to be met in order to succeed.
+In some cases, capabilities can be compensatory–meeting one demand adequately can account for failing another demand, i.e., the system’s capabilities compensate for each other.
+In other cases, they can be non-compensatory–both capabilities are necessary to reach some level of performance.
+In our experiments, our capabilities were non-compensatory, so we often opted for taking the product of probabilities.
+However, more complex relations such as taking the minimum or maximum of a set of probabilities is available by utilising a generalised mean (preserving differentiability of the link functions).
+The other major property that needs to identified in order to construct the measurement layout is the sources of possible bias.
+In principle, this is similar to creating the structure for capabilities: The meta-features that may affect performance need being identified, however, now those that are not necessary for the task are also required.
+A bias “performance” is now also included for each instance (how present was the bias in this instance, and in which direction was it directed?), but this is now included in the sigmoid function for the relevant capability it may affect: $\sigma(capability-demand+bias)$.
+Once all of the demands and biases are accounted for, a final combination can give an output for overall task performance.
+Breaking a complex task down into demands and possible biases is not always straightforward.
+Care, and domain knowledge of what constitutes success in the task, needs to be used to identify the core broad capabilities required, and how these break down to identifiable meta-features.
+We recommend an iterative approach: if a particular measurement layout cannot explain system behaviour reliably (with a low standard deviation on the resulting posterior distributions), then there is either too few evaluation instances, or the measurement layout is missing an important capability or bias.
+When evaluating artificial intelligence, and for that matter non-human natural intelligence, we need to be careful to avoid anthropomorphising the system.
+We do not want to mistakenly ascribe capabilities or explain behaviours based on a crude facsimile of human behaviour.
+AI systems, in particular, have the potential to solve tasks using significantly different internal mechanisms.
+We have been careful to avoid this mistake in our work by focusing on the required demands for a task rather than presupposing any mechanisms for how the system internally processes meta-features.
+7 Related Work
+Psychometrics and latent variable modelling.
+Measurement layouts are inspired by psychometric traditions that infer latent constructs from patterns of performance.
+Classic techniques such as Item Response Theory (IRT; Gustafsson and Undheim 1996) and Structural Equation Modelling (SEM; Keith and Reynolds 2010) infer latent abilities by analysing populations tested on relatively few items.
+These approaches typically assume homogeneous item types, linear link functions, and population-level data.
+In contrast, machine learning evaluation usually involves thousands of heterogeneous task instances but only a single subject (the model) of interest.
+Our method adapts psychometric ideas to this setting: we derive capability priors from domain knowledge, and infer a cognitive profile for an individual system by triangulating across varied task demands.
+This allows richer inference than task taxonomies or category aggregates, and avoids post-hoc interpretation of statistical factors as “capabilities.” Related Bayesian frameworks such as HiBayES (Luettgau et al., 2025) focus on hierarchical GLM modelling of aggregate performance and uncertainty quantification; our emphasis is instead on capability-oriented inference from instance-level meta-features.
+The ADeLe framework (Zhou et al., 2025) also identifies capabilities in relation to task demands, but does not allow for the complex combination of demands with non-linear linking functions, limiting its explanatory power compared to measurement layouts.
+Cognitive modelling.
+Hierarchical Bayesian models are widely used in cognitive science to explain human behaviour (Lee, 2011).
+These models are often simple at the individual level, designed to capture limited behavioural paradigms.
+By contrast, measurement layouts scale hierarchical Bayesian inference to complex, multi-demand tasks with thousands of instances, and support non-linear, non-compensatory interactions between capabilities.
+Our contribution is to treat capabilities as explicit latent variables linked to meta-features of tasks, rather than black-box latent factors.
+Evaluation frameworks in machine learning.
+Recent ML evaluation efforts emphasise broad benchmarking rather than structured inference.
+Frameworks such as HELM (Liang et al., 2022), Dynabench (Kiela et al., 2021), and related initiatives highlight the need for dynamic and holistic evaluation, but typically report aggregated performance metrics or leaderboards.
+These aggregates are useful but opaque: they predict average success rates without explaining why a model succeeds or fails.
+Our approach is complementary: measurement layouts retain the generality of diverse benchmarks while providing explanatory, capability-level inferences that can predict behaviour on new tasks.
+This goes beyond static aggregate statistics, enabling interpretable and predictive evaluation.
 8 Discussion
-Increasingly more machine learning systems are (pre-)trained once and deployed for a range of tasks, Task-oriented evaluation based on aggregated performance cannot robustly evaluate these systems across the many situations in which they might be deployed, nor predict behaviour for new tasks and changing distributions.
+Increasingly, machine learning systems are (pre-)trained once and deployed for a range of tasks.
+Task-oriented evaluation based on aggregated performance cannot robustly evaluate these systems across the many situations in which they might be deployed, nor predict behaviour for new tasks and changing distributions.
 However, we can often identify abstract capabilities that are not directly measurable, such as ‘object permanence’ in agents or ‘handling negation’ in language models.
-In the behavioural sciences this is a common approach, but many techniques are populational.
-Triangulation for single individuals by contrasting against the task demands is common in animal cognition research, but it is unusual to see complex inferences from thousands of items, as is common in machine learning.
+In the behavioural sciences this is a common approach, but many techniques assume a population of participants tested on a small number of tasks.
+Inferring capabilities from the behaviour of single individuals by contrasting their performance against the task demands is common in animal cognition research, but it is unusual to see complex inferences from thousands of items, as is common and required in machine learning and AI research.
 In this paper we have integrated several elements to make this inferential exercise possible and shown it can identify the capabilities and biases of each model—its cognitive profile—independently.
@@
 We advocate incremental approaches, where each capability is inferred with subsets of tasks before moving to tasks with more complex dependencies.
-Nonetheless, we have shown that even a posteriori, as in the AAIO, inferences can be achieved simultaneously.
-Here we covered only the domain of agentic systems, but the same ideas can be extrapolated to language, vision and other domains.
+Here we applied our method to the domain of agentic systems, but our method applies equally to any evaluation setting, including in natural language processing, computer vision, speech modelling, and reinforcement learning.
 Ultimately, the biggest challenge for the construction of measurement layouts is the need for detailed domain knowledge and robust evaluation batteries.
 Both can be costly, but are urgently needed to improve AI evaluation.
-Currently, evaluation is focused on advancing SOTA performance on an increasingly small selection of benchmarks that often lack construct validity and are limited in scope (Koch et al.
-2021; Raji et al.
-2021).
-Our approach also highlights the need for a clear distinction between training datasets, which can be massive, unannotated, and built in an adversarial way (Kiela et al.
-2021) (Jia and Liang 2017; Zellers et al.
-2019), and evaluation batteries, which should be devised carefully with systematic variation across the elements that might affect performance.
-Our vision is that, from well-constructed batteries such as O-PIAAGETS, building measurement layouts and triangulating with them should become routine.
-Appendices
-Here we include some supplementary material to the paper.
-First we include more details about PyMC, the Bayesian inference engine we are using and the decisions made for the estimation algorithm, and then we give more details and results for the two scenarios; first, the Animal AI Olympics and second, O-PIAAGETS.
-Appendix A PyMC
-PyMC (Salvatier, Wiecki, and Fonnesbeck 2016) is a probabilistic programming language for building Bayesian models and fitting them with Markov Chain Monte Carlo (MCMC) methods (Andrieu et al.
-2003).
-Complex Bayesian models often require solving intractable integrals for inference such as the normalisation constant in Bayes’ theorem,
-[…]
-among others. Monte Carlo (MC) Integration provides a method to approximate these intractable integrals using random sampling.
-In order to evaluate
-[…]
-we can rewrite it as an expectation
-[…]
-for any density function $p$ such that $p(x)>0$.
-MC Integration takes a random sample $X=\{x_{i}\}_{i=1}^{N}\in\Omega$ with distribution $p(x)$ and uses the Law of Large Numbers to approximate the integral as:
-$I\approx\frac{1}{N}\sum_{i=1}^{N}\frac{f(x_{i})}{p(x_{i})}$
-MC Integration is reliant on being able to sample efficiently from $\Omega$ according to $p(x)$.
-MCMC algorithms aim to do this, scaling to complex, hierarchical models by building a Markov Chain with equilibrium distributions approximating the desired probability density function $p$.
-Sampling from the Markov Chain can then be used as a stand-in for sampling $p(x)$.
-The prototypical example of an MCMC algorithm is the Metropolis Hastings algorithm (Hastings 1970).
-Hamiltonian Monte Carlo methods (Betancourt and Girolami 2013) improve on typical MCMC algorithms by replacing the random-walk element exploring the Markov Chain state-space with Hamiltonian Mechanics.
-PyMC requires all expressions to be differentiable, which means that functions such as maximum and minimum are implemented as generalised means, and logical operators are implemented with products and sums.
-A.1 Measurement Layouts in PyMC
-In our work, we utilise the No U-Turn Sampler (Hoffman and Gelman 2014) built into PyMC, an extension to Hamiltonian MC methods which eliminates the need to set particular hyper-parameters.
-With the NUTS sampler we use 2 sampling chains, tuned for $1000$ samples that are discarded in order to optimise the step-size within NUTS, and then draw $1000$ samples to estimate the posterior.
-We define the model once for each scenario and then we estimate the parameters of the cognitive profile and some other INON elements for each system being evaluated independently.
-As mentioned in the introduction, data and code will become available upon publication.
-Appendix B Animal AI Environment and Competition
+Currently, evaluation is focused on advancing state-of-the-art performance on an increasingly small selection of benchmarks that often lack construct validity and are limited in scope (Koch et al., 2021; Raji et al., 2021) or do not actually measure what the benchmark designers originally intended (Browning and LeCun, 2023).
+Further care is needed when evaluating cognitive capabilities; it can be inappropriate to simply apply the same tests that would be used for human subjects (Mitchell, 2023) and there are numerous other pitfalls to avoid and good practices to adhere to (Ivanova, 2023).
+Our approach also highlights the need for a clear distinction between training datasets, which can be massive, unannotated, and built in an adversarial way (Kiela et al., 2021; Jia and Liang, 2017; Zellers et al., 2019), and evaluation batteries, which should be devised carefully perspective, and with systematic variation across the elements that might affect performance.
+Our vision is that, from well-constructed batteries such as those used here, building measurement layouts and triangulating predictive and explanatory capability profiles with them should become routine.
+Acknowledgments
+This work was funded by US DARPA HR00112120007 (RECoG-AI), the Future of Life Institute, FLI, under grant RFP2-152, a ESRC DTP scholarship (ES/P000738/1), the EU (FEDER) and Spanish grant RTI2018-094403-B-C32 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”, Generalitat Valenciana under PROMETEO/2019/098, and EU’s Horizon 2020 research and innovation programme under grant agreement No.
+952215 (TAILOR)
+Appendix A Psychometric Underpinnings
+Measurement layouts are informed by, but distinct from, classic psychometric approaches:
+  1. 1.
+    We do not simply break out performance by ‘categories’.
+Related work employs a ‘taxonomy of tasks’ (see e.g., Table 1 in Liang et al., 2022), each placed into one or more categories.
+This limits the inferential capacity and predictability of these category aggregates for new tasks.
+Instead, we evaluate based on instance-level demands.
+  2. 2.
+    Unlike psychometric approaches such as factor analysis (FA), structural equation modelling (SEM) or item response theory (IRT), we do not rely on populational data (Gustafsson and Undheim, 1996).
+Instead, we infer the cognitive profile of a single subject from their performance data alone.
+  3. 3.
+    We do not simply extract latent variables statistically and then post-hoc interpret them as capabilities, which has been done hierarchically elsewhere (e.g., the Cattell-Horn-Carroll model (Keith and Reynolds, 2010)).
+Instead, first identify prerequisite capabilities based on domain knowledge, and following this, derive general linking functions that allow us to infer and measure the subject’s latent capabilities from the demands of the task.
+  4. 4.
+    Probabilistic models currently used in cognitive modelling (Lee, 2011) are often simple at the level of the individual and cannot convey the required probabilistic expressions connecting individual performance, capabilities and task features.
+Instead, complex interactions allow for Bayesian triangulation which we use for informative evaluation.
+Together, these differences adapt psychometric reasoning to the machine learning setting, where evaluation involves many heterogeneous task instances but often only a single subject of interest.
+Appendix B Experimental Materials
+To illustrate how a measurement layout identifies cognitive profiles, we first select a domain with simple task characterisations that can be straightforwardly related to cognitive profiles to predict performance.
+We chose the Animal-AI (AAI) (Beyret et al., 2019; Voudouris et al., 2025) environment as this domain for the scalability of task complexity it provides.
 Animal-AI is an environment for training and testing AI systems.
-It is built in the Unity engine (Juliani et al.
-2018).
-Animal-AI provides a first-person perspective in a 3D environment.
+It is built in the Unity engine (Juliani et al., 2018).
+Animal-AI provides a single agent with a first-person perspective in a 3D environment.
 The environment can contain a number of pre-specified objects, representing rewards, obstacles, tools, or dangers.
-Tasks can be defined using a domain-specific language to procedurally generate a large number of tasks instances.
-The agent’s interaction with the world is limited to moving around in the 3D space, though it can exploit physics interactions to exert more influence over the environment.
-Fig.
-7 shows two screenshots from separate task instances.
-In the Animal-AI Olympics (Crosby et al.
-2020), an example arena containing the possible objects was provided to contestants.
-They were instructed that their agents would be evaluated on a variety of problems constructed from the objects in the example arena.
-The contestants then needed to build an agent—using any method of their choosing—that they believed would succeed at a wide variety of tasks.
-The tasks that ultimately made up the evaluation suite were based on adaptations of tasks from comparative and developmental literature.
-More information can be found at (http://animalai.org/AAI/).
-For our work with the AAIO data-set we utilised a subset of tasks instances that were solvable with simple goal-directed behaviour. The specific task instances we utilised were:
-[1-1-1, 1-1-2, 1-1-3, 1-2-1, 1-2-2, 1-2-3, 1-3-1, 1-3-2, 1-3-3, 1-4-1, 1-4-2, 1-4-3, 1-5-1, 1-5-2, 1-5-3, 1-6-1, 1-6-2, 1-6-3, 1-7-1, 1-7-2, 1-7-3, 1-8-1, 1-8-2, 1-8-3, 1-9-1, 1-9-2, 1-9-3, 1-10-1, 1-10-2, 1-10-3, 1-11-1, 1-11-2, 1-11-3, 1-12-1, 1-12-2, 1-12-3, 1-13-1, 1-13-2, 1-13-3, 1-14-1, 1-14-2, 1-14-3, 1-15-1, 1-15-2, 1-15-3, 1-16-1, 1-16-2, 1-16-3, 1-17-1, 1-17-2, 1-17-3, 7-1-1, 7-1-2, 7-1-3, 7-2-1, 7-2-2, 7-2-3, 7-3-1, 7-3-2, 7-3-3, 7-4-1, 7-4-2, 7-4-3, 7-5-1, 7-5-2, 7-5-3, 7-6-1, 7-6-2, 7-6-3]
-The types of arenas that these instances correspond to can be found at (http://animalai.org/AAI/testbed).
-Figure 7: Screenshots from two different instances in the Animal AI Olympics that we include in our analysis.
-Appendix C AAIO Agent Results
-This section gives more details for the first domain we examined in Section 4, the Animal-AI Olympics competition dataset. We also give more detail about the specifics of the measurement layouts.
-C.1 Patterns of Performance
-Figure 8: Agent characteristic grid displaying results for some selected agents.
-Top row, from left to right: ‘ironbar’, ‘sparklemotion’, ‘Juohmaru’.
-Bottom row, from left to right: ‘y.yang’, ‘daydayup’, ‘41Animals’.
-Note that the scales of the $x$-axis and $y$-axis are normalised and not the same as in the main body of text.
-To show the patterns of performance of several agents, we use a simple visualisation known as agent characteristic grids, as described in (Burnell et al.
-2022).
-The plots place each feature as a dimension and the success rate represented in colours from green (success) to red (failure).
-Values are grouped into bins of appropriate width in the desired dimensions for analysis.
-Each cell also shows the number of observed instances in that bin.
-Fig.
-8 shows some of the agents discussed in the paper, as well as some other agents that will serve us to illustrate a few points.
-For instance, ‘ironbar’ has 99% accuracy (the grid is mostly green everywhere) for this set of tasks (like ‘Trrrr’, the winner of the competition) and is identified as saturating both capabilities in Fig.
-4.
-The next one, ‘sparklemotion’, with 36% accuracy, is very non-conformant, with navigationAbility $0.40\pm 0.38$ and visualAbility $0.59\pm 0.45$, and high noise $0.31\pm 0.13$.
-Proportionally, navigationAbility is worse as the grid looks even less conformant on the distance of the reward than the reward size.
-The rightmost on the top row is ‘Juohmaru’, with 54% accuracy, which is a very conformant agent, with red concentrating on the top and right of the grid (the objectively most difficult instances).
-Its navigation ability is $2.50\pm 1.08$ and visual ability is $0.92\pm 0.44$, and medium noise: $0.26\pm 0.17$.
-We have specifically chosen some particular agents in the second row for which the extracted cognitive profile does not capture the whole performance patterns completely.
-Leftmost on the second row, ‘y-yang’ has 70% accuracy and a very strange grid, with red cells in a diagonal, indicating a very non-conformant agent.
-This translates into navigationAbility $3.82\pm 0.45$ and visualAbility $1.68\pm 0.21$, but medium noise $0.22\pm 0.17$.
-This is so because overall the agent has high accuracy, so a bit of noise can explain the few anomalies on the top right of the grid.
-In the middle of the second row is ‘daydayup’ with 52% accuracy.
-This agent shows a higher navigation ability than visual ability, as can be seen in the grid, where the size of the reward is the predictive feature and reward distance is less relevant.
-This translates into navigationAbility $3.48\pm 1.28$ (high than average but with uncertainty) and visualAbility $0.92\pm 0.44$ (low), with some a high level of random noise affecting performance $0.48\pm 0.20$.
-Finally, ‘41Animals’ with 87% accuracy, shows relatively high values and low deviations for both navigationAbility $4.37\pm 0.59$ and visualAbility $1.81\pm 0.11$, with very little noise $0.05\pm 0.04$ despite having some non-conformant behaviour in the diagonal.
-In this case the left-right bias is higher ($0.45\pm 0.91$) and may account for this.
-In general, for those agents where accuracy is too low or too high, the differences are more subtle.
-Of course, the interesting part of Fig.
-9 is the agents in the middle.
-This suggests that estimating capabilities can benefit from a more focused analysis of instances around the demand thresholds.
-For ’Juohmaru’, this would amount to selecting instances by difficulty, instead of simply selecting all instances for analysis.
-C.2 Animal AI Olympics Measurement Layout
-Table 2 provides more information about the task meta-features that we utilised to construct the measurement layouts.
-Note that these ranges determine the ranges of the capabilities, so that we have units.
-For instance, if distance to the reward were measured in metres then the navigation capability would have metres as a unit.
-Let the navigation capability of an agent be, say, 20 metres.
-Since it is defined with a logistic function of distance, that means that the agent can achieve excellent performance when well below 20, performing at chance around the 20 metre mark, and failing every instance when the reward is well beyond 20 metres away.
-Dimension
-Description
-Range/Values
-Reward Size
-The size of the reward (the sign is modified so lower is larger)
-$[0,1.9]$
-Reward Distance
-Euclidean distance from the agent’s start position to the reward
-$[0,5.3]$
-Reward Behind
-Whether the reward originates behind (1), left or right (0.5) or in front (0) of the agent
-$\{0,0.5,1\}$
-XPos
-Whether the reward is to the left ($-1$), the right (1) or is centred (0) relative to the agent’s start position
-$\{-1,0,1\}$
-Table 2: AAIO Task Characterisation details. The dimensions on which instances vary in our measurement layouts.
+Tasks can be defined using a domain-specific language to procedurally generate a large number of task instances.
+The agent’s interaction with the world is limited to moving around in 3D space, though it can exploit physical interactions (such as pushing certain objects) to generate more complex behavioural sequences.
+In the main text, we explore two evaluation test batteries developed in AAI.
+In Experiment 1 (Section 3), we design some simple tasks for measuring the navigation and visual acuity capabilities of hand-designed agents.
+In Experiments 2 and 3 (Sections 4 and 5), we use the Object-Permanence in Animal-Ai: GEneralisable Test Suites (O-PIAAGETS) test bed (Voudouris et al., 2022).
+We describe each test battery in turn.
+B.1 Simple Navigation And Visual Acuity Test Battery
+To demonstrate the utility of measurement layouts, we use a simulation data from agents with known capabilities.
+The aim of the measurement layout is to infer an agent’s visual acuity (its capability to see) and its navigation ability (its capability to navigate towards rewarding objects).
+Figure 10: A series of tasks with the agent in a fixed position and goals of different sizes in 1 of 100 different positions in the arena.
+In the Animal-AI Environment, a battery of tasks can be easily defined.
+The agent is placed in the arena at a fixed position and random starting rotation, along with a single rewarding GoodGoal of a certain size placed at some location in the arena.
+The agent must locate the goal and navigate towards it.
+The agent passes the instance if they obtain the goal, and fails otherwise.
+We procedurally generated 1000 instances of arenas in Animal-AI, with the agent in a fixed position at $(20,0,0.5)$ and goals in 100 evenly spaced positions across the arena.
+There are 10 possible $x$-coordinates and 10 possible $z$-coordinates, at $[2,6,10,14,18,22,26,30,34,38]$ for each.
+The Euclidean distance between the centre of the agent and the centre of the goal was computed, encoding the distance demand.
+There were 10 different sizes for each goal, $[0.2,0.4,0.6,0.8,1.0,1.2,1.4,1.6,1.8,2.0]$.
+Figure 10 presents three tasks with goals of different sizes placed in different positions.
+The smaller the goal or the further it is from the agent, the higher the visual acuity required for the agent to see it.
+This is because the ‘retinal image’ of the object decreases for smaller, more distant goals.
+There are two task demands, therefore, goal distance and goal size.
+We want the latent visual acuity capability estimate to be higher if the agent tends to pass instances where the goal is further away and/or smaller.
+As such, we want the overall demands of the instance (some combination of size and distance) to be higher.
+Retinal size, which can be defined as size divided by distance, decreases for ‘harder’ tasks.
+So an alternative is to invert retinal size, by dividing distance by size.
+This value increases when the goal has a smaller retinal size from the starting point.
+In summary, by defining the visual acuity demand as the inverse of the retinal size, we can produce a scale for latent visual acuity estimates with interpretable units, and therefore, interpretable magnitudes.
+Figure 2 b) presents a measurement layout relating these two task demands to performance.
+B.2 O-PIAAGETS
+To demonstrate the Measurement Layout framework’s ability to handle more complex tasks and infer complex psychological capabilities, we also apply it to a subset of the O-PIAAGETS test battery (Voudouris et al., 2022).
+O-PIAAGETS is also built in the Animal-AI environment and is designed to assess object permanence.
+O-PIAAGETS consists of three broad task types from the animal and developmental sychological literature for assessing object permanence and a set of control tasks testing for basic skills that are required to solve more complex OP tasks.
+In all experiments, two experimental paradigms were selected from O-PIAAGETS: the Primate Cognition Test Battery and Chiandetti & Vallortigara Tests of Intuitive Physics.
+The corresponding instances from the control battery for these experimental paradigms were also used, as well as a series of simple instances.
+We describe each of these tasks below.
+In all figures, red arrows denote the location of rewards and grey arrows denote the location of the agent.
+B.2.1 Basic Controls
+A series of instances are used that involve the basic objects in Animal-AI 3 in simple combinations.
+In this chapter, these are collectively called the Basic Control Tasks.
+This includes instances where the goal rolls from one side to the other, or is on a blue platform obtainable by navigating up a ramp.
+There are some tasks that require detouring around opaque and transparent walls and death zones, or to make simple choices (see Figure 11).
+Figure 11: Basic Control Tasks.
+Top Left: A forced choice task—the agent is spawned on a platform with a green reward on the right and a death zone on the left.
+Top Right: An avoidance task—the agent must navigate around the death zone to obtain the green reward.
+Bottom: A selection of further simple tasks.
+B.2.2 Chiandetti & Vallortigara Tests of Intuitive Physics
+The instances in this paradigm are referred to collectively in this chapter as the CV Chick Tasks.
+Figure  12 presents one task.
+Here, the agent is frozen on a platform, watching a reward roll directly away from it through a death zone.
+Then the lights go out, during which time the goal is deflected behind the wall on the left.
+When the lights come back on the agent is unfrozen and must infer that the goal is occluded on the left.
+This demands OP, because the agent must behave as though the object continues to exist even though it is (a) not visible, and (b) not conceivably occluded where it was last observed (in the death zone).
+This task varies: the colour of the occluding wall, the side on which the goal is occluded, and the presence of another wall on the other side.
+These walls are either: transparent, thinner than the width of the goal, or lower than the height of the goal, meaning that they would not be able to occlude the goal were it there.
+There are two special cases of this test.
+In one, there are two identical occluders on each side and no lights out period.
+The agent watches the reward roll behind one occluder, and must navigate towards it.
+In the second, there are two identical occluders which fall backwards during the lights out period.
+One occluder falls (almost) flat to the ground, meaning that the goal could not be behind it or under it.
+The other occluder falls only slightly, meaning the goal could be (and is) underneath it.
+These tasks also vary in wall colour and whether the goal is to be found on the right or the left.
+Figure 12: An instance of a CV Chick Task from the Object Permanence test battery. A–C: Three stages of the task.
+Figure 13 presents a version of a CV Chick Task that does not demand OP.
+The arena is constructed identically, with the agent on a platform and a death zone in the middle.
+The agent is frozen, and watches the goal roll to the left in front of a wall.
+There are variants with and without lights out periods.
+The goal is therefore never occluded.
+These tasks vary: the size of the goal, the colour of any wall objects, the side that the goal rolls to, the presence of a lights out period, the height or width of potential occluders, and the presence of transparent walls.
+There are variants of the tasks with falling walls, except that the wall falls forwards and the goal is always visible.
+Figure 13: An instance of a CV Chick Task from the Room Practice Control test battery, which does not test for OP.
+A–C: Three stages of the instance.
+Every instance has a lights out and a no lights out variation.
+B.2.3 The Primate Cognition Test Battery
+The two kinds of task inspired by the Primate Cognition Test Battery (PCTB) from O-PIAAGETS were included in this study.
+Instances from the first kind are collectively called Three-Cup Tasks here.
+The left image of Figure 14 presents an instance for testing OP.
+The agent is spawned in front of three ramps, separated by coloured walls.
+These simulate the cups under which the experimenter hides rewards in the original Primate Cognition Test Battery tasks.
+Goals fall from a height into one or more of the cups, dropping behind a wall so that they become occluded.
+In cases where only one goal is dropped, the ramps allow the agent to enter the cup but not exit it.
+When two goals drop, there are two ramps in cups with multiple goals, but only one ramp for the incorrect ramp, so if the agent makes the incorrect choice, they fail.
+These tasks vary: the distance of the agent from the cups, the type of goal (yellow or green), the size of the goal, the colour of the occluding walls, the number of goals, and the presence of death zones next to the ramps (to make navigation harder).
+Figure 14: Instances of a PCTB Cup Task. Left: A version of the task from the Object Permanence test battery. Right: A version of the task from the Room Practice Control test battery.
+The right image of Figure 14 presents a control version of the PCTB Three-Cup Task, which does not require OP.
+In this case, the walls between the cups are transparent, and there are no walls behind which the goal is occluded once it has dropped.
+These tasks vary: the distance of the agent from the cups, the type of goal (yellow or green), the size of the goal, the colour of any walls in the arena, the number of goals, the presence of death zones next to the ramps, the presence of ramps, and the presence of transparent walls in front of the goal(s).
+Instances from the second kind of PCTB task are collectively called PCTB Grid Tasks here.
+The agent is spawned atop a large ramp, with a view over a series of holes in the floor, demarcated by white rims.
+A single green goal drops down one of the holes, and the agent must navigate to the correct one.
+In the left image of Figure 15, the goal drops down one of 12 holes.
+Once it has dropped, it is no longer visible, testing OP.
+In the right image of Figure 15 is a version of the Grid Task that does not test OP.
+Here, the holes are shallow, goals continue to drop from a height, but remain visible in their position.
+Both the OP and control tasks vary: the number of holes (4, 8, 12), the size of the goal, and the hole into which the goal drops.
+Figure 15: Instances of a PCTB Grid Task. Left: A version of the task from the Object Permanence test battery. Right: A version of the task from the Room Practice Control test battery.
+B.2.4 Task Overview
+These tests are relatively simple in terms of what is required of the agent compared to some other tasks in O-PIAAGETS.
+They also contain a large number of variants, controlling for several alternative hypotheses.
+Moreover, the PCTB Tasks and the CV Chick Tasks are an important dyad, because they control for distinct hypotheses about behaviour.
+One sophisticated policy that could lead to good performance on PCTB Tasks would be to navigate to where the goal was last observed.
+This is not necessarily the same as having OP: the agent need not have any representation that the same goal exists at the last observation location; they could represent the goal as having disappeared while having learnt that they tend to get rewarded with a ‘new’ goal when they navigate to the last observation location.
+This would be a successful strategy for the PCTB Tasks, but not the CV Chick Tasks, which require that the agent represent the goal as continuing to exist, so that they can make the inference that it must be occluded behind one of the walls.
+In the other direction, the CV Chick Task could be solved by a policy that navigates behind the largest wall in the scene.
+This policy would lead to success on many CV Chick tasks, but it would not lead to success on the PCTB tasks, where there are either three walls of equal width in the case of the Three-Cup tasks, or no walls at all in the case of the Grid tasks.
+Thus, using both the PCTB tests and the CV Chick tests is a particularly useful subsection of O-PIAAGETS to focus on.
+Appendix C Evaluation Subjects
+For each experiment, we used a different set of subjects (agents) for evaluation. We describe each in turn.
+C.1 Experiment 1: A Simple Navigation Task
+Each agent has two parameters: $p$, where $p\times p\times 3$ is the size of the pixel input from the environment with three colour channels (R, G, B), and $\nu$, where $1-\nu$ is the probability of taking an action at a time step and $\nu$ is the probability of taking no action (i.e., staying still).
+The agent rotates left in its starting position.
+If there are green pixels to the right or left of the centre of the image, it rotates accordingly until they are in the centre, with probability $1-\nu$.
+If green pixels are present in the centre of the image, it navigates forward with probability $1-\nu$.
+The pixel input size constrains the agents visual acuity, by limiting what it can see in the environment.
+The navigation noise constrains the ability of the agent to navigate towards a reward, with higher values lowering the agents navigational ability.
+We simulate 100 agents on our 1000 tasks, using pixel inputs ranging from $\times4$ to $\times40$ in increments of 4 and navigational noise ranging from 0 to 0.9 in increments of 0.1.
+C.2 Experiment 2: An Object Permanence Task
+Object Permanence (OP) is the understanding that objects continue to exist when they go out of view.
+Behaviourally, an agent has object permanence if they behave as though objects continue to exist even when they cannot directly perceive them.
+It is a key component of physical common-sense (Shanahan et al., 2020; Lake et al., 2017; Baillargeon et al., 2010a), allowing biological agents to more successfully predict and interact with their environment.
+Object permanence is well studied in cognitive science, where several experimental designs are used to detect its presence in human and non-human animals (Scholl, 2007; Flombaum and Scholl, 2006; Hare and Tomasello, 2005; Chiandetti and Vallortigara, 2011).
+Many of these paradigms are represented in O-PIAAGETS, which contains over 22,000 instances that test OP (as well as over 250,000 control instances that test other skills necessary to complete the main battery of instances).
+We selected three paradigms that test for allocentric object permanence, where objects are occluded independently of the observer’s actions.
+We also include basic control tasks within the test-suite to evaluate the agent’s non-OP capabilities and make the inference process more robust.
+Details about the test-bed design criteria, the included tasks, and why object permanence is difficult to identify are given in Appendix E.
+This scenario tests whether our methods can recover complex cognitive profiles from performance data of agents for which we know the ground truth.
+To do so, the the authors of this paper were divided into two teams, Team A and Team B.
+Team A generated the synthetic dataset, while Team B defined the measurement layout and attempted to recover the cognitive profiles of the synthetic agents.
+In total, we included 2188 unique instances from O-PIAAGETS, varying along the 12 dimensions outlined in Table  1.
+Note the asterisks in the table, denoting that a particular feature was not used in the measurement layouts used for evaluation in this paper.
+This occurred because when constructing the measurement layout, Team B either did not realise the feature was present in the environment or because it was missed as a possible source of bias or capability.
+Such discrepancies are an unsurprising result of the two-team blind set-up, but also mirrors a situation in which the full ground truth of a population is unnkown.
+Nevertheless, this actually highlights a strength of our evaluation framework; imperfect measurement layouts can still be useful for both inferring agent capabilities and predicting future agent performance.
+Dimension Description Values
+Reward Visibility* Is the reward visible at the choice point? 0,1{0,1}
+Time Under Occlusion Time reward is occluded $[0,4]$
+Positions Number of alternative reward places $[1,11]$
+Reward Distance Manhattan distance to the goal $[0,56]$
+Right/left Position Is reward left, right or straight from the agent? $\{-1,0,1\}$
+Reward Size How large is the reward? $[0.5,4]$
+Occluder Presence Are there any occluders hiding the rewards? $\{0,1\}$
+Lava Presence Is there any lava in the arena? 0,1{0,1}
+Platform Presence Are there any platforms in the arena? 0,1{0,1}
+Ramp Presence Are there any ramps in the arena? 0,1{0,1}
+Transparent Walls Are there any transparent walls in the arena? 0,1{0,1}
+Occluder Redness* What is the R value of occluders in the arena? $[0,255]$
+Occluder Greenness* What is the G value of occluders in the arena? $[0,255]$
+Occluder Blueness* What is the B value of occluders in the arena? $[0,255]$
+Table 1: Object Permanence Task Characterisation details: The dimensions on which the instances vary. Dimensions with an asterisk are not included in the measurement layout presented in this paper.
+Team A generated 30 synthetic agents of four different types to ensure a diversity of behaviour.
+Reference Agents act as reference points, e.g., a perfect agent, an agent with xx% success for various values of $x$.
+Achilles Heel Agents have low ability on some dimension, but otherwise perform well and can be said to have object permanence.
+Context-Specific Agents have mixed OP performance, doing well in some specific tests/contexts but not others.
+Fraudster Agents behave in a manner that might look like object permanence, but are actually using rigid heuristics such as “go to last seen location of reward, then search” (Užgiris and Hunt, 1975).
+Tables  5,  6, and 7 in Appendix E.1 show the specifics of each agent.
+This information was not disclosed to Team B.
+C.3 Experiment 3: Extending Measurement Layouts To Real Data
+We use the dataset from Voudouris et al.
+(2024) to fit measurement layouts for realistic data collected from deep reinforcement learning agents and human children, as well as a heuristic agent and a random action agent.
+We use the random agent to define a baseline for chance performance.
+This agent takes one of the nine available actions available in the Animal-AI Environment and repeats it for a number of steps, $S$ sampled from a uniform distribution: $S\sim U(0,20)$, after which it randomly selects a new action.
+It executes an action for that many steps before selecting a new one.
+There are no biases to select certain actions.
+We use the heuristic agent as a more meaningful baseline, as an agent that lacks object permanence, but possesses many other capabilities that would lead to success in the Animal-AI Environment.
+We do not expect random walkers to fail tests of OP because they lack only the capability of OP.
+Rather, we expect them to fail because they lack many other capabilities, such as the capability to move towards an appetitive reward or away from an aversive stimulus (goal-directedness).
+The heuristic agent follows a small set of rules and heuristics.
+This agent can view objects in the $x$-$z$ plane that are in the forward-facing 60° of its view using raycasts.
+When it detects a green or a yellow goal in this viewing angle, it orients itself towards it and moves forward to obtain it.
+When it detects a red goal or a death zone, it orients itself away from it and moves backwards.
+If the agent is stationary or if there is a wall in front of it, the agent moves either forwards and left or forwards and right with a probability pp=0.5, perseverating that movement subsequently to navigate around the object until the agent is no longer stationary or there is no longer a wall in front of it.
+If the agent does not detect any goals, death zones, or walls, and if it is not stationary, it perseverates its previous action.
+Data from children aged 4–7 years old performing on a subset of the O-PIAAGETS tasks was also used as a meaningful baseline for an agent that is error-prone but possesses object permanence and other ancillary capabilities necessary to complete tasks in Animal-AI Baillargeon et al.
+(2010b); Užgiris and Hunt (1975); Piaget (1923).
+Data from children was collected between January and March (inclusive) 2023 by the authors of Voudouris et al.
+(2024), with a minimum target sample size of 30.
+The study was conducted at a university, in the presence of researchers and the participants’ guardians.
+Guardians were provided with an information sheet and the opportunity to consent to their child’s participation in the study.
+If they consented, the study was initiated.
+Participants watched a three-minute video introducing the study as a game called ‘Get the Fruit!’.
+Participants were invited to assist ‘Farmer John’ in finding all the green and yellow ‘apples’ in the farmyard, while avoiding any poisonous red fruit and ‘large puddles of lava rain.’ They were informed that they would receive stickers in return for fruit collection, although they actually received one sticker every ten trials, regardless of whether they successfully found the goal, and an extra sticker upon completion of the study.
+After the video, participants commenced play with the Animal-AI Environment, interacting with it through a small hand-held controller connected to a computer.
+First, they played a ‘tutorial’ round, consisting of 11 tasks.
+They could play these tasks as many times as they wished, and ask for help throughout.
+Then, they played the ‘test’ round, consisting of 38 Test tasks.
+These tasks were ordered randomly, with half of participants playing the fixed random order and the other half playing the reverse of that order.
+Every 10 tasks, the participant could take a break from the game and receive their stickers.
+The break consisted of a simple task with a single large green goal and no time limit.
+When the participant was ready to continue playing, they simply had to navigate to the goal to initiate the next batch of tasks.
+During the test round, guardians filled out a short survey about the participant, with questions on age, gender, and videogame playing habits.
+Upon completion of the test round, participants and their guardians were debriefed and offered a book voucher for their participation.
+The deep reinforcement learning agents, Dreamer-v3 and PPO, were each trained on 5 different curricula, resulting in 5 different trained agents (see Table 2).
+Table 2: The 5 different trained agents for each architecture (Dreamer-v3 & PPO; 10 total), their training curricula, and the number of training steps.
+Stratification was done by experimental paradigm.
+Agent Name Training Curriculum Training Steps (Millions)
+PPO/Dreamer 1 All Basic tasks 2M
+PPO/Dreamer 2 All Basic tasks ++ a stratified random sample of 300 Control tasks 4M
+PPO/Dreamer 3 All Basic tasks ++ all Control tasks (in 3 randomly sampled batches) 8M
+PPO/Dreamer 4 All Basic tasks ++ a stratified random sample of 300 Control tasks ++ a stratified random sample of 300 Test tasks 6M
+PPO/Dreamer 5 All Basic tasks ++ all Control tasks (in 3 randomly sampled batches) ++ a stratified random sample of 300 Test tasks 10M
+Appendix D Measurement Layout Specifications
+D.1 Experiment 1: A Simple Navigation Task
+The measurement layout for the simple navigation task consists of two capability nodes (navigation, visual acuity) and two meta-feature nodes (goal distance, goal size), as depicted in Figure 2B.
+Each of the capability nodes has an associated instance-performance node, an intermediate non-observable node (INON).
+To compute probabilities at these nodes, we use the sigmoid function over the difference between estimated capability and the current meta-feature value.
+For visual acuity, this is the goal distance divided by the goal size.
+We apply the natural logarithm to this difference to improve numerical stability.
+For navigation, the meta-feature is simply the goal distance.
+We assume a non-compensatory relationship between navigation and visual acuity, and therefore multiply these marginal INON probabilities to condition the Bernoulli distribution over binary (success/failure) outputs.
+We use the half-normal prior over capabilities with a $\sigma$ equal to half of the maximum relevant meta-feature or meta-feature combination.
+D.2 Experiment 2: An Object Permanence Task
+The measurement layout for OP consists of seven capability nodes, one bias node, and one robustness node (stochastic noise on success/failure).
+Again, each of the capability and bias nodes has an associated instance-performance node, an intermediate non-observable node (INON).
+The noise is applied at the end to the final performance probability.
+A full description is given in Table 3 including all the value ranges, distributional priors and linking functions.
 Name Type Range/Values Prior/Linking Function
-rewardSize Meta-Feature $[0,1.9]$
-rewardDistance Meta-Feature $[0,5.3]$
-rewardBehind Meta-Feature $\{0,0.5,1\}$
-XPos Meta-Feature $\{-1,0,1\}$
-navigationAbility Capability $[0,5.3]$ $ScaledBeta(1,1,0,5.3)$
-visualAbility Capability $[0,1.9]$ $ScaledBeta(1,1,0,1.9)$
-rightleftBias Bias [,][$\infty$,$\infty$][ - ∞ , ∞ ] $\mathcal{N}(0,1)$
+rampPresence Meta-feature $\{0,1\}$
+lavaPresence Meta-feature $\{0,1\}$
+platformPresence Meta-feature $\{0,1\}$
+goalDistance Meta-feature $[0,56]$
+goalSize Meta-feature $[0.4,5]$
+rightleftPosition Meta-feature $\{-1,0,1\}$
+occluderPresence Meta-feature $\{0,1\}$
+timeUnderOcc Meta-feature $[0,4]$
+mumberOfPositions Meta-feature $[1,11]$
+OPAbility Capability $[0,48.4]$ $Uniform(0,\sf memoryAbility\times maxPositions)$
+memoryAbility Capability $[0,4.4]$ $Uniform(0,4.4)$
+visualAbility Capability $[0,6]$ $Uniform(0,6)$
+rampAbility Capability $[0,1]$ $Uniform(0,1)$
+lavaAbility Capability $[0,1]$ $Uniform(0,1)$
+platformAbility Capability $[0,1]$ $Uniform(0,1)$
+flatNavAbility Capability $[0,56]$ $\mathcal{N}(0,1)$
+rightleftBias Bias $\infty$$\infty$] $\mathcal{N}(0,1)$
 noiseLevel Robustness $[0,1]$ $Uniform(0,1)$
-rightleftEffect INON $[-1,1]$ ${\sf rightleftBias}\times{\sf XPos}$
-navigationPerformance INON $[0,1]$ $\sigma({\sf NavigationAbility}-{\sf rewardDistance}\times(\frac{1}{2}{\sf rewardBehind}+1)+{\sf rightleftEffect})$
-visualPerformance INON $[0,1]$ $\sigma({\sf visualAbility}-{\sf rewardSize})$
-taskPerformance Observed $[0,1]$ $p\sim Bernoulli(\omega[{\sf noiseLevel}]({\sf navigationPerformance}\times{\sf visualPerformance},\nu))$
-Table 3: Nodes in the Animal AI Measurement Layout, including range/values, priors and linking function.
-The measurement layout for this suite of tasks consists of two capability nodes (navigation and visual acuity), one bias node (whether the goal is initially on the left or the right of the agent), and one robustness node (stochastic noise on success/failure).
-Each of the capability and bias nodes has an associated instance-performance node, its intermediate non-observable node (INON).
-The noise is applied at the end.
-In Table 3 we have complete details of all the nodes in the measurement layout. The function $\sigma$ represents the standard logistic function:
-$\sigma(x)=\frac{1}{1+e^{-x}}$
-while $\omega$ represents a weighting:
-$\omega[\alpha](b,c)=(1-\alpha)\cdot b+\alpha\cdot c$
-The value $\nu$ represents a noise prior that represents a constant model set to 1 -- average performance.
-Finally, $ScaledBeta$ represents a scaled beta distribution to an interval different from $[0,1]$.
-Note that a beta distribution has two parameters that need to be estimated.
-Figure 9: Inferred cognitive profile elements, along with the success rate, for the 68 agents in the AAIO.
-The gradient colour red-yellow-green denotes how high the means are (green is highest) while the gradient circles white-grey-black denote how high the standard deviations are (black is highest).
-Finally, Fig.
-9 shows the complete inference results for the 68 agents in the AAIO.
-The agents are ordered by the aggregate success rate at the collection of task instances.
-We see that, for the agents in the middle of the table in particular, our approach is able to decompose agent capabilities at this task into the constituent navigation and visual abilities.
-Bias has a more diverse behaviour throughout the table, with strong biases for extremely low scoring agents too, such as with agents ‘ahorizon’ and ‘Octopus’.
-Equally, we can see that the detected bias can account for poor performance.
-Compare ‘Thursday’ and ‘MadStorks’ which share a similar overall success rate (0.30 vs.
-0.28) despite the fact that ‘Thursday’ has much more navigation ability than ‘MadStorks’ and they have similar visual abilities.
-This is explained by ‘Thursday”s much larger bias towards moving to the right (0.73 vs 0.4) even if the reward is to the left.
-Appendix D O-PIAAGETS Battery
-Now we look at the object permanence test battery in more detail.
-We look at the overall design criteria for the test-bed, the types of tasks involved, and how the tasks are characterised.
-We also give full details for the measurement layout devised for this task and an extended view of the results inferred by our approach.
-D.1 Design Criteria
+rampPerformance INON $[0,1]$ $\mu({\sf rampAbility},{\sf rampPresence})$
+platformPerformance INON $[0,1]$ $\mu({\sf platformAbility},{\sf platformPresence})$
+lavaPerformance INON $[0,1]$ $\mu({\sf lavaAbility},{\sf lavaPresence})$
+flatnavPerformance INON $[0,1]$ $\sigma({\sf flatNavAbility}-{\sf goalDistance}+{\sf rightLeftEffect})$
+navigationPerformance INON $[0,1]$ ${\sf rampPerformance}\times{\sf platformPerformance}\times{\sf lavaPerformance}\times{\sf flatnavPerformance}$
+memoryPerformance INON $[0,1]$ $\sigma({\sf memoryAbility}-{\sf timeUnderOcc})$
+OPPerformance INON $[0,1]$ ${\sf memoryPerformance}\times\sigma(48.4-((48.4-({\sf OPAbility}-{\sf 4}\times{\sf numPositions}))\times{\sf occluderPresence}))$
+visualAcuityPerformance INON $[0,1]$ $\sigma({\sf visualAbility}-({\sf 5}-{\sf goalSize}))$
+rightleftEffect INON $\infty$$\infty$] ${\sf rightLeftBias}\times{\sf rightLeftPosition}$
+taskPerformance Observed $[0,1]$ $p\sim Bernoulli(\omega[{\sf noiseLevel}]({\sf OPPerformance}\times{\sf navigationPerformance}\times{\sf visualAcuityPerformance},\nu))$
+Table 3: Nodes in the Object Permanence Measurement Layout, including priors and linking function.
+The function $\sigma$ is the standard logistic function while $\omega$ is a parameterised weighting function, where the value $\nu$ represents a noise prior that represents a constant model set to 1 - average performance.
+The function $\mu$ is a margin, defined as
+$\mu(a,b)=1-((1-a)\cdot b)$
+where $a$ is an ability and $b$ a binary feature.
+This means that when the binary feature is 0 then the margin represents $p(success)=1$.
+If the binary feature = 1 then $p(success)=a$.
+The units of some of the abilities derive from the meta-features, like the memory ability, which is seconds, as it comes from the opposing metafeature timeUnderOcc.
+Similarly, visualAbility has the same units as goalSize, and flatNavAbility has the same units as goalDistance.
+OPAbility comes from a product between timeUnderOcc and numPositions so its unit is seconds times positions.
+Clearly, the most complicated linking function is the one for OPPerformance, which basically is opposed to (subtracted by) the maximum time under occlusion (4) times the number of positions.
+This is subtracted from the maximum OPAbility (48.4) and multiplied by the occluderPresence which is subtracted again from the maximum OPAbility (48.4).
+A logistic function is applied as usual but then this is multiplied by memoryPerformance, representing the dependency between object permanence and memory.
+D.3 Experiment 3: Extending Measurement Layouts To Real Data
+To accommodate the data from Voudouris et al. (2024), we updated the measurement layout, which is presented in Figure 16.
+Figure 16: The updated measurement layout to accommodate the meta-features and observed variables collected in Voudouris et al. (2024).
+There are three central capabilities: OP, navigation, and visual acuity.
+There are four further capabilities that are pertinent to navigation: the capability to avoid death zones, and the capabilities to navigate to goals placed to the left, directly ahead, and to the right of the agent’s starting position.
+A full description is given in Table 4 including all the value ranges, distributional priors and linking functions.
+Name Type Range/Values Prior/Linking Function
+Number of Positions Meta-feature $\{0,12\}$
+Goal Occluded Meta-feature $\{0,1\}$
+Goal Distance Meta-feature $[9,118]$
+Number of Turns To Goal Meta-feature $[0,13]$
+Choice Distance Meta-feature $[1,53.5]$
+Number of Turns To Choice Meta-feature $[0,6]$
+Goal Size Meta-feature $[0.5,5]$
+Death Zone Presence Meta-feature $\{0,1\}$
+Goal Ahead Meta-feature $\{0,1\}$
+Goal Left Meta-feature $\{0,1\}$
+Goal Right Meta-feature $\{0,1\}$
+OPAbility Capability $0,\infty]$ $\text{HalfNormal}(503.1)$
+Flat Navigation Ability Capability $[0,4.4]$ $\text{HalfNormal}(163.80)$
+Visual Acuity Capability $0,\infty]$ $\text{HalfNormal}(47.80)$
+Lava Ability Capability $0,\infty]$ $\text{HalfNormal}(1)$
+Ahead Ability Capability $0,\infty]$ $\text{HalfNormal}(1)$
+Right Ability Capability $0,\infty]$ $\text{HalfNormal}(1)$
+Left Ability Capability $0,\infty]$ $\text{HalfNormal}(1)$
+OPSuccessPerformance INON $[0,1]$ $\sigma({\sf OPAbility},{\sf GoalOccluded*(GoalDistance-ChoiceDistance)*NumberOfPositions})$
+OPChoicePerformance INON $[0,1]$ $\sigma({\sf OPAbility},{\sf GoalOccluded*ChoiceDistance*NumberOfPositions})$
+FlatNavGoalPerformance INON $[0,1]$ $\sigma({\sf FlatNavAbility},{\sf GoalDistance*NumberTurnsGoal})$
+FlatNavChoicePerformance INON $[0,1]$ $\sigma({\sf FlatNavAbility},{\sf ChoiceDistance*NumberTurnsChoice})$
+visualAcuityPerformance INON $[0,1]$ $\sigma({\sf visualAbility}-\text{log}({\sf goalDistance}/{\sf goalSize}))$
+lavaPerformance INON $[0,1]$ $\mu({\sf lavaAbility},{\sf lavaPresence})$
+AheadPerformance INON $[0,1]$ $\mu({\sf AheadAbility},{\sf AheadPresence})$
+LeftPerformance INON $[0,1]$ $\mu({\sf LeftAbility},{\sf LeftPresence})$
+RightPerformance INON $[0,1]$ $\mu({\sf RightAbility},{\sf RightPresence})$
+SuccessPerformance Observed $[0,1]$ $p\sim Bernoulli({\sf FlatNavGoalPerformance}\times{\sf OPSuccessPerformance}\times{\sf visualAcuityPerformance}\times{\sf lavaPerformance})$
+ChoicePerformance Observed $[0,1]$ $p\sim Bernoulli({\sf FlatNavChoicePerformance}\times{\sf OPChoicePerformance}\times{\sf visualAcuityPerformance}\times{\sf AheadPerformance}\times{\sf LeftPerformance}\times{\sf RightPerformance})$
+Table 4: Nodes in the Object Permanence Measurement Layout, including priors and linking function.
+Appendix E Evaluating the Presence of Object Permanence
 Object Permanence (OP), and similar cognitive capabilities such as episodic memory and theory of mind, are difficult to robustly identify in AI systems because there are often several competing alternative explanations for why an agent is succeeding or failing at a particular task.
@@
 Conversely, an agent lacking OP might pass this task, simply because it has been rewarded previously for going forwards some number of steps and then turning left.
-This is the problem of the underdetermination of theory by behaviour (Stanford 2006), in which an observed behaviour can be explained in multiple ways.
+This is the problem of the underdetermination of theory by behaviour (Stanford, 2006), in which an observed behaviour can be explained in multiple ways.
 The use of well-informed measurement layouts based on internally valid test batteries allows us to more robustly eliminate incorrect alternative explanations and triangulate on the underlying causal explanations of AI system behaviour.
@@
     Instances are not necessarily evenly distributed across the levels of a dimension, leading to the potential for sampling biases.
-These considerations were used as design criteria for the test-bed presented in the next subsection.
-D.2 Object Permanence Tasks
-The tasks that we use in our OP test-bed are based on a subset of the O-PIAAGETS object permanence test battery introduced in (Voudouris et al.
-2022).
-The battery in its entirety is vast (including over 13,000 task instances requiring object permanence and over 250,000 basic task instances that don’t require OP, but help control for other potential explanations.
-Here, we adapt the following three separate task types from O-PIAAGETS.
-Primate Cognition Test Battery Cup Tasks
-Herrmann et al.
-developed the Primate Cognition Test Battery (PCTB) for testing aspects of physical and psychological common-sense in primates (Herrmann et al.
-2007).
-One task in this battery was designed to assess OP.
-The participant is presented with three identical inverted cups, under which the experimenter hides a reward.
-Once the reward has been occluded, the participant is invited to lift a cup.
-If they fail to locate the reward, the trial ends.
-The reasoning here is that if the participant correctly locates the reward without any evidence of trial-and-error, there is evidence that they have OP, understanding that the reward continues to exist even when it is out of view.
-Figure 10: Primate Cognition Test Battery 3-Cup (’Cup’) Task. Top: Top-down view; bottom: Agent-view. Red arrow indicate reward location, grey arrows indicate agent location.
-O-PIAAGETS has a battery of 598 instances based on this experimental set-up, including 359 that require object permanence to be solved and 239 that do not.
-In the OP tests, the agent observes a green reward drop from a height behind an occluder.
-They must traverse a ramp, sometimes avoiding lava, to obtain it.
-If they traverse the wrong ramp, they cannot return, and they will fail the instance, making this a forced choice test.
-See Fig.
-10.
-In the instances that do not require OP, the reward is always visible, either through transparent objects or due to the lack of occluders.
-These instances serve as controls, so that we can evaluate whether poor performance on tasks requiring OP can be explained not by a lack of object permanence per se, but rather by poor performance traversing ramps, avoiding lava, or navigating towards rewards.
-Primate Cognition Test Battery Grid Tasks
-Extending the general structure for testing OP, Crosby et al.
-developed a task where rewards can be occluded in one of up to 12 different positions, through holes in the floor (Crosby et al.
-2020).
-O-PIAAGETS contains 430 instances based on this design, with 4, 8, or 12 holes (see Fig.
-11).
-191 instances require OP to be solved, with the remaining 239 not requiring OP, where the holes are not deep enough to occlude the goal.
-These latter instances serve as controls.
-Figure 11: Primate Cognition Test Battery Grid (’Grid’) tasks.
-Top: Top-down view; bottom: Agent-view.
-Red arrow indicate reward location, grey arrows indicate agent location.
-Reproduced from Voudouris et al., 2022.
-Chiandetti & Vallortigara Chick Tasks
-Chiandetti and Vallortigara developed a series of tests for investigating physical common-sense in day-old chicks (Chiandetti and Vallortigara 2011), which Crosby et al.
-developed for use in the Animal-AI Environment (Crosby et al.
-2020).
-In the example in Fig.
-12, the agent observes a reward rolling away (panel A).
-Then the lights go out, blocking all visual input for several time steps (panel B).
-During this lights out period, the reward is deflected either left and rolls behind the occluder.
-When the lights turn on, the reward is fully occluded (panel C).
-If the agent understands that the reward continues to exist even when it is out of view, they ought to reason that it is behind the occluder on the left, as there is nowhere it could be hiding on the right.
-Figure 12: Chiandetti and Vallortigara (’CV’) Task. Top: Reward Rolling away; middle: ”Lights out” period; bottom: Reward occluded. Reproduced from Voudouris et al., 2022
-We used 36 instances from O-PIAAGETS that tested OP in this way, and included 1195 further tasks that did not require OP to be solved, to serve as controls.
-Basic Tasks
-All of these tasks require an agent to be able to navigate towards green goals, around obstacles, while traversing platforms and ramps, and avoiding lava.
-When testing for OP in biological agents such as primates or chicks, it is often safe to assume that they understand the value of rewards such as food or that they are able to navigate or search for rewarding objects.
-This is not a safe assumption when evaluating artificial agents, particularly those trained purely with techniques from machine learning.
-Therefore, the testbed must include instances that test for these basic skills.
-We included a further 521 instances testing basic abilities to navigate towards reward, avoid lava, and use ramps.
-These are extensions and variations of the tests used in section  4.
-In total, we included 2188 unique instances from O-PIAAGETS, varying along the 12 dimensions outlined in Table  4.
-Note the asterisks in the table, denoting that a particular feature was not used in the measurement layouts used for evaluation in this paper.
-This occurred because Team B either didn’t realise the feature was present in the environment or because it was missed as a possible source of bias or capability.
-Nevertheless, this actually highlights a strength of our evaluation framework; imperfect measurement layouts can still be useful for both inferring agent capabilities and predicting future agent performance.
-Dimension
-Description
-Range/Values
-Reward Visibility*
-Is the reward visible at the choice point?
-${0,1}$
-Time Under Occlusion
-Time reward is occluded
-$[0,4]$
-Positions
-Number of alternative reward places
-$[1,11]$
-Reward Distance
-Manhattan distance to the goal
-$[0,56]$
-Right/left Position
-Is reward left, right or straight from the agent?
-$\{-1,0,1\}$
-Reward Size
-How large is the reward?
-$[0.5,4]$
-Occluder Presence
-Are there any occluders hiding the rewards?
-$\{0,1\}$
-Lava Presence
-Is there any lava in the arena?
-${0,1}$
-Platform Presence
-Are there any platforms in the arena?
-${0,1}$
-Ramp Presence
-Are there any ramps in the arena?
-${0,1}$
-Transparent Walls
-Are there any transparent walls in the arena?
-${0,1}$
-Occluder Redness*
-What is the R value of occluders in the arena?
-$[0,255]$
-Occluder Greenness*
-What is the G value of occluders in the arena?
-$[0,255]$
-Occluder Blueness*
-What is the B value of occluders in the arena?
-$[0,255]$
-Table 4: Object Permanence Task Characterisation details. The dimensions that instances vary on. The dimensions with an asterisk are not included in the measurement layout presented in this paper.
-This test-bed satisfies our criteria for capturing and evaluating OP that we outlined above:
-  1. 1.
-    Object permanence is difficult to place on an ordinal scale. In cognitive science, it is usually discussed in terms of its presence or absence.
-  2. 2.
-    Object permanence relies on working memory capacity and spatial cognition, and successful performance on tests of OP often require good navigational skills.
-The dimensions defining each task vary on their relevance to OP and these precursor abilities (working memory, spatial cognition, navigational ability).
-There is therefore a complex interaction between latent capabilities and task dimensions.
-  3. 3.
-    Instances were not defined on all dimensions, introducing missingness. For example, in many instances, there were no occluders, so Occluder Colour is undefined.
-  4. 4.
-    Due to the structure of Animal-AI and O-PIAAGETS, some instance-types are represented more than others.
-For example, we have 36 Chiandetti & Vallortigara Chick Tasks testing OP, compared to 359 Primate Cognition Test Battery Cup Tasks, almost 10 times more.
-D.3 Object Permanence Measurement Layout
-This section extends the information about the measurement layout in section 5.2 and Fig.
-5.
-The measurement layout for OP consists of seven capability nodes, one bias node, and one robustness node (stochastic noise on success/failure).
-As in the AAIO case, each of the capability and bias nodes has an associated instance-performance node, an intermediate non-observable node (INON).
-The noise is applied at the end to the final performance probability.
-Name Type Range/Values Prior/Linking Function
-rampPresence Meta-feature $\{0,1\}$
-lavaPresence Meta-feature $\{0,1\}$
-platformPresence Meta-feature $\{0,1\}$
-goalDistance Meta-feature $[0,56]$
-goalSize Meta-feature $[0.4,5]$
-rightleftPosition Meta-feature $\{-1,0,1\}$
-occluderPresence Meta-feature $\{0,1\}$
-timeUnderOcc Meta-feature $[0,4]$
-mumberOfPositions Meta-feature $[1,11]$
-OPAbility Capability $[0,48.4]$ $Uniform(0,48.4)$
-memoryAbility Capability $[0,4.4]$ $Uniform(0,4.4)$
-visualAbility Capability $[0,6]$ $Uniform(0,6)$
-rampAbility Capability $[0,1]$ $Beta(1,1)$
-lavaAbility Capability $[0,1]$ $Beta(1,1)$
-platformAbility Capability $[0,1]$ $Beta(1,1)$
-flatNavAbility Capability $[0,56]$ $\mathcal{N}(0,1)$
-rightleftBias Bias [,][$\infty$,$\infty$][ - ∞ , ∞ ] $\mathcal{N}(0,1)$
-noiseLevel Robustness $[0,1]$ $Uniform(0,1)$
-rampPerformance INON $[0,1]$ $\mu({\sf rampAbility},{\sf rampPresence})$
-platformPerformance INON $[0,1]$ $\mu({\sf platformAbility},{\sf platformPresence})$
-lavaPerformance INON $[0,1]$ $\mu({\sf lavaAbility},{\sf lavaPresence})$
-flatnavPerformance INON $[0,1]$ $\sigma({\sf flatNavAbility}-{\sf goalDistance}+{\sf rightLeftEffect})$
-navigationPerformance INON $[0,1]$ ${\sf rampPerformance}\times{\sf platformPerformance}\times{\sf lavaPerformance}\times{\sf flatnavPerformance}$
-memoryPerformance INON $[0,1]$ $\sigma({\sf memoryAbility}-{\sf timeUnderOcc})$
-OPPerformance INON $[0,1]$ ${\sf memoryPerformance}\times\sigma(48.4-((48.4-({\sf OPAbility}-{\sf 4}\times{\sf numPositions}))\times{\sf occluderPresence}))$
-visualAcuityPerformance INON $[0,1]$ $\sigma({\sf visualAbility}-({\sf 5}-{\sf goalSize}))$
-rightleftEffect INON [,][$\infty$,$\infty$][ - ∞ , ∞ ] ${\sf rightLeftBias}\times{\sf rightLeftPosition}$
-taskPerformance Observed $[0,1]$ $p\sim Bernoulli(\omega[{\sf noiseLevel}]({\sf OPPerformance}\times{\sf navigationPerformance}\times{\sf visualAcuityPerformance},\nu))$
-Table 5: Nodes in the Object Permanence Measurement Layout, including priors and linking function.
-Table 5 gives the complete details, broken down by node type.
-As in the AAIO case, the function $\sigma$ is the standard logistic function while $\omega$ is a parameterised weighting function, where the value $\nu$ represents a noise prior that represents a constant model set to 1 -- average performance.
-The function $\mu$ is a margin, defined as
-$\mu(a,b)=1-((1-a)\cdot b)$
-where $a$ is an ability and $b$ a binary feature.
-This means that when the binary feature is 0 then the margin represents $p(success)=1$.
-If the binary feature = 1 then $p(success)=a$.
-The units of some of the abilities derive from the meta-features, like the memory ability, which is seconds, as it comes from the opposing metafeature timeUnderOcc.
-Similarly, visualAbility has the same units as goalSize, and flatNavAbility has the same units as goalDistance.
-OPAbility comes from a product between timeUnderOcc and numPositions so its unit is seconds times positions.
-Clearly, the most complicated linking function is the one for OPPerformance, which basically is opposed to (subtracted by) the maximum time under occlusion (4) times the number of positions.
-This is subtracted from the maximum OPAbility (48.4) and multiplied by the occluderPresence which is subtracted again from the maximum OPAbility (48.4).
-A logistic function is applied as usual but then this is multiplied by memoryPerformance, representing the dependency between object permanence and memory.
-D.4 Synthetic Agents and Comparison
-The synthetic agents Team A designed can be found in Table  6 and Table  7.
+E.1 Synthetic Agents and Comparison
+The synthetic agents Team A designed can be found in Table  5, Table 6, and Table 7.
 Name Agent Class Agent Description
-Agent 1
-Reference
-Passes every instance.
-Agent 2
-Reference
-Passes each instance with a probability of 0.5.
-Agent 3
-Reference
-Passes each instance with a probability of 0.1.
-Agent 4
-Reference
-Approximates a random walker. Passes a small proportion of instances. It is more likely to pass ’safe’ instances that don’t have lots of dangerous holes to fall down or lava to run into.
-Agent 5
-Reference
-Passes each instance with a probability of 0.9. This is a more human-like failure rate than agent 1.
-Agent 6
-Achilles Heel
-Has OP, passing an instance with a probability of 0.95 unless there is lava present, in which case it passes with a probability of 0.1.
-Agent 7
-Achilles Heel
-Has OP, passing an instance with a probability of 0.95 unless there is lava or an occluder with a ‘redness‘ over 200 present, in which case it passes with a probability of 0.1.
-Agent 8
-Achilles Heel
-Has OP, passing an instance with a probability of 0.95 unless there is an occluder with a ‘greenness‘ over 200 present, in which case it passes with a probability of 0.1.
-Agent 9
-Achilles Heel
-Has OP, passing an instance with a probability of 0.95 unless there is a ramp present, in which case it passes with a probability of 0.1.
-Agent 10
-Achilles Heel
-Has OP but lacks fine control of actions.
+Agent 1 Reference Passes every instance.
+Agent 2 Reference Passes each instance with a probability of 0.5.
+Agent 3 Reference Passes each instance with a probability of 0.1.
+Agent 4 Reference Approximates a random walker.
+Passes a small proportion of instances.
+It is more likely to pass ’safe’ instances that don’t have lots of dangerous holes to fall down or lava to run into.
+Agent 5 Reference Passes each instance with a probability of 0.9. This is a more human-like failure rate than agent 1.
+Agent 6 Achilles Heel Has OP, passing an instance with a probability of 0.95 unless there is lava present, in which case it passes with a probability of 0.1.
+Agent 7 Achilles Heel Has OP, passing an instance with a probability of 0.95 unless there is lava or an occluder with a ‘redness‘ over 200 present, in which case it passes with a probability of 0.1.
+Agent 8 Achilles Heel Has OP, passing an instance with a probability of 0.95 unless there is an occluder with a ‘greenness‘ over 200 present, in which case it passes with a probability of 0.1.
+Agent 9 Achilles Heel Has OP, passing an instance with a probability of 0.95 unless there is a ramp present, in which case it passes with a probability of 0.1.
+Agent 10 Achilles Heel Has OP but lacks fine control of actions.
 If there is lava present, it only passes those instances with a probability of 0.5.
@@
 For the non-OP instances and the remaining instances, it solves an instance with a probability of 0.8.
-Agent 11
-Achilles Heel
-Has OP but poor visual acuity.
+Table 5: The agents Team A designed and synthesised performances for. Agents 1-10.
+Name Agent Class Agent Description
+Agent 11 Achilles Heel Has OP but poor visual acuity.
 If goals are 0.5 or smaller, they only pass the instance with a probability of 0.1.
 For goals of size 1, the probability is 0.3, it is 0.5 for size 1.5, 0.7 for size 2, 0.8 for size 2.5, and 0.9 for a goal size 3 or larger.
-Agent 12
-Achilles Heel
-Has OP but is biased to the right. Probability of passing a task is 0.95 when the goal is to the right, 0.3 when it is to the left, and 0.6 when it is centrally located.
-Agent 13
-Achilles Heel
-Has OP but is biased to the left. Probability of passing a task is 0.95 when the goal is to the left, 0.3 when it is to the right, and 0.6 when it is centrally located.
-Agent 14
-Achilles Heel
-Has OP, but is easily confused when there are several positions that are similar looking (either identical or mirror images of each other).
-Agent 15
-Fraudster
-Does not have OP, but goes to where they last saw the reward. Deterministically passes all instances where this policy works (all tasks except some of CV Chick tasks), and fails the rest.
-Agent 16
-Fraudster
-Does not have OP, but goes to where they last saw the reward.
+Agent 12 Achilles Heel Has OP but is biased to the right. Probability of passing a task is 0.95 when the goal is to the right, 0.3 when it is to the left, and 0.6 when it is centrally located.
+Agent 13 Achilles Heel Has OP but is biased to the left. Probability of passing a task is 0.95 when the goal is to the left, 0.3 when it is to the right, and 0.6 when it is centrally located.
+Agent 14 Achilles Heel Has OP, but is easily confused when there are several positions that are similar looking (either identical or mirror images of each other).
+Agent 15 Fraudster Does not have OP, but goes to where they last saw the reward.
+Deterministically passes all instances where this policy works (all tasks except some of CV Chick tasks), and fails the rest.
+Agent 16 Fraudster Does not have OP, but goes to where they last saw the reward.
 Passes all instances where this policy works with probability of 0.95 (all tasks except some of CV Chick tasks), and passes the rest with probability 0.25.
-Table 6: The agents Team A designed and synthesised performances for.
-Name Agent Class Agent Description
-Agent 17
-Achilles Heel
-Has OP, but has low memory, so struggles when the goal is occluded for longer.
-Agent 18
-Achilles Heel
-Has OP, but is easily confused when there are more positions that the goal could be occluded in.
+Agent 17 Achilles Heel Has OP, but has low memory, so struggles when the goal is occluded for longer.
+Agent 18 Achilles Heel Has OP, but is easily confused when there are more positions that the goal could be occluded in.
 Passes instances with more than 10 other positions available with probability of 0.3.
@@
 Less than 3, passes with probability of 0.95.
-Agent 19
-Context-Specific
-Non-robust OP. It can solve the navigation tasks and all the CV tasks, but fails the other instances.
-Agent 20
-Context-Specific
-Non-robust OP. It solves the navigation tasks and all the CV tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
-Agent 21
-Context-Specific
-Non-robust OP. It can solve the navigation tasks and all the Grid tasks, but fails the other instances.
-Agent 22
-Context-Specific
-Non-robust OP. It can solve the navigation tasks and all the Grid tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
-Agent 23
-Context-Specific
-Non-robust OP. It can solve the navigation tasks and all the 3 cup tasks, but fails the other instances.
-Agent 24
-Context-Specific
-Non-robust OP. It can solve the navigation tasks and all the 3 cup tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
-Agent 25
-Context-Specific
-Non-robust OP. It can solve the navigation tasks and all the CV and grid tasks, but fails the other instances.
-Agent 26
-Cognitive Pathology
-Non-robust OP. It can solve the navigation tasks and all the CV and grid tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
-Agent 27
-Cognitive Pathology
-Non-robust OP. It can solve the navigation tasks and all the 3 Cup and grid tasks, but fails the other instances.
-Agent 28
-Cognitive Pathology
-Non-robust OP. It can solve the navigation tasks and all the 3 cup and grid tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
-Agent 29
-Cognitive Pathology
-Non-robust OP. It can solve the navigation tasks and all the CV and 3 cup tasks, but fails the other instances.
-Agent 30
-Cognitive Pathology
-Non-robust OP. It can solve the navigation tasks and all the CV and 3 cup tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
-Table 7: The agents Team A designed and synthesised performances for (continued).
+Agent 19 Context-Specific Non-robust OP. It can solve the navigation tasks and all the CV tasks, but fails the other instances.
+Agent 20 Context-Specific Non-robust OP. It solves the navigation tasks and all the CV tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
+Table 6: The agents Team A designed and synthesised performances for. Agents 11-20.
+Name Agent Class Agent Description
+Agent 21 Context-Specific Non-robust OP. It can solve the navigation tasks and all the Grid tasks, but fails the other instances.
+Agent 22 Context-Specific Non-robust OP. It can solve the navigation tasks and all the Grid tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
+Agent 23 Context-Specific Non-robust OP. It can solve the navigation tasks and all the 3 cup tasks, but fails the other instances.
+Agent 24 Context-Specific Non-robust OP. It can solve the navigation tasks and all the 3 cup tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
+Agent 25 Context-Specific Non-robust OP. It can solve the navigation tasks and all the CV and grid tasks, but fails the other instances.
+Agent 26 Cognitive Pathology Non-robust OP. It can solve the navigation tasks and all the CV and grid tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
+Agent 27 Cognitive Pathology Non-robust OP. It can solve the navigation tasks and all the 3 Cup and grid tasks, but fails the other instances.
+Agent 28 Cognitive Pathology Non-robust OP. It can solve the navigation tasks and all the 3 cup and grid tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
+Agent 29 Cognitive Pathology Non-robust OP. It can solve the navigation tasks and all the CV and 3 cup tasks, but fails the other instances.
+Agent 30 Cognitive Pathology Non-robust OP. It can solve the navigation tasks and all the CV and 3 cup tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
+Table 7: The agents Team A designed and synthesised performances for. Agents 21-30.
+E.2 Object Permanence Qualitative Results
 Figs.
-13 and 14 show the models inferred values for each element of the cognitive profile.
+19 and 20 show the models inferred values for each element of the cognitive profile.
 Fig 19 gives the mean of the inferred distribution and Fig .20 the standard deviation.
@@
 Equally, when looking at Agents 13 and 14, the right-left bias was high (with low standard deviation) leading Team B to believe that these agents were massively favouring movement in particular directions.
+We also present Forest Plots for agents, showing their 95% HDI across the 8 core capabilities in Figures 17 and 18.
+(a) Object permanence
+(b) Flat navigation
+(c) Visual acuity
+(d) Lava
+Figure 17: 95% HDIs for all agents across Object Permanence, Flat Navigation, Visual Acuity, and Lava.
+(a) Platform
+(b) Ramp
+(c) Right Left Bias
+(d) Memory
+Figure 18: 95% HDIs for all agents across Platforms, Ramps,
 In Fig.
@@
 The relatively lower accuracy here was because task type was not included in the measurement layout.
-Therefore, many of the biases recovered by Team B reflect the particular construction of the paradigms used by Team A (see legend of Fig.
-15).
+Therefore, many of the biases recovered by Team B reflect the particular construction of the paradigms used by Team A (see legend of Figure 21).
 Team B did not accurately recover the cognitive profiles of the ‘fraudster’ agents.
@@
 This results in generally good performance in all tasks except the Chiandetti & Vallortigara tasks, where this policy would result in falling into lava.
-Therefore, the performance distributions look similar to a high performing object permanence agent.
+When this ‘Fraudster’ policy would be successful, these agents succeed with a probability of 0.95, if the policy would not be applicable, the agents succeeds with probability 0 (for agent 15) or $0.25$ (for agent 16).
+Therefore, the performance distributions look similar to a high performing object permanence agent and our approach attributed higher values of OP than were warranted.
 To successfully discriminate these sophisticated agents, we would need to weight the importance of certain sub-groups of instances, such that, for example, Chiandetti & Vallortigara tasks are of most inferential importance in the case where there is good performance on the remaining paradigms.
-This requires targeted inference and diligent use and application of measurement layouts.
+This requires good alignment between those designing the task and those creating the measurement layouts to make sure that controls established in the task design are reflected in the analysis.
 Overall, however, Team B was able to extract a lot of information about a the majority of the synthetic agents—despite an imperfectly designed measurement layout.
 Also noteworthy is that when the measurement provides correct inferences, they tend to be consistent across that particular agent.
-That is, if the measurement layout identifies a particular weakness in an agent, it won’t also pick a second, nonexistent weakness also.
+That is, if the measurement layout identifies a particular weakness in an agent, it won’t also pick a second, nonexistent weakness.
+Capability OP FlatNav Visual Lava Platform Ramp Memory
+Overall RMSE 0.28 0.18 0.16 0.20 0.28 0.38 0.15
+Included RMSE 0.13 0.11 0.24 0.17 0.13 0.27 0.16
+Table 8: RMSE between the ground truth cognitive profile for Team A’s synthetic agents and the inferred values by Team B’s measurement layout.
+Overall: All 30 agents.
+Included: 13 agents generated using only the features included in the model.
 Figure 19: Mean inference results for the cognitive profile elements, along with the mean success rate, for the 30 synthetic agents in the OP task.
@@
 In a gradient from green-yellow-red, with green representing a successful identification and red an unsuccessful one.
-Appendix E Prediction Results
-This section gives the full results for prediction on unseen instances.
+E.3 Object Permanence Predictive Results
 The synthetic dataset was partitioned into an 80/20 train-test split.
@@
 For each instance from the test-set we apply forward inference to yield a success probability.
-We evaluate these probabilities using a Brier score (Brier 1950), also making use of the decomposition into Calibration and Refinement (Murphy 1973).
-These scores are given in Table 8, with the better performing approach for prediction in bold for each agent.
+We evaluate these probabilities using a Brier score (Brier, 1950), also making use of the decomposition into Calibration and Refinement (Murphy, 1973).
+Figure 5(a) plots Brier score against average success rate.
+The measurement layout model is consistently a better predictor of success than relying on the aggregate measure.
+We see better prediction performance for agents with less certain success rates.
+These are the agents for which we are more interested in improving prediction as there is more capacity to reduce uncertainty.
+These scores are also given fully in Table 9, with the better performing approach for prediction in bold for each agent.
 As expected, the aggregate predictions have no refinement at all, and the values are close to 0.25 for balanced cases (the lower numbers only appear when the proportion of right/wrong performance is more imbalanced).
@@
 When we look at calibration, however, we see aggregate predictions are almost perfect, since train and test come from the same distribution.
-But the calibration from the measurement layout is also quite good, and overall there are many more wins for the measurement layouts approach.
+But the calibration from the measurement layout is quite high, and overall there are many more cases where the measurement layout approach is superior than the aggregate.
 Moreover, it is important to emphasise that unlike this table, train and test do not have to come from the same distribution.
 For instance, if we removed the easy instances for many of the tasks, we would get similar results for the measurement layout prediction approach (calibration and refinement should not change dramatically), but the results for the aggregate prediction approach would be much worse, since the calibration would be completely lost.
-Looking at the particular agent properties that cause the measurement layouts to be more or less capable at prediction can provide some insight into the strengths and weaknesses of our approach.
-Agent 1 was designated as the perfect agent that will always pass every instance.
-It is no surprise then that predictions using the aggregate success rate were very successful, achieving the best possible Brier score.
-Our model’s predictions were still highly accurate (Brier score of 0.022), but the model lacked the extreme certainty of the aggregate measure—likely because using Bayes’ rule, beginning with an uninformed prior it’s impossible to update the posterior to exclude one of the outcomes.
-Agent 5 is another case where the aggregate is better at prediction than our measurement layouts.
-Agent 5 passed all instances with probability 0.9.
-We believe our model struggled to perform as well as the aggregate here due to there not being sufficiently many instances to move the initial prior (uniform distribution) to the more extreme distribution predicting success at a rate of 0.9.
-Our model was able to cope with other fixed pass rates (such as with Agent 2, 3, and 4), so this may just be a matter of degree.
+However, doing this
+Looking at the particular agent properties that result in the measurement layouts to be more or less capable at prediction can provide some insight into the strengths and weaknesses of our approach.
+Agents 1 and 5 were designed to pass all instances with probability 1 and 0.9 respectively.
+It’s no surprise that predictions using the aggregate success rate were very successful.
+Our model’s predictions were still highly accurate (Brier scores of 0.022 and 0.099 respectively), but the model lacked the extreme certainties of the aggregate measures.
+For Agent 1, this would be impossible—using Bayes’ rule and an uninformed prior it’s not possible to update the posterior to completely exclude one of the outcomes.
+For agent 5 we believe that our model didn’t achieve the same level of certainty of success due to there not being sufficiently many instances to move the initial prior (uniform distribution) to the more extreme distribution predicting success at a rate of 0.9.
+Our model was able to cope with other fixed pass rates (such as with agents 2, 3, and 4), so this may just be a matter of degree.
 It’s worth noting that for any fixed pass rate agent, it’s actually impossible to predict more accurately than the aggregate in expectation.
@@
 These two agents were designed as ‘Fraudster’ agents that mimic object permanence by going to the last position where they saw the reward.
-This is sufficient to solve many object permanence tasks except for some of the CV chick tasks.
-When this ‘Fraudster’ policy would be successful, these agents succeed with a probability of 0.95, if the policy would not be applicable, the agents succeeds with probability 00 (for agent 15) or $0.25$ (for agent 16).
-We believe it is the high overall success rate that causes our model to be less predictive here, similar to Agent 5 (and to a lesser extent Agent 7 and 8), again because it takes more task instances to move an uninformative policy to extremes such as predicting success with a probability greater than $0.9$.
+We discussed these agents and their strategy in the Section 4 about Qualitative Validity.
+Since these agents had very high overall success rates, we believe that this causes our model to be less predictive, similar to Agent 5 (and to a lesser extent Agent 7 and 8), because it takes more task instances to move an uninformative policy to extremes such as predicting success with a probability greater than $0.9$.
 Agent ML BS Agg. BS ML Cal. Agg. Cal. ML Ref. Agg. Ref.
@@
 30 0.205 0.232 0.013 0.000 0.191 0.232
-Table 8: Brier Scores (BS) and the breakdown into calibration (Cal.) and refinement (Ref.) for the Measurement Layouts (ML) and aggregate predictions (Agg.)
-Appendix F Methodology for Designing Measurement Layouts
-For complex, multifaceted tasks measurement layouts have the potential to grow and become large and complex themselves.
-It’s therefore important that to apply them to other domains (and reap the benefits of the additional evaluation information they provide) to have a general methodology for constructing the appropriate measurement layout structure.
-To this end, we developed a methodology for constructing measurement layouts that we wish to explicate here more clearly.
-As stated in the main paper, the core of our approach is to identify the demands of a task, and to compare these against the capabilities of the system being evaluated.
-Tasks are composed of multiple instances that vary in the degree of demands.
-In our methodology we tie demands with meta-features in order to give the demands values that are tied to objective properties of the task.
-For each demand in a task, a system will have varying levels of performance at dealing with the demand based on its capability.
-we model this as a probability of the system successfully handling the aspect of the task with its meta-feature value.
-This is where the linking function comes into play.
-We often leveraged the sigmoid function $\sigma(capabiltiy-demand)$.
-This approach is beneficial because it ties capabilities to probabilistic performance based on the demand.
-This also causes the capabilities to have the same unit as the demand, and allows us to interpret a capability with value $x$ as consistently succeeding on the demand with meta-feature $x$ in $50\%$ of instances.
-Certain demands also require that other demands have been met.
-This requires a dependency understanding of the demands of the task.
-Performances for dependent task demands can be incorporated into the sigmoid margin calculation or can have an effect on the resulting performance value.
-The exact approach for combining demand performances will depend on the task itself, and whether the demands are compensatory—whether performance can be achieved with one capability compensating for lacking another.
-Our environments were non-compensatory, so we often opted for taking the product of probabilities.
-However, more complex relations such as taking the minimum or maximum of a set of probabilities is available by utilising a generalised mean.
-The other major aspect that we need to identify to construct the measurement layout is the sources of possible bias.
-In principle, this is similar to creating the structure for capabilities: We still identify meta-features that may affect performance, however, now we are looking for those that are not necessary for the task.
-We still include a bias ”performance” for each instance (how present was the bias in this instance, and in which direction was it directed?) , but this is now included in the sigmoid function for the relevant capability it may affect: $\sigma(capability-demand+bias)$.
-Once all of the demands and biases are accounted for, a final combination can give an output for overall task performance.
-Breaking a complex task down into demands and possible biases is not always straightforward.
-Care, and domain knowledge of what constitutes success in the task, needs to be utilised to identify the core broad capabilities required, and how these break down to identifiable meta-features.
-We recommend an iterative approach: If a particular measurement layout cannot explain system behaviour reliably (with a low standard deviation on the resulting posterior distributions), then there is either too few evaluation instances, or the measurement layout is missing an important capability or bias.
-When evaluating Artificial Intelligence systems, and for that matter non-human Biological Intelligences, we need to be careful to avoid anthropomorphising the system.
-We do not want to mistakenly ascribe capabilities or explain behaviours based on a crude facsimile of human behaviour.
-AI systems, in particular, have the potential to solve tasks using incredibly different internal methods and skills.
-We have been careful to avoid this mistake in our work by focusing on the required demands for a task rather than any model for how the system internally processes meta-features.
-Appendix G Computation Code, and Datasets
-Our experiments did not require a large amount of compute, We ran all of our experiments on Google Colaboratory (Google 2017) using the free tier of service without utilising the GPU.
+Table 9: Brier Scores (BS) and the breakdown into calibration (Cal.) and refinement (Ref.) for the Measurement Layouts (ML) and aggregate predictions (Agg.).
+Bold indicates whether the Measurement Layout or Aggregate Brier score was lower (and therefore better).
+Appendix F Supplementary Experiment: Animal AI Olympics
+This experiment exemplifies that evaluation suites that weren’t designed for this type of nuanced analysis can still provide considerable insights.
+The Animal AI Olympics (AAIO) was an open competition evaluating AI agents in the AAI 3D environment, where agents had to get a reward by locating it and navigating in an arena.
+In the Animal-AI Olympics (Crosby et al., 2020), an example arena containing the possible objects was provided to contestants.
+They were instructed that their agents would be evaluated on a variety of problems constructed from the objects in the example arena.
+The contestants then needed to build an agent—using any method of their choosing—that they believed would succeed at a wide variety of tasks.
+The tasks that ultimately made up the evaluation suite were based on adaptations of tasks from comparative and developmental literature.
+More information can be found at http://animalai.org/animal-ai-olympics.
+For our work with the AAIO dataset we used a subset of 69 tasks instances that were solvable with simple goal-directed behaviour.
+The exact instances can be found in Appendix F.1.
+We include 68 participant agents for which we know nothing about their design.
+F.1 Animal AI Olympics Instances
+For our work with the AAIO dataset we used a subset of tasks instances that were solvable with simple goal-directed behaviour. The specific task instances we used were:
+[1-1-1, 1-1-2, 1-1-3, 1-2-1, 1-2-2, 1-2-3, 1-3-1, 1-3-2, 1-3-3, 1-4-1, 1-4-2, 1-4-3, 1-5-1, 1-5-2, 1-5-3, 1-6-1, 1-6-2, 1-6-3, 1-7-1, 1-7-2, 1-7-3, 1-8-1, 1-8-2, 1-8-3, 1-9-1, 1-9-2, 1-9-3, 1-10-1, 1-10-2, 1-10-3, 1-11-1, 1-11-2, 1-11-3, 1-12-1, 1-12-2, 1-12-3, 1-13-1, 1-13-2, 1-13-3, 1-14-1, 1-14-2, 1-14-3, 1-15-1, 1-15-2, 1-15-3, 1-16-1, 1-16-2, 1-16-3, 1-17-1, 1-17-2, 1-17-3, 7-1-1, 7-1-2, 7-1-3, 7-2-1, 7-2-2, 7-2-3, 7-3-1, 7-3-2, 7-3-3, 7-4-1, 7-4-2, 7-4-3, 7-5-1, 7-5-2, 7-5-3, 7-6-1, 7-6-2, 7-6-3]
+The types of arenas that these instances correspond to can be found at (http://animalai.org/animal-ai-olympics).
+Figure 22: Screenshots from two different instances in the Animal AI Olympics that we include in our analysis.
+Dimension Description Range/Values
+Reward Size The size of the reward (the sign is modified so lower is larger) $[0,1.9]$
+Reward Distance Euclidean distance from the agent’s start position to the reward $[0,5.3]$
+Reward Behind Whether the reward originates behind (1), left or right (0.5) or in front (0) of the agent $\{0,0.5,1\}$
+XPos Whether the reward is to the left (1-1), the right (1) or is centred (0) relative to the agent’s start position $\{-1,0,1\}$
+Table 10: AAIO Task Characterisation details. The dimensions on which instances vary in our measurement layouts.
+Note that these ranges determine the ranges of the capabilities, so that we have units.
+For instance, if distance to the reward were measured in metres then the navigation capability would have metres as a unit.
+Let the navigation capability of an agent be, say, 20 metres.
+Since it is defined with a logistic function of distance, that means that the agent can achieve excellent performance when well below 20, performing at chance around the 20 metre mark, and failing every instance when the reward is well beyond 20 metres away.
+Figure 23: Inferred cognitive profile elements, along with the success rate, for the 68 agents in the AAIO.
+The gradient colour red-yellow-green denotes how high the means are (green is best) while the gradient circles white-grey-black denote how high the standard deviations are (black is highest).
+Fig.
+23 shows the complete inference results for the 68 agents in the AAIO.
+The agents are ordered by the aggregate success rate at the collection of task instances.
+We see that, for the agents in the middle of the table in particular, our approach is able to decompose agent capabilities at this task into the constituent navigation and visual abilities.
+Bias has a more diverse behaviour throughout the table, with strong biases for extremely low scoring agents too, such as with agents ‘ahorizon’ and ‘Octopus’.
+Equally, we can see that the detected bias can account for poor performance.
+Compare ‘Thursday’ and ‘MadStorks’ which share a similar overall success rate (0.30 vs.
+0.28) despite the fact that ‘Thursday’ has much more navigation ability than ‘MadStorks’ and they have similar visual abilities.
+This is explained by ‘Thursday”s much larger bias towards moving to the right (0.73 vs 0.4) even if the reward is to the left.
+F.2 Cognitive Profiles
+To illustrate how a measurement layout identifies cognitive profiles, we explore a range of tasks in the Animal-AI Environment (Beyret et al., 2019; Voudouris et al., 2025).
+Full details of the environment can be found in Appendix B.
+we first select a simple navigation task where agents must navigate to a goal.
+In total, there are 69 selected navigation tasks and screenshots for two tasks are shown in Fig.
+22.
+We include 68 agents submitted to the AAI Olympics competition (Crosby et al., 2020).
+The profiles of agents will be composed of navigation and perception skills and whether they have any bias towards goals occuring on the left vs.
+the right.
+Following the analysis performed in (Burnell et al., 2022), we selected the set of meta-features $X$ to characterise task instances as rewardSize, rewardDistance, Xpos and rewardBehind.
+These represent the reward’s size, distance, and if it is left, right, in front, or behind the agent.
+Full details of the ranges of these variables and precise descriptions are presented in Appendix B, Table 10.
+The cognitive profile is composed of two capabilities, navigationAbility and visualAbility, respectively representing the skill to move around purposefully and the visual acuity of the agent, one bias rightLeftBias, representing any preference for left or right movements, and one robustness level noiseLevel, accounting for any unexplained variability.
+We then connect the task characterisation to the cognitive profile using a Measurement Layouts, that captures the interaction between demands and capabilities/bias.
+Fig.
+24(a) presents the Measurement Layout we built.
+The metafeatures $X$ appear as observed (shaded, rectangular) nodes in the model, appearing within the plate, with 69 instances.
+The elements of the cognitive profile appear as external latent (unshaded) variables.
+The derived nodes lie at the intersection of the meta-features and the external cognitive profile nodes.
+(a) Measurement layout for the Animal‑AI Olympics.
+(b) Inferred cognitive profiles for the 68 AAIO agents. Error bars show s.d.; dot area encodes mean accuracy; colour encodes noiseLevel (red=high).
+Figure 24:
+Not visible in the figure are the linking functions associated with each arrow.
+Broadly, the relevant metafeature task demands combine to increase difficulty, and are opposed by relevant capabilities, yielding partial performance estimates in the forms of probabilities.
+In this task, because success requires both identidying the goal and navigating too it, the partial performances are multiplied to calculate the final probability of success.
+Note that the ranges of these two abilities are given by the ranges of the task metafeatures, with navigationAbility in $[0,5.3]$ and with visualAbility in $[0,1.9]$ respectively, representing how distant and how big a reward—with their corresponding distance and size units—the agent could be measured as having in this task.
+This provides an intuitive property to the Measurement Layouts, in that the capabilities have meaningful units.
+We detail all of the nodes and linking functions comprising this Measurement Layout in Table 11. The function $\sigma$ represents the standard logistic function, while $\omega$ represents a weighting:
+$\omega[\alpha](b,c)=(1-\alpha)\cdot b+\alpha\cdot c$
+The value $\nu$ represents a noise prior that represents a constant model set to 1 - average performance.
+Finally, $ScaledBeta$ represents a scaled beta distribution to an interval different from $[0,1]$.
+Note that a beta distribution has two parameters that need to be estimated.
+We describe our choices for building the measurement Layout in this way in Appendix F.
+Name Type Range/Values Prior / Linking Function
+rewardSize Meta-Feature $[0,1.9]$
+rewardDistance Meta-Feature $[0,5.3]$
+rewardBehind Meta-Feature $\{0,0.5,1\}$
+XPos Meta-Feature $\{-1,0,1\}$
+navigationAbility Capability $[0,5.3]$ $ScaledBeta(1,1,0,5.3)$
+visualAbility Capability $[0,1.9]$ $ScaledBeta(1,1,0,1.9)$
+rightleftBias Bias $\infty$$\infty$] $\mathcal{N}(0,1)$
+noiseLevel Robustness $[0,1]$ $Uniform(0,1)$
+rightleftEffect INON [1,1][-1,1] ${\sf rightleftBias}\times{\sf XPos}$
+navigationPerformance INON $[0,1]$ $\sigma({\sf NavigationAbility}-{\sf rewardDistance}\times(\frac{1}{2}{\sf rewardBehind}+1)+{\sf rightleftEffect})$
+visualPerformance INON $[0,1]$ $\sigma({\sf visualAbility}-{\sf rewardSize})$
+taskPerformance Observed $[0,1]$ $p\sim Bernoulli(\omega[{\sf noiseLevel}]({\sf navigationPerformance}\times{\sf visualPerformance},\nu))$
+Table 11: Nodes in the AAIO Measurement Layout, including range/values, priors and linking function.
+F.3 Inference and Analysis
+Using the measurement layout and the 69 instances, we can use PyMC to infer the values of the cognitive profile for each of the 68 agents.
+Fig.
+24(b) summarises the inferred capabilities and noise levels.
+We see the two main capabilities are correlated for this population, however, we are also able to identify agents with weak navigation skill, but strong visual acuity (e.g.,cocel) and vice versa.
+We can also identify agents with strong biases (e.g., ‘Thursday’ and ‘KoozytHiperdyne’ with bias $1.14$ and 0.78-0.78).
+A full table with all of the inferred results for each agent is presented in Appendix F.
+Where most evaluations end, with a results table of identified scores, Measurement Layouts are novel insofar as they can go on to predict agent performance on unseen tasks instances.
+This happens by the forwards inference process of the HBN, using the To evaluate this, we randomly held out 20% of the task instances from the backwards inference stage and evaluated the predicted probabilities of success on them.
+We evaluate these probabilities using a Brier score (Brier, 1950).
+That is:
+$\textrm{Brier Score}(p,o)=\frac{1}{N}\sum_{i=1}^{N}(p_{i}-o_{i})^{2}$
+Where $p_{i}$ and $o_{i}$ are the predicted probabilities of success and observed outcomes respectively.
+The Brier Score is essentially the mean square error applied to probabilities and binary events.
+Due to the small size of the dataset, we repeated this process 15 times and took the mean Brier score for each agent.
+Figure 25 summarises these scores, which are lower than predictions based on aggregate performance for agents with success rates in the range $[0.2,0.7]$, exactly where there is variability to explain.
+Figure 25: Brier score of model predictions against agent success rates for AAIO.
+The orange parabola is the expected value for the Brier score on aggregate predictions BS=s(1s)BS=s(1-s) for success rate $s$.
+The orange points are the observed Brier Scores for each agent using the aggregate prediction.
+These fall on the parabola as expected.
+The blue points are the Brier scores for our predictions using the Measurement Layout.
+F.4 AAIO Patterns of Performance
+Figure 26: Agent characteristic grid displaying results for some selected agents.
+Top row, from left to right: ‘ironbar’, ‘sparklemotion’, ‘Juohmaru’.
+Bottom row, from left to right: ‘y.yang’, ‘daydayup’, ‘41Animals’.
+Note that the scales of the $x$-axis and $y$-axis are normalised and not the same as in the main body of text.
+To show the patterns of performance of several agents, we use a simple visualisation known as agent characteristic grids, as described in (Burnell et al., 2022).
+The plots place each feature as a dimension and the success rate represented in colours from green (success) to red (failure).
+Values are grouped into bins of appropriate width in the desired dimensions for analysis.
+Each cell also shows the number of observed instances in that bin.
+Fig.
+26 shows some of the agents discussed in the paper, as well as some other agents that will serve us to illustrate a few points.
+For instance, ‘ironbar’ has 99% accuracy (the grid is mostly green everywhere) for this set of tasks (like ‘Trrrr’, the winner of the competition) and is identified as saturating both capabilities in Fig.
+24(b).
+The next one, ‘sparklemotion’, with 36% accuracy, is very non-conformant, with navigationAbility $0.40\pm 0.38$ and visualAbility $0.59\pm 0.45$, and high noise $0.31\pm 0.13$.
+Proportionally, navigationAbility is worse as the grid looks even less conformant on the distance of the reward than the reward size.
+The rightmost on the top row is ‘Juohmaru’, with 54% accuracy, which is a very conformant agent, with red concentrating on the top and right of the grid (the objectively most difficult instances).
+Its navigation ability is $2.50\pm 1.08$ and visual ability is $0.92\pm 0.44$, and medium noise: $0.26\pm 0.17$.
+We have specifically chosen some particular agents in the second row for which the extracted cognitive profile does not capture the whole performance patterns completely.
+Leftmost on the second row, ‘y-yang’ has 70% accuracy and a very strange grid, with red cells in a diagonal, indicating a very non-conformant agent.
+This translates into navigationAbility $3.82\pm 0.45$ and visualAbility $1.68\pm 0.21$, but medium noise $0.22\pm 0.17$.
+This is so because overall the agent has high accuracy, so a bit of noise can explain the few anomalies on the top right of the grid.
+In the middle of the second row is ‘daydayup’ with 52% accuracy.
+This agent shows a higher navigation ability than visual ability, as can be seen in the grid, where the size of the reward is the predictive feature and reward distance is less relevant.
+This translates into navigationAbility $3.48\pm 1.28$ (high than average but with uncertainty) and visualAbility $0.92\pm 0.44$ (low), with some a high level of random noise affecting performance $0.48\pm 0.20$.
+Finally, ‘41Animals’ with 87% accuracy, shows relatively high values and low deviations for both navigationAbility $4.37\pm 0.59$ and visualAbility $1.81\pm 0.11$, with very little noise $0.05\pm 0.04$ despite having some non-conformant behaviour in the diagonal.
+In this case the left-right bias is higher ($0.45\pm 0.91$) and may account for this.
+In general, for those agents where accuracy is too low or too high, the differences are more subtle.
+Of course, the interesting part of Fig.
+23 is the agents in the middle.
+This suggests that estimating capabilities can benefit from a more focused analysis of instances around the demand thresholds.
+For ’Juohmaru’, this would amount to selecting instances by difficulty, instead of simply selecting all instances for analysis.
+Appendix G Code and Datasets
+Our experiments did not require a large amount of compute, We ran all of our experiments on Google Colaboratory (Google, ) using the free tier of service.
 All corresponding code, datasets and raw results will be released on publication to ensure ease of replication.